### v1

# latexOverwriting file

Abstract
For algorithms based on interacting particle systems that admit a mean-field description, convergence analysis is often more accessible at the mean-field level. In order to transpose convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics. In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling. Our approach extends Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems.
Keywords. Mean-field limits, Interacting particle systems, Consensus-Based Optimization, Consensus-Based Sampling, Coupling methods, Wasserstein stability estimates.
AMS subject classification. 35Q93, 65C35, 70F45, 35K55.
1 Introduction
This work focuses on the rigorous derivation of quantitative mean-field limits for recently proposed algorithms known as Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants. The former, initially proposed in [46], is an algorithm for global optimization tasks, while the latter, proposed in [11], can be employed for both sampling and optimization purposes. Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels. This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
Since Consensus-Based Sampling can be viewed as a variant on Consensus-Based Optimization, we focus predominantly on the former and then briefly discuss the latter.
1.1 Consensus-Based Optimization (CBO)
Given an objective function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, we consider the following version of Consensus-Based Optimization (CBO):
$\mathrm{d}X^{(j)}_{t}$ $=-\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}t+S\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}W^{(j)}_{t},\qquad t\geqslant 0,\qquad j\in,$ (1.1)
where $J>1$ is a natural number, :={$:=\{1,\dots,J\}$ and $\mu^{J}_{t}:=\frac{1}{J}\sum_{i=1}^{J}\delta_{X^{(i)}_{t}}$ denotes the empirical measure associated with the particle system. For a probability measure $\mu\in\mathcal{P}(\mathbf{R}^{d})$, the notation $\mathcal{M}_{\beta}(\mu)$ denotes the mean of $\mu$ after reweighting by $\operatorname{e}^{-\beta f}$. Its precise definition, together with its expression when $\mu=\mu^{J}_{t}$, is given by
$\mathcal{M}_{\beta}(\mu):=\frac{\int_{\mathbf{R}^{d}}x\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)},\qquad\mathcal{M}_{\beta}(\mu^{J}_{t}):=\frac{\sum_{j=1}^{J}X^{(j)}\operatorname{e}^{-\beta f\left(X^{(j)}_{t}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}_{t}\right)}}.$ (1.2)
The weighted mean $\mathcal{M}_{\beta}(\mu^{J}_{t})$ may be viewed as an approximation of the particle position $X^{(j)}_{t}$ such that $f(X^{(j)}_{t})$ is minimal. In (1.1), the processes $\bigl{(}W^{(j)}_{t}\bigr{)}_{j\in}$ are independent $\mathbf{R}^{d}$-valued Brownian motions and $S\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ is a globally Lipschitz continuous function. The two diffusion types commonly studied in the literature are isotropic diffusion [46, 10, 23], with $S(u)=\sqrt{2\theta}\lvert u\rvert\operatorname{I}_{d}$, and anisotropic or component-wise [13, 24] diffusion with $S(u)=\sqrt{2\theta}\operatorname{diag}(u)$. Here $\operatorname{I}_{d}\in\mathbf{R}^{d\times d}$ is the identity matrix and $\operatorname{diag}\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ the operator mapping a vector onto a diagonal matrix with the vector as its diagonal. It was shown in [13] that CBO with anisotropic diffusion is well-suited for high-dimensional global optimization problems, for example those arising in machine learning. Commonly a constant coefficient is included in front of the drift term in (1.1), but since this coefficient can be removed by rescaling time upon redefining the diffusion parameter $\theta$, we choose to omit it.
By design, CBO allows passage to the corresponding mean-field description. Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [37], it is possible to formally write down the corresponding mean-field equation:
[…]
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [29]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [34, 10, 23]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [10]. The distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large exponential parameter $\beta$. The proof in [10] covers also non-convex cost functions $f$.
The convergence, in an appropriate sense, of the empirical measure associated with the CBO finite-size particle system (1.1) to the solution of the mean-field equation (1.3) in the large particle limit $J\to\infty$ was first proved rigorously in a recent paper [34] using a compactness argument, as we discuss further in Subsection 1.3. However, the method of proof employed in [34] does not provide an explicit rate of convergence with respect to the number of particles. Obtaining such an explicit rate of convergence is highly desirable, as it would enable to transfer convergence results concerning the mean-field equation (1.3) to the finite ensemble size setting (1.1) by means of a triangle inequality argument, modulo appropriate interfacing of the metrics. This is the approach proposed in [23, Theorem 13], where the authors obtain a partial mean-field result in a set of high probability, enabling them to prove convergence with high probability to the global minimizer in the finite ensemble size setting. Here, we provide a stronger version of the mean-field limit result given in [23, Proposition 15], extending to more general diffusion terms and providing scope for global convergence results in a stronger metric than convergence in probability.
A number of extensions of CBO have been proposed in the literature. As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [13]. Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [30, 31, 41]. (On this subject, we point out that the results in [31] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [35], memory effects [49], and Lévy noise [36], as well as extensions to multi-objective optimization [40, 7] and global optimization with constraints [21, 22, 12, 8, 32, 4]. Moreover, a variant tailored for objective functions with several minima was proposed in [9], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [29, 17]. Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [48].
1.2 Consensus-Based Sampling (CBS)
Consensus-Based Sampling is a sampling method inspired by the Ensemble Kalman Sampler (EKS) [26] and by consensus-based approaches such as CBO. It is based on an interacting particle system with a noise term preconditioned by the square root of the empirical covariance. The method in the finite ensemble size and continuous time setting reads
$\mathrm{d}X^{(j)}_{t}=-\Bigl{(}X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\,\mathrm{d}W^{(j)}_{t},\qquad j\in,$ (1.4)
with parameters $\beta,\lambda>0$. Here, for a probability measure $\mu\in\mathcal{P}(\mathbf{R}^{d})$, the notation $\mathcal{C}_{\beta}(\mu)$ denotes the covariance of $\mu$ after reweighting by $\operatorname{e}^{-\beta f}$; see (1.9) in Subsection 1.5 for the precise definition. It is possible to formally write down the mean-field equation corresponding to (1.4):
[…]
When the parameters satisfy $\lambda=(1+\beta)^{-1}$, the interacting particle system (1.4) enables to generate approximate samples from the probability distribution with density proportional to $\operatorname{e}^{-f}$, that is from
$\rho_{f}:=\frac{\operatorname{e}^{-f}}{\mathcal{Z}},\qquad\mathcal{Z}:=\int_{\mathbf{R}^{d}}\operatorname{e}^{-f(x)}\,\mathrm{d}x.$
More precisely, the authors of [11] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$. In addition, it was noticed in [11] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [46]. Since the convergence analysis in [11] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$. This has so far been left as an open problem. Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [9], and for rare event estimation in [2].
1.3 Propagation of chaos
The CBO (1.1) and CBS (1.4) dynamics can both be recast as interacting particle systems of the form
$\mathrm{d}X^{(j)}_{t}=b\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}t+\sigma\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}W^{(j)},\qquad j\in\llbracket 1,J\rrbracket.$ (1.6)
with drift coefficient $b$ and diffusion coefficient $\sigma$ depending on $\mu$ through a convolution with appropriate kernels $K_{b}$ and $K_{\sigma}$:
$b(x,\mu)=\widetilde{b}(x,K_{b}\star\mu),\qquad\sigma(x,\mu)=\widetilde{\sigma}(x,K_{\sigma}\star\mu).$ (1.7)
Although the consensus-based dynamics can be written in the form (1.6), it is not possible to do so in such a way that $\widetilde{b},\widetilde{\sigma},K_{b}$ and $K_{\sigma}$ are all globally Lipschitz continuous and at the same time $K_{b}$ and $K_{\sigma}$ are globally bounded. Therefore, the classical McKean’s theorem [15, Theorem 3.1] is not directly applicable, and a different approach is needed. We briefly discuss hereafter the existing mean-field results for CBO, noting that no results have been obtained so far for CBS.
  • In [34], propagation of chaos for CBO is proved through a compactness argument based on Prokhorov’s theorem. However, this method comes without convergence rates in terms of the number of particles $J$.
  • Later, propagation of chaos by means of synchronous coupling was shown with optimal rates in [21] for a variant of CBO constrained to hypersurfaces.
  • Recent results in [23] also show a probabilistic mean-field approximation of the form
    $\sup_{t\in[0,T]}\mathbf{E}\left[\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{2}\mathsf{1}_{\Omega\setminus\Omega_{M}}\right]\leqslant\frac{C}{J},\qquad\Omega_{M}:=\left\{\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\max\left\{\left\lvert X^{(j)}_{t}\right\rvert^{4},\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{4}\right\}\geqslant M\right\}.$
    It is additionally proved in [23], by application of Markov’s inequality, that the set of realizations $\Omega_{M}$ has probability close to 0 for large $M$. Specifically, it holds that $\mathbf{P}(\Omega_{M})=\mathcal{O}(M^{-1})$.
  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [36]. In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1. There are, however, several differences between our results and those presented in [36]. Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [1, 18] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3). Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [36].
1.4 Our contributions
The main contributions of this paper are the following.
  • We obtain quantitative results on the mean-field limit for CBO and CBS. More precisely, we prove results with an explicit rate of convergence with respect to the number $J$ of particles. Investigating the roles of dimension and other problem parameters on the value of the constant prefactor is left for future work.
  • We present two approaches for dealing with the weighted mean (in the case of CBO) and the weighted covariance (in the case of CBS). One of these approaches is based on the careful use of indicator sets, whereas the other is based on stopping times. The latter approach is inspired by a method initially proposed in [33] in order to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
  • In order to prove mean-field results, we generalize existing theorems and prove new results on the well-posedness of the particle systems and their mean-field limit. We also present novel stability estimates for the weighted mean and the weighted covariance.
The rest of this paper is organized as follows. In Section 2, we present our main results for the consensus-based dynamics (1.1) and (1.4). They are based on a number of auxiliary results, which are stated precisely and proved in Section 3. These include Wasserstein stability estimates for the weighted mean and the weighted covariance, as well as moment estimates for the particle systems. In Section 4, we prove well-posedness results for the particle systems and their mean-field limits. Section 5 is reserved for extensions and perspectives for future work. Additional technical results are presented in Appendix A. The interdependence of the sections is illustrated in Figure 1.
Main theorems on the mean-field limits for CBO and CBS. Section 2 Wasserstein stability estimates for weighted moments. Subsection 3.1 Moment estimates for empirical measures. Subsection 3.2 Convergence of the weighted moments for i.i.d. samples. Subsection 3.3 Well-posedness result for the mean-field equations. Section 4 Technical bound on weighted moments. Subsection A.2
Figure 1: Interdependence of the results contained in this paper.
1.5 Notation
[…]
$\left(\int_{E}\lvert x\rvert^{p}\,\mu(\mathrm{d}x)\right)^{\frac{1}{p}}\leqslant R.$
Note that $\mathcal{P}_{q,R}(E)\subset\mathcal{P}_{p,R}(E)$ for all $0\leqslant p<q$. If the metric space $E$ is not specified, then it is implicitly assumed that $E=\mathbf{R}^{d}$. For example, we sometimes write $\mathcal{P}_{p,R}$ as a shorthand notation for $\mathcal{P}_{p,R}(\mathbf{R}^{d})$. For a probability measure $\mu\in\mathcal{P}(\mathbf{R}^{d})$ we use the following notation for the mean and covariance:
$\mathcal{M}(\mu)=\int_{\mathbf{R}^{d}}x\,\mu(\mathrm{d}x),\qquad\mathcal{C}(\mu)=\int_{\mathbf{R}^{d}}\bigl{(}x-\mathcal{M}(\mu)\bigr{)}\otimes\bigl{(}x-\mathcal{M}(\mu)\bigr{)}\,\mu(\mathrm{d}x).$ (1.8)
The notations $\mathcal{M}_{\beta}(\mu)$ and $\mathcal{C}_{\beta}(\mu)$ denote respectively the mean and covariance after reweighting by $\operatorname{e}^{-\beta f}$. More precisely, we use the notation
$\mathcal{M}_{\beta}(\mu)=\mathcal{M}(L_{\beta}\mu),\qquad\mathcal{C}_{\beta}(\mu)=\mathcal{C}(L_{\beta}\mu),\qquad L_{\beta}\mu:=\frac{\operatorname{e}^{-\beta f}\,\mu}{\int\operatorname{e}^{-\beta f}\,\mathrm{d}\mu}.$ (1.9)
For a probability measure $\mu\in\mathcal{P}(E)$ and a function $f\colon E\to\mathbf{R}$, we use the short-hand notation
$\mu[f]=\int_{E}f(x)\,\mu(\mathrm{d}x).$
By a slight abuse of notation, we sometimes write $\mu[f(x)]$ instead of $\mu[f]$ for convenience. For example, for a probability measure $\mu\in\mathcal{P}_{2}(\mathbf{R})$, the notation $\mu\left[x^{2}\right]$ refers to the second raw moment of $\mu$. Furthermore, for $J\in\mathbf{N}^{+}$ we denote by $\mu^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{X^{(j)}}$ and $\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}}$ the empirical measures of the particles $X^{(1)},\dots,X^{(J)}$ and $\overline{X}^{(1)},\dots,\overline{X}^{(J)}$, respectively. Finally, throughout this paper, the notation $\Omega$ refers to the sample space, and the notation $C$ refers to a constant whose exact value is irrelevant in the context and may change from line to line.
2 Main results
This section is organized as follows. In Subsection 2.1, we present the assumptions used throughout this work. Then, after stating well-posedness results for CBO and CBS in Subsection 2.2, we prove in Subsection 2.3 a key preparatory result for the mean-field limits. We then prove quantitative mean-field results for Consensus-Based Optimization and Sampling in Subsections 2.4 and 2.5, respectively.
2.1 Main assumptions
We assume throughout this paper that the diffusion operator $S\colon\mathbf{R}^{d}\rightarrow\mathbf{R}^{d\times d}$ appearing in the CBO dynamics (1.1) is globally Lipschitz continuous. For the objective function $f$, we use the following assumption.
Assumption H.
We assume the following properties for the target function $f$.
  1. (H1)
    The function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ is bounded from below by $f_{\star}=\inf f$ and satisfies
    $\forall x,y\in\mathbf{R}^{d},\qquad|f(x)-f(y)|\leqslant L_{f}\bigl{(}1+|x|+|y|\bigr{)}^{s}|x-y|,$ (2.1)
    for some constants $L_{f}>0$ and $s\geqslant-1$.
  2. (H2)
    There exist constants $c_{\ell},c_{u}>0$ and $u\geqslant\ell\geqslant 0$ such that
    $\forall x\in\mathbf{R}^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
    $\forall x\in\mathbf{R}^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
Remark 2.1.
A few remarks are in order.
  • When $\ell=u=0$, assumption (H2) implies that $f$ is bounded from above.
  • Observe that (H1) implies (2.2a) with $u=s+1$.
  • In the particular case where $s=1$ and either $u=\ell=0$ or $u=\ell=2$, the assumptions in Assumption H coincide with those in [10]. The well-posedness and mean-field results proved in this paper generalize beyond this setting, as explained below.
In the rest of this paper, we denote by $\mathcal{A}(s,\ell,u)$ the set of objective functions that satisfy Assumption H with these parameters. Our main results hold only for $f\in\mathcal{A}(s,\ell,\ell)$, i.e. under Assumption H with $\ell=u$. Many of the results we present hold only for values of $p$ in a range dictated by the interval of validity of the Wasserstein stability estimates presented in Subsection 3.1. In order to avoid repetitions, we introduce the associated notation
$p_{\mathcal{M}}(s,\ell):=\begin{cases}s+2,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$},\end{cases}\qquad p_{\mathcal{C}}(s,\ell):=\begin{cases}s+3,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$}.\end{cases}$ (2.3)
2.2 Well-posedness for CBO and CBS
In this section, we state well-posedness results for CBO and CBS, referring for the proofs to Section 4. While for CBO, the following theorems are only generalizations of already existing results [10], the well-posedness proofs for the CBS particle and mean-field systems are new.
Theorem 2.2 (Existence and uniqueness for the particle models).
Assume that $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ is locally Lipschitz continuous and fix $J\in\mathbf{N}^{+}$. Then, the stochastic differential equations (1.1) and (1.4) have unique strong solutions $\bigl{\{}{\mathbf{X}}_{t}^{(J)}\,|\,t\geqslant 0\bigr{\}}$ for any initial condition ${\mathbf{X}}_{0}^{(J)}$ which is independent of the Brownian motions $\mathbf{W}_{t}=\bigl{(}W^{(1)}_{t},\dots,W^{(J)}_{t}\bigr{)}$. The solutions are almost surely continuous.
In the case of CBO, Theorem 2.2 was already proved in [10, Theorem 2.1]. We now present well-posedness results for the mean-field SDEs, beginning with CBO. For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [10, Theorems 3.1 and 3.2].
Theorem 2.3 (CBO: Existence and uniqueness for the mean-field SDE).
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant-1$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.3) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
[…]
and the function $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$. Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C\left([0,T],\mathcal{P}_{p}(\mathbf{R}^{d})\right)$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
$\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}\Bigr{)}+\frac{1}{2}\nabla\cdot\nabla\cdot\Bigl{(}D(\overline{\rho},x)\overline{\rho}\Bigr{)}.$ (2.5)
where $D(\overline{\rho},x):=S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}^{\mathsf{T}}$.
Theorem 2.4 (CBS: Existence and uniqueness for the mean-field SDE).
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant-1$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.5) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,\qquad\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\overline{\rho}_{t})\right\rvert<\infty,\qquad\sup_{t\in[0,T]}\left\lVert\mathcal{C}_{\beta}(\overline{\rho}_{t})\right\rVert<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.6)
and the functions $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$, $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$. Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C([0,T],\mathcal{P}_{p}(\mathbf{R}^{d}))$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
$\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}+\lambda^{-1}\mathcal{C}_{\beta}(\overline{\rho})\nabla\overline{\rho}\Bigr{)}.$ (2.7)
2.3 Key preparatory lemma
We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of a standard non-asymptotic version of the central limit theorem known as Bikelis’ theorem, see for example [28, Theorem 3.7].
Lemma 2.5 (Bound on the probability of large excursions).
Let $(Z_{j})_{j\in\mathbf{N}^{+}}$ be a family of i.i.d. $\mathbf{R}$-valued random variables such that $\mathbf{E}\left[\left\lvert Z_{1}\right\rvert^{3}\right]<\infty$. Then for all $R>\mathbf{E}\bigl{[}\left\lvert Z_{1}\right\rvert\bigr{]}$, there exists a constant $C>0$ such that
$\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant\frac{C}{J^{2}}.$
Proof.
Denote by $m$ and $\sigma^{2}$ the mean and variance of $Z_{j}$, respectively. Bikelis’ theorem provides the following non-asymptotic error estimate:
$\forall z\in\mathbf{R},\qquad\bigl{\lvert}\phi_{J}(z)-\phi(z)\bigr{\rvert}\leqslant\frac{1}{\sqrt{J}}\,\left(\frac{\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]}{\sigma^{3}(1+\lvert z\rvert)^{3}}\right),\qquad\phi_{J}(z):=\mathbf{P}\left[\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\frac{Z_{j}-m}{\sigma}\leqslant z\right],$
where $\phi$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ Gaussian law. Thus, if $R>m$, we obtain
$\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]$ $=\mathbf{P}\left[\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\frac{Z_{j}-m}{\sigma}\geqslant\sqrt{J}\left(\frac{R-m}{\sigma}\right)\right]$
$=1-\phi_{J}\left(\sqrt{J}k\right),\qquad k:=\frac{R-m}{\sigma}$
$\leqslant 1-\phi\left(\sqrt{J}k\right)+\frac{1}{\sqrt{J}}\,\left(\frac{\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]}{\sigma^{3}(1+\sqrt{J}k)^{3}}\right)=\mathcal{O}(J^{-2}),$
which completes the proof. ∎
2.4 Quantitative mean-field limit for CBO
We now present and prove the mean-field results, starting with CBO. Synchronous coupling for McKean-Vlasov diffusions with drift $b\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d}$ and diffusion $\sigma\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d\times d}$ assumes the following setting [14, 15]. Given i.i.d. Brownian motions $\bigl{(}W^{(j)}\bigr{)}_{j\in\mathbf{N}^{+}}$ and initial conditions $\bigl{(}X^{(j)}_{0}\bigr{)}_{j\in\mathbf{N}^{+}}$ sampled i.i.d. from some $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$, we consider for each $J\in\mathbf{N}^{+}$ the interacting particle system
$\forall j\in,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
to which we couple the system of i.i.d. mean-field McKean-Vlasov diffusions
$\forall N\in\mathbf{N}^{+},\forall j\in,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
driven by the same Brownian motions and initial conditions as (2.8). In the case of CBO, the drift and diffusion coefficients for (2.8) and (2.9) are given by
$b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=S\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)},$ (2.10)
where the diffusion operator $S\colon\mathbf{R}^{d}\rightarrow\mathbf{R}^{d\times d}$ is, as always in this work, a globally Lipschitz continuous function.
Theorem 2.6 (Mean-field limit for CBO).
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$, and consider the systems (2.8) and (2.9) with the coefficients given by (2.10). Then for all $p\geqslant 2$ such that
$\frac{q}{3}\geqslant p\geqslant p_{\mathcal{M}}(s,\ell),$ (2.11)
where $p_{\mathcal{M}}$ is given in (2.3), there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\min\left\{\frac{p}{2},2-\frac{2p}{q}\right\}}.$
Remark 2.7.
In particular, we obtain a decay bound of order $\mathcal{O}\left(\frac{1}{J}\right)$ if $p=2$. If $p\geqslant 4$, then for each $\varepsilon>0$, decay of order $\mathcal{O}\left(\frac{1}{J^{2-\varepsilon}}\right)$ holds provided that $\overline{\rho}_{0}\in\mathcal{P}_{q}(^{d})$ with $q\geqslant\frac{2p}{\varepsilon}\vee 3p$.
Remark 2.8.
Theorem 2.6 implies pathwise Wasserstein-$p$ chaos. More precisely, it holds that
$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{2}{p}-\frac{2}{q}\right\}}.$
where $\rho^{J}_{[0,T]}$ is the pathwise law of the microscopic particle system over $[0,T]$ and $\overline{\rho}_{[0,T]}$ is the pathwise mean-field law over $[0,T]$. That is to say,
[…]
Here, $W_{p}$ denotes the pathwise Wasserstein-$p$ distance, associated with the following metric on $\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}$:
$d(\mathbf{X},\mathbf{Y})=\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}-Y^{(j)}_{t}\right\rvert,\qquad\mathbf{X},\mathbf{Y}\in\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}.$
See [14, Definitions 3.1 and 3.5] for more details.
Before presenting the proof, we summarize below a number of useful auxiliary results which are valid for $f\in\mathcal{A}(s,\ell,\ell)$. These results are stated precisely and proved in Section 3.
  1. (a)
    Wasserstein stability estimate for the weighted mean. We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant p_{\mathcal{M}}(s,\ell)$, where $p_{\mathcal{M}}(s,\ell)\in^{+}$ is given in (2.3), there exists $\widetilde{L}_{\mathcal{M}}=\widetilde{L}_{\mathcal{M}}(R,p)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,W_{p}(\mu,\nu).$
  2. (b)
    Moment estimate for the particle system. We prove in Lemma 3.5 that, if the initial condition satisfies $\overline{\rho}_{0}\in\mathcal{P}_{q}(^{d})$, then there exists a constant $\kappa>0$ independent of $J$ such that
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}\right\rvert^{q}\right]\quad\vee\quad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\mu^{J}_{t})\right\rvert^{q}\right]\leqslant\kappa.$
  3. (c)
    Convergence of the weighted mean for i.i.d. samples. We prove in Lemma 3.7 that, for any reals $r>p>0$ and any $\mu\in\mathcal{P}_{r}(\mathbf{R}^{d})$, there is $C$ such that
    $\mathbf{E}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
Proof of Theorem 2.6.
Fix $j\in$. As in Sznitman’s argument [15, Theorem 3.1], we begin by writing
$\bigl{\lvert}X_{t}^{(j)}-\overline{X}_{t}^{(j)}\bigr{\rvert}^{p}\leqslant$  $\,2^{p-1}\left\lvert\int_{0}^{t}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
$+2^{p-1}\left\lvert\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$
By the Burkholder–Davis–Gundy inequality, see [44, Theorem 7.3], it holds for any $t\in[0,T]$ that
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant 2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert\,\mathrm{d}s\right)^{p}$
$\qquad+C_{\rm BDG}2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad+\,C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (2.12)
By the triangle inequality, the first expectation on the right-hand side can be bounded as follows:
$\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$ $\leqslant 2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\,\right\rvert^{p}\mathrm{d}s$
$\qquad+2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s.$ (2.13)
Bounding the first term in (2.13)
From (2.10) we have
$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left\lvert\Bigl{(}\overline{X}^{(j)}_{s}-X^{(j)}_{s}+\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\Bigr{)}\right\rvert^{p}$
$\leqslant 2^{p-1}\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+2^{p-1}\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}.$ (2.14)
In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [15, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
$\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\},$ (2.15)
where $R>\mathbf{E}[Z_{j}]$ is fixed. Define
$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}$
Since $q\geqslant 3p$, Theorem 2.3 implies that $\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]<\infty.$ Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
$\forall t\in[0,T],\qquad\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\Omega_{J,t}\right]\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant\frac{C}{J^{2}}.$
We can split the second term on the right-hand of (2.14) as follows:
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]$
$\qquad+\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right].$ (2.16)
Let us now bound each term on the right-hand side of 2.16 separately.
  • Since the $p$-th moment of $\overline{\mu}_{s}^{J}$ is bounded from above by $R$ for all $\omega\in\Omega\setminus\Omega_{J,s}$, we can use the stability estimate on the weighted mean given (see Item (a) and Corollary 3.3) to obtain
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\left(\mu^{J}_{s},\overline{\mu}^{J}_{s}\right)^{p}\right]$
    for the first term in (2.16). Since in addition
    $\mathbf{E}\left[W_{p}\left(\mu^{J}_{s},\overline{\mu}^{J}_{s}\right)^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
    it follows that
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}.$
  • We now bound the second term in (2.16). From Item (b) (see Lemma 3.5) and (2.4) we have
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right)\leqslant 2^{q}\kappa.$
    Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{2(q-p)}{q}}.$
We have thus shown that
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{2(q-p)}{q}},$
which in turn by (2.14) implies that
$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{2(q-p)}{q}}.$
Bounding the second term in (2.13)
For the second term on the right-hand side of (2.13), we have by Item (c) (see Lemma 3.7) that
$\mathbf{E}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}=\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\rho}_{s}\right)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}}.$
Conclusion
Combining the bounds on both terms in (2.13), we deduce
$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
Using that $S\colon\mathbf{R}\rightarrow\mathbf{R}^{d\times d}$ is globally Lipschitz continuous, we obtain analogously
$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
Substituting in (2.12), we finally obtain
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
The conclusion then follows from Grönwall’s lemma. ∎
Remark 2.9.
The proof is similar to that of McKean’s theorem [15, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
Remark 2.10.
A crucial detail in the proof is that Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below. This allows us to define $\Omega_{J,t}$ in (2.15) only in terms of the i.i.d. mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ need not be included in the definition of $\Omega_{J,t}$
Remark 2.11.
In his famous lecture notes [47], Sznitman shows for a laboratory example that
$\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [16]. Although CBO does not fit within Sznitman’s simplified framework, Item (c) enables to show a bound with the same rate for the second term on the right-hand side of (2.13),
$\mathbf{E}\biggl{[}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,t}}\biggr{]}\leqslant\frac{C}{J^{\frac{p}{2}}}.$
The limiting factor, for the overall rate of convergence in Theorem 2.6, is the bound on the probability $\mathbf{P}[\Omega_{J,t}]$.
2.5 Quantitative mean-field limit for CBS
Using the same reasoning as that in the proof of Theorem 2.6, we can prove a mean-field result for the CBS dynamics (1.4). We state the theorem and, given the similarity of the proof with that of Theorem 2.6, only list the key ingredients needed for the proof.
Theorem 2.12 (Mean-field limit for CBS).
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s\geqslant-1$, $\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$, and consider the systems (2.8) and (2.9) with the coefficients given by
$b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}(\mu)}.$ (2.17)
Then for all $p\geqslant 2$ such that
$\frac{q}{3}\geqslant p\geqslant 2p_{\mathcal{M}}(s,\ell)=\begin{cases}2s+4,\,&\text{if $\ell=0$},\\ 2,\,&\text{if $\ell>0$},\end{cases}$
there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{\frac{p}{2},2-\frac{2p}{q}\right\}}.$
We summarize below the additional auxiliary results necessary for the proof, which are valid for an objective function $f\in\mathcal{A}(s,\ell,\ell)$. These results are stated precisely and proved in Section 3.
  • Wasserstein stability estimate for the weighted covariance. We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant 2p_{\mathcal{M}}(\ell,s)$, there exists $\widetilde{L}_{\sqrt{\mathcal{C}}}(R,p)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$
  • Moment estimate for the particle system. We prove in Lemma 3.5 that, under the assumptions of Theorem 2.12, there exists a constant $\kappa>0$ independent of $J$ such that for all $J\in\mathbf{N}^{+},\qquad$
    $\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}\right\rvert^{q}\right]\quad\vee\quad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\mu^{J}_{t})\right\rvert^{q}\right]\quad\vee\quad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\sqrt{\mathcal{C}_{\beta}(\mu^{J}_{t})}\right\rvert^{q}\right]\leqslant\kappa.$
  • Convergence of the square root of the weighted covariance for i.i.d. samples. We prove in Lemma 3.8 that, for any reals $r>2p>0$ and any $\mu\in\mathcal{P}_{r}(\mathbf{R}^{d})$ such that $\mathcal{C}_{\beta}(\mu)\succ 0$, then there is $C$ such that
    $\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
  • No collapse in finite time. In order to apply the third item with $\mu=\overline{\rho}_{t}$, we prove in Proposition 3.9 that the assumption $\mathcal{C}(\overline{\rho}_{0})\succ 0$ implies that
    $\eta:=\inf_{t\in[0,T]}\lambda_{\text{min}}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$
3 Auxiliary results
This section is organized as follows. In Subsection 3.1, we improve on existing Wasserstein stability estimates for the weighted mean and prove novel ones for the weighted covariance and the square root of the weighted covariance. In Subsection 3.2, we present moment estimates for the empirical measures associated with the CBO (1.1) and CBS (1.4) dynamics. In Subsection 3.3, we prove convergence estimates for the weighted moments when particle positions are drawn independently from a given probability distribution, using results from the statistics literature. Finally, in Subsection 3.4, we prove that if $\mathcal{C}_{\beta}(\overline{\rho}_{0})\succ 0$, then the covariance matrix $\mathcal{C}_{\beta}(\overline{\rho}_{t})\succ 0$ associated with the solution to the mean-field CBS dynamics (1.5) remains strictly positive definite for all times.
3.1 Wasserstein stability estimates for the weighted moments
In this subsection, we generalize [10, Lemma 3.2] to the more general conditions in Assumption H and, more importantly, to the weighted covariance and its square root.
Proposition 3.1 (Basic stability estimates for the weighted mean and the weighted covariance).
Suppose that $f\in\mathcal{A}(s,\ell,u)$ with $s\geqslant-1$ and $u\geqslant\ell\geqslant 0$. Then:
  1. (a)
    For all $R>0$ and for all $p\geqslant p_{\mathcal{M}}(s,\ell)$, where $p_{\mathcal{M}}(s,\ell)$ is given in (2.3), there exists $L_{\mathcal{M}}=L_{\mathcal{M}}(p,R)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant L_{\mathcal{M}}\,W_{p}(\mu,\nu).$ (3.1)
  2. (b)
    For all $R>0$ and for all $p\geqslant p_{\mathcal{C}}(s,\ell)$, where $p_{\mathcal{C}}(s,\ell)$ is given in (2.3), there exists $L_{\mathcal{C}}=L_{\mathcal{C}}(p,R)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L_{\mathcal{C}}\,W_{p}(\mu,\nu).$ (3.2)
Proof.
We prove first (a) and then (b). Both proofs rely on the technical result Lemma A.1 in the appendix.
Proof of (a)
In the notation of Lemma A.1 we can write
$\mathcal{M}_{\beta}(\mu)=\frac{\mu[g]}{\mu[h]},\qquad g(x):=\operatorname{e}^{-\beta f(x)}x,\qquad h(x):=\operatorname{e}^{-\beta f(x)}.$
Condition A.1 is satisfied for $\varsigma:=s+1$ if $\ell=0$ or $\varsigma:=0$ if $\ell>0$. Therefore, the claim follows from Lemma A.1.
Proof of (b)
First observe that
$\mathcal{C}_{\beta}(\mu)=\frac{\mu[g]}{\mu[h]}-\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu),\qquad g(x):=\operatorname{e}^{-\beta f(x)}(x\otimes x),\qquad h(x):=\operatorname{e}^{-\beta f(x)}.$
By the triangle inequality, we have that
$\bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\bigr{\rVert}_{\rm F}\leqslant\Bigl{\lVert}\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\otimes\mathcal{M}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}+\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert_{\rm F}\,.$ (3.3)
We shall use that for all $(x,y)\in\mathbf{R}^{d}\times\mathbf{R}^{d}$, it holds that
$\left\lVert x\otimes x-y\otimes y\right\rVert_{\rm F}$ $=\left\lVert x(x-y)^{\mathsf{T}}+(x-y)y^{\mathsf{T}}\right\rVert_{\rm F}$
$\leqslant\left\lVert x(x-y)^{\mathsf{T}}\right\rVert_{\rm F}+\left\lVert(x-y)y^{\mathsf{T}}\right\rVert_{\rm F}=\lvert x\rvert\cdot\lvert x-y\rvert+\lvert y\rvert\cdot\lvert x-y\rvert=\bigl{(}\lvert x\rvert+\lvert y\rvert\bigr{)}\lvert x-y\rvert.$ (3.4)
Using this inequality, that $p_{\mathcal{C}}\geqslant p_{\mathcal{M}}$ and (a), we obtain for the first term in (3.3) that
$\Bigl{\lVert}\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\otimes\mathcal{M}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}$ $\leqslant\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\cdot\bigl{\lvert}\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\bigr{\rvert}$
$\leqslant L_{\mathcal{M}}\,\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\,W_{p}(\mu,\nu)$
$\leqslant L_{\mathcal{M}}^{2}\,\Bigl{(}W_{p}(\mu,\delta_{0})+W_{p}(\nu,\delta_{0})\Bigr{)}\,W_{p}(\mu,\nu)\leqslant 2L_{\mathcal{M}}^{2}R\,W_{p}(\mu,\nu).$
In order to bound the second term in (3.3), we again use (3.4) to deduce that the function $g$ satisfies the local Lipschitz estimate
$\forall x,y\in\mathbf{R}^{d},\qquad\bigl{\lvert}g(x)-g(y)\bigr{\rvert}\leqslant L\bigl{(}1+|x|+|y|\bigr{)}^{\varsigma}\lvert x-y\rvert,\qquad\varsigma:=\begin{cases}s+2\quad&\text{ if\leavevmode\nobreak\ $\ell=0$},\\ 0\quad&\text{ if\leavevmode\nobreak\ $\ell>0$}.\end{cases}$
We deduce that the functions $g$ and $h$ satisfy the assumption (A.1) of Lemma A.1 with this $\varsigma$. Therefore, it holds that
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\frac{\mu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\mu[\operatorname{e}^{-\beta f(x)}]}-\frac{\nu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\nu[\operatorname{e}^{-\beta f(x)}]}\right\rVert_{\rm F}\leqslant CW_{p}(\mu,\nu),$
which completes the proof. ∎
Proposition 3.2 (Basic stability estimate for the square root covariance).
Suppose that $f\in\mathcal{A}(s,\ell,u)$ with $s\geqslant 0$ and $u\geqslant\ell\geqslant 0$. Then for all $R>0$ and all $p\geqslant 2p_{\mathcal{M}}(s,\ell)$, there exists $L_{\sqrt{\mathcal{C}}}=L_{\sqrt{\mathcal{C}}}(p,R)>0$ such that
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant L_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$ (3.5)
Proof.
It is sufficient to check the claim for measures $\mu$ and $\nu$ of the form
$\mu^{J}=\frac{1}{J}\sum_{j=1}^{J}\delta_{X^{(j)}},\qquad\nu^{J}=\frac{1}{J}\sum_{j=1}^{J}\delta_{Y^{(j)}},\qquad J\in\mathbf{N}^{+}.$ (3.6)
Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$. By [51, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$. Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [51, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$. Then
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu})-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant$ $\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu^{J})}-\sqrt{\mathcal{C}_{\beta}(\nu^{J})}\right\rVert_{\rm F}$
$\qquad+\left\lVert\sqrt{\mathcal{C}_{\beta}(\nu)}-\sqrt{\mathcal{C}_{\beta}(\nu^{J})}\right\rVert_{\rm F}+\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\mu^{J})}\right\rVert_{\rm F}.$
The first term is bounded from above by $CW_{p}(\mu^{J},\nu^{J})$ by the base case, while the other two terms converge to 0 in the limit as $J\to\infty$ by Proposition 3.1. Taking the limit $J\to\infty$, we deduce that
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant CW_{p}(\mu,\nu).$
Proof of the statement for empirical measures
Fix $R>0$. By [52, p.5], the Wasserstein distance between empirical measures $\mu^{J}$ and $\nu^{J}$ of the form (3.6) is equal to
$W_{p}(\mu^{J},\nu^{J})=\min_{\sigma\in\mathcal{S}_{J}}\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{\sigma(j)}\right\rvert^{p}\right)^{\frac{1}{p}},$
where $\mathcal{S}_{J}$ denotes the set of permutations in $\{1,\dotsc,J\}$. Thus, the claim will follow if we can prove that, for any pair $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,R}$ of the form (3.6), it holds that
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu^{J})}-\sqrt{\mathcal{C}_{\beta}(\nu^{J})}\right\rVert_{\rm F}\leqslant C\,\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{(j)}\right\rvert^{p}\right)^{\frac{1}{p}}.$ (3.7)
We henceforth drop the superscript $J$ in $\mu^{J},\nu^{J}$ for simplicity of notation, and write $\mathbf{X}=\left(X^{(1)},\dotsc,X^{(J)}\right)$ and $\mathbf{Y}=\left(Y^{(1)},\dotsc,Y^{(J)}\right)$. First note that
$\mathcal{C}_{\beta}(\mu)=\sum_{j=1}^{J}w_{j}(\mathbf{X})\Bigl{(}X^{(j)}-\mathcal{M}_{\beta}(\mu)\Bigr{)}\otimes\Bigl{(}X^{(j)}-\mathcal{M}_{\beta}(\mu)\Bigr{)},\qquad w_{j}(\mathbf{X}):=\frac{\operatorname{e}^{-\beta f(X^{(j)})}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f(X^{(k)})}}.$
It is easily seen that $\mathcal{C}_{\beta}(\mu)=M_{\mathbf{X}}M_{\mathbf{X}}^{\mathsf{T}}$, where $M_{\mathbf{X}}$ is the following $\mathbf{R}^{d\times J}$ matrix
$M_{\mathbf{X}}:=\begin{pmatrix}\sqrt{w_{1}(\mathbf{X})}\left(X^{(1)}-\mathcal{M}_{\beta}(\mu)\right)&\ldots&\sqrt{w_{J}(\mathbf{X})}\left(X^{(J)}-\mathcal{M}_{\beta}(\mu)\right)\end{pmatrix}.$
Proceeding in the same manner, we construct a matrix $M_{\mathbf{Y}}\in\mathbf{R}^{d\times J}$ such that $\mathcal{M}_{\beta}(\nu)=M_{\mathbf{Y}}M_{\mathbf{Y}}^{\mathsf{T}}$. A result by Araki and Yamagami [3], later generalized by Kittaneh [39] and Bhatia [6], states for any two matrices $A$ and $B$ with the same shape, it holds that
$\left\lVert\sqrt{A^{\mathsf{T}}A}-\sqrt{B^{\mathsf{T}}B}\right\rVert_{\rm F}\leqslant\sqrt{2}\left\lVert A-B\right\rVert_{\rm F},$
see also [5, Theorem VII.5.7] for a textbook presentation. This yields
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}$
$\qquad\leqslant\sqrt{2}\left(\sum_{j=1}^{J}\left\lvert\sqrt{w_{j}(\mathbf{X})}\left(X^{(j)}-\mathcal{M}_{\beta}(\mu)\right)-\sqrt{w_{j}(\mathbf{Y})}\left(Y^{(j)}-\mathcal{M}_{\beta}(\nu)\right)\right\rvert^{2}\right)^{\frac{1}{2}}$
$\qquad\leqslant\sqrt{2}\left(\sum_{j=1}^{J}\left\lvert\sqrt{w_{j}(\mathbf{X})}X^{(j)}-\sqrt{w_{j}(\mathbf{Y})}Y^{(j)}\right\rvert^{2}\right)^{\frac{1}{2}}+\sqrt{2}\left(\sum_{j=1}^{J}\left\lvert\sqrt{w_{j}(\mathbf{X})}\mathcal{M}_{\beta}(\mu)-\sqrt{w_{j}(\mathbf{Y})}\mathcal{M}_{\beta}(\nu)\right\rvert^{2}\right)^{\frac{1}{2}},$
where we applied the triangle inequality for the Frobenius norm. Let $\pi\in\Pi(\mu,\nu)\subset\mathcal{P}(\mathbf{R}^{d}\times\mathbf{R}^{d})$ denote the probability measure
$\pi=\frac{1}{J}\sum_{j=1}^{J}\delta_{\left(X^{(j)},Y^{(j)}\right)}$
and set $\gamma=\frac{\beta}{2}$. We may rewrite the previous inequality as
$\frac{1}{\sqrt{2}}\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}$ $\leqslant\left(\iint_{\mathbf{R}^{d}\times\mathbf{R}^{d}}\left\lvert\frac{\operatorname{e}^{-\gamma f(x)}x}{\sqrt{\mu\left[\operatorname{e}^{-\beta f}\right]}}-\frac{\operatorname{e}^{-\gamma f(y)}y}{\sqrt{\nu\left[\operatorname{e}^{-\beta f}\right]}}\right\rvert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{1}{2}}$
$\qquad+\left(\iint_{\mathbf{R}^{d}\times\mathbf{R}^{d}}\left\lvert\frac{\operatorname{e}^{-\gamma f(x)}\mathcal{M}_{\beta}(\mu)}{\sqrt{\mu\left[\operatorname{e}^{-\beta f}\right]}}-\frac{\operatorname{e}^{-\gamma f(y)}\mathcal{M}_{\beta}(\nu)}{\sqrt{\nu\left[\operatorname{e}^{-\beta f}\right]}}\right\rvert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{1}{2}}.$
The conclusion then follows from Lemma A.2 in the appendix:
  • For the first term we let $g(x)=\operatorname{e}^{-\gamma f(x)}x$, $h(x)=\operatorname{e}^{-\beta f(x)}$, and $\mathcal{S}(\mu)=1$. The conditions of Lemma A.2 are satisfied for $\varsigma:=s+1$ if $\ell=0$ and $\varsigma:=0$ if $\ell>0$.
  • For the second term let $g(x)=\operatorname{e}^{-\gamma f(x)}$, $h(x)=\operatorname{e}^{-\beta f(x)}$ and $\mathcal{S}(\mu)=\mathcal{M}_{\beta}(\mu)$. The conditions of Lemma A.2 are satisfied for $\varsigma:=s$ if $\ell=0$ and $\varsigma:=0$ if $\ell>0$.
This completes the proof. ∎
The key for our mean-field proofs is the following improvement of Propositions 3.1 and 3.2. Recall that
$p_{\mathcal{M}}:=p_{\mathcal{M}}(s,\ell):=\begin{cases}s+2,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$}.\end{cases}$
Corollary 3.3 (Improved stability estimates).
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$. Then
  1. (a)
    For all $R>0$ and for all $p\geqslant p_{\mathcal{M}}$, there exists $\widetilde{L}_{\mathcal{M}}=\widetilde{L}_{\mathcal{M}}(R,p)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,W_{p}(\mu,\nu).$
  2. (b)
    For all $R>0$ and $p\geqslant 2p_{\mathcal{M}}$, there exists $\widetilde{L}_{\mathcal{C}}=\widetilde{L}_{\mathcal{C}}(R,p)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\Bigr{\rVert}_{\rm F}\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$
Remark 3.4.
Observe that in contrast to Propositions 3.1 and 3.2, no upper bounds on the $p$-th moments of $\nu$ are needed in Corollary 3.3, it is only required that they are finite.
Proof.
Let $\mathcal{T}:=\mathcal{M}_{\beta}$ or $\mathcal{T}:=\sqrt{\mathcal{C}_{\beta}}$, and denote $p_{\mathcal{T}}=p_{\mathcal{M}}$ if $\mathcal{T}=\mathcal{M}_{\beta}$ or $p_{\mathcal{T}}=2p_{\mathcal{M}}$ if $\mathcal{T}=\sqrt{\mathcal{C}_{\beta}}$. Note that the following statements hold in either case.
  • For all $R>0$ and $p\geqslant p_{\mathcal{T}}$, there exists $L_{\mathcal{T}}(R)>0$ such that
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)-\mathcal{T}(\nu)\bigr{\rVert}\leqslant L_{\mathcal{T}}(R)\,W_{p}(\mu,\nu).$ (3.8)
[…]
  • For all $p\geqslant p_{\mathcal{T}}$, there is $K>0$ such that
    $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)\bigr{\rVert}\leqslant KW_{p}(\mu,\delta_{0}).$ (3.9)
    When $\mathcal{T}=\mathcal{M}_{\beta}$, this follows immediately from the inequality $\operatorname{e}^{-\beta f^{*}}\leqslant\operatorname{e}^{-\beta f(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})}\leqslant\operatorname{e}^{-\beta f_{*}}$ if $\ell=0$ and from Proposition A.3 if $\ell>0$, where $f^{*}$ the supremum of $f$ over ${}^{d}$. When $\mathcal{T}=\sqrt{\mathcal{C}_{\beta}}$, the inequality (3.9) follows in the same way after noticing that
    $\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}\Bigr{\rVert}_{\rm F}^{2}=\operatorname*{tr}\Bigl{(}\mathcal{C}_{\beta}\left(\mu\right)\Bigr{)}=\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{2}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}-\Bigl{\lvert}\mathcal{M}_{\beta}(\mu)\Bigr{\rvert}^{2}\leqslant\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{2}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}$ (3.10)
    for all $\mu\in\mathcal{P}_{p}(\mathbf{R}^{d})$.
Now, assume for contradiction that there exists $R>0$ and $\{(\mu_{i},\nu_{i})\}_{i\in\mathbf{N}}$ in $\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}$ such that
$\forall i\in\mathbf{N},\qquad\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{W_{p}\left(\mu_{i},\nu_{i}\right)}\geqslant i.$ (3.11)
In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of 3.11 would be bounded from above uniformly in $i$ by (3.8). Therefore $W_{p}(\nu_{i},\delta_{0})\to\infty$. On the other hand, by the triangle inequality,
$W_{p}\left(\mu_{i},\nu_{i}\right)\geqslant W_{p}\left(\nu_{i},\delta_{0}\right)-W_{p}\left(\mu_{i},\delta_{0}\right)\geqslant W_{p}\left(\nu_{i},\delta_{0}\right)-R.$
Therefore, by (3.9), we deduce that
$\limsup_{i\to\infty}\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{W_{p}\left(\mu_{i},\nu_{i}\right)}\leqslant K\limsup_{i\to\infty}\frac{W_{p}(\nu_{i},\delta_{0})+R}{W_{p}\left(\nu_{i},\delta_{0}\right)-R}=K,$
which contradicts (3.11). ∎
3.2 Moment estimates for the empirical measures
Lemma 3.5.
Let $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$, and suppose that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for some $p\geqslant 2$.
  1. (a)
    Consider the particle system (1.1) with $\overline{\rho}_{0}^{\otimes J}$-distributed initial data, and let $\mu^{J}_{t}$ denote the corresponding empirical measures. Then there exists a constant $\kappa>0$ independent of $J$ such that for all $J\in\mathbf{N}^{+}$
    $\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}\right\rvert^{p}\right]\quad\vee\quad\mathbf{E}\left[\sup_{t\in[0,T]}\mu^{J}_{t}\bigl{[}|x|^{p}\bigr{]}\right]\quad\vee\quad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\mu^{J}_{t})\right\rvert^{p}\right]\leqslant\kappa.$
  2. (b)
    Consider the particle system (1.4) with $\overline{\rho}_{0}^{\otimes J}$-distributed initial data, and let $\mu^{J}_{t}$ denote the corresponding empirical measures. Then there exists a constant $\kappa>0$ independent of $J$ such that for all $J\in\mathbf{N}^{+}$
    $\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}\right\rvert^{p}\right]\vee\mathbf{E}\left[\sup_{t\in[0,T]}\mu^{J}_{t}\bigl{[}|x|^{p}\bigr{]}\right]\vee\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\mu^{J}_{t})\right\rvert^{p}\right]\vee\mathbf{E}\left[\sup_{t\in[0,T]}\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu^{J}_{t})}\right\rVert^{p}\right]\leqslant\kappa.$
Proof.
The result can be proved using the same reasoning as in [10, Lemma 3.4] or [23, Lemma 15]. For convenience, we denote by $C$ any constant depending only on $(f,\beta,\lambda,p,T)$. We first observe by (3.10) and Proposition A.3 that
$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\Bigl{\lvert}\mathcal{M}_{\beta}(\mu)\Bigr{\rvert}\vee\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}\Bigr{\rVert}_{\rm F}\leqslant C\left(\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}.$ (3.12)
Writing the drift and diffusion coefficients for (a) and (b) as in (2.10) and (2.17), respectively, it holds in both cases that
$\forall x\in\mathbf{R}^{d},\qquad\forall\mu\in\mathcal{P}(\mathbf{R}^{d}),\qquad\bigl{\lvert}b(x,\mu)\bigr{\rvert}\vee\bigl{\lVert}\sigma(x,\mu)\bigr{\rVert}_{\rm F}\leqslant C\left(\lvert x\rvert^{p}+\int_{\mathbf{R}^{d}}|y|^{p}\,\mu(\mathrm{d}y)\right)^{\frac{1}{p}}.$
Arguing in parallel for (a) and (b), we fix $j\in$ and apply the Burkholder–Davis–Gundy inequality to find for all $t\in[0,T]$ that
$\frac{1}{3^{p-1}}\mathbf{E}\left[\sup_{s\in[0,t]}\left\lvert X_{s}^{(j)}\right\rvert^{p}\right]$ $\leqslant\mathbf{E}\left\lvert X_{0}^{(j)}\right\rvert^{p}+T^{p-1}\int_{0}^{t}\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad\qquad+C_{\rm BDG}T^{\frac{p}{2}-1}\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$
$\leqslant\mathbf{E}\left\lvert X_{0}^{(j)}\right\rvert^{p}+C\int_{0}^{t}\mathbf{E}\left[\left\lvert X_{s}^{(j)}\right\rvert^{p}+\int_{\mathbf{R}^{d}}|y|^{p}\,\mu^{J}_{s}(\mathrm{d}y)\right]\,\mathrm{d}s.$
Since the common law of $\left(X^{(J)},\dots,X^{(J)}\right)$ is invariant under permutations of $\llbracket 1,J\rrbracket$, its marginals corresponding to the particles are all the same. Therefore, given that $\mu^{J}_{s}=\frac{1}{J}\sum_{j=1}^{J}\delta_{X^{(j)}_{s}}$, we find that
$\mathbf{E}\left[\int_{\mathbf{R}^{d}}|y|^{p}\,\mu^{J}_{s}(\mathrm{d}y)\right]=\mathbf{E}\left\lvert X^{(j)}_{s}\right\rvert^{p}.$
Consequently, it holds that
$\mathbf{E}\left[\sup_{s\in[0,t]}\lvert X_{s}^{(j)}\rvert^{p}\right]$ $\leqslant C\mathbf{E}\left\lvert X_{0}^{(j)}\right\rvert^{p}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X_{u}^{(j)}\right\rvert^{p}\right]\,\mathrm{d}s.$
Grönwall’s inequality then gives that
$\mathbf{E}\left[\sup_{t\in[0,T]}\bigl{\lvert}X_{t}^{(j)}\bigr{\rvert}^{p}\right]\leqslant\kappa.$
The second bound in (a) and in (b) follows from the inequality
$\mathbf{E}\left[\sup_{t\in[0,T]}\mu^{J}_{t}\bigl{[}\left\lvert x\right\rvert^{p}\bigr{]}\right]=\mathbf{E}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\bigl{\lvert}X^{(j)}_{t}\bigr{\rvert}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\bigl{\lvert}X^{(j)}_{t}\bigr{\rvert}^{p}\right]=\mathbf{E}\left[\sup_{t\in[0,T]}\bigl{\lvert}X^{(j)}_{t}\bigr{\rvert}^{p}\right],$
while the third bound in (a) as well as the third and fourth bound in (b) follow from (3.12). ∎
3.3 Convergence of weighted moments for i.i.d. samples
For the proof of the lemmas in this section, we will need [18, Theorem 1], which we recall for the reader’s convenience in Theorem 3.6 below, slightly adapted to the i.i.d. setting. We also refer to [1, Theorem 2.3], where similar results for $p=1$ and $p=2$ are proved in the context of importance sampling, with explicit constant prefactors.
Theorem 3.6 ($L^{p}$ convergence of weighted sums).
Let $(w_{j},V_{j})_{j\in\mathbf{N}}$ be i.i.d. random variables with values in $\mathbf{R}\times\mathbf{R}^{d}$ or $\mathbf{R}\times\mathbf{R}^{d\times d}$ with $w_{j}>0$ almost surely, and set
$\widehat{N}_{J}=\frac{1}{J}\sum_{j=1}^{J}w_{j}V_{j},\qquad\widehat{D}_{J}=\frac{1}{J}\sum_{j=1}^{J}w_{j},\qquad\widehat{R}_{J}=\frac{\widehat{N}_{J}}{\widehat{D}_{J}}.$
Let also $N=\mathbf{E}\left[\widehat{N}_{J}\right]$, $D=\mathbf{E}\left[\widehat{D}_{J}\right]$, and $R=N/D$. Let $0<p<q$ and assume that for some $c>0$,
$\left\lVert w_{1}\right\rVert_{L^{q}(\Omega)}\leqslant c,\qquad\left\lVert V_{1}\right\rVert_{L^{r}(\Omega)}\leqslant c,\qquad\left\lVert w_{1}V_{1}\right\rVert_{L^{s}(\Omega)}\leqslant c,\qquad r:=\frac{p(q+2)}{q-p},\qquad s:=\frac{pq}{q-p}.$
Then there is $C$ depending only on $(c,p,q,r,s)$ such that
$\bigl{\lVert}\widehat{R}_{J}-R\bigr{\rVert}_{{L^{p}(\Omega)}}\leqslant CJ^{-\frac{1}{2}}.$
Using Theorem 3.6, we now prove the convergence of the weighted mean and weighted covariance for empirical measures formed from i.i.d. samples, in the limit of infinitely many samples.
Lemma 3.7 (Convergence of the weighted mean for i.i.d. samples).
Suppose that $f\colon^{d}\to$ is bounded from below and let $0<p<r$. Then for all $\mu\in\mathcal{P}_{r}(\mathbf{R}^{d})$, there is $C$ depending only on $(f,\beta,p,r)$ and the $r$-th moment of $\mu$ such that for all $J\in\mathbf{N}$
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}\right)-\mathcal{M}_{\beta}\bigl{(}\mu\bigr{)}\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
Proof of Lemma 3.7.
We apply Theorem 3.6 with $q=\frac{p(r+2)}{r-p}$, in which case $\frac{p(q+2)}{q-p}=r.$ In our setting
$(w_{j},V_{j})=\left(\operatorname{e}^{-\beta f\left(\overline{X}^{(j)}\right)},\overline{X}^{(j)}\operatorname{e}^{-\beta f\left(\overline{X}^{(j)}\right)}\right).$
Since $w_{j}$ is bounded from above almost surely, it obviously holds that $\left\lVert w_{1}\right\rVert_{L^{q}(\Omega)}<\infty$. Since $\mu\in\mathcal{P}_{r}(^{d})$, it also holds that $\left\lVert V_{1}\right\rVert_{L^{r}(\Omega)}<\infty$ and $\left\lVert w_{1}V_{1}\right\rVert_{L^{s}(\Omega)}<\infty$ since $s:=\frac{pq}{q-p}<r$. Consequently, the assumptions of Theorem 3.6 are satisfied, and we obtain the statement.
Lemma 3.8 (Convergence of the weighted covariance for i.i.d. samples).
Suppose that $f\colon^{d}\to$ is bounded from below and let $0<2p<r$. Then for all $\mu\in\mathcal{P}_{r}(\mathbf{R}^{d})$ satisfying $\mathcal{C}_{\beta}(\mu)\succcurlyeq\eta\operatorname{I}_{d}\succ 0$, there is $C$ depending only on $(f,\beta,p,r,\eta)$ and the $r$-th moment of $\mu$ such that for all $J\in\mathbf{N}^{+}$
$\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert_{\rm F}^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{X^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
Proof.
By assumption it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$. By an inequality due to van Hemmen and Ando [50], see also [5, Problem X.5.5], it holds that
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert_{\rm F}\leqslant\frac{1}{\eta}\Bigl{\lVert}\mathcal{C}_{\beta}(\overline{\mu}^{J})-\mathcal{C}_{\beta}(\mu)\Bigr{\rVert}_{\rm F}.$
Using the triangle inequality, we have that
$\Bigl{\lVert}\mathcal{C}_{\beta}(\overline{\mu}^{J})-\mathcal{C}_{\beta}(\mu)\Bigr{\rVert}_{\rm F}^{p}$ $=\Bigl{\lVert}L_{\beta}\overline{\mu}^{J}\left[x\otimes x\right]-L_{\beta}\mu\left[x\otimes x\right]-\mathcal{M}_{\beta}(\overline{\mu}^{J})\otimes\mathcal{M}_{\beta}(\overline{\mu}^{J})+\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu)\Bigr{\rVert}_{\rm F}^{p}$
$\leqslant 2^{p-1}\Bigl{\lVert}L_{\beta}\overline{\mu}^{J}\left[x\otimes x\right]-L_{\beta}\mu\left[x\otimes x\right]\Bigr{\rVert}_{\rm F}^{p}$
$\qquad\qquad+2^{p-1}\Bigl{\lVert}\mathcal{M}_{\beta}(\overline{\mu}^{J})\otimes\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu)\Bigr{\rVert}_{\rm F}^{p}.$
Convergence to zero of the expectation of the first term with rate $J^{-\frac{p}{2}}$ follows from Theorem 3.6, this time with $q=\frac{p(r+4)}{r-2p}$, in which case $\frac{p(q+2)}{q-p}=\frac{r}{2}.$ For the second term, we use (3.4) to obtain that
$\mathbf{E}\Bigl{\lVert}\mathcal{M}_{\beta}(\overline{\mu}^{J})\otimes\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu)\Bigr{\rVert}_{\rm F}^{p}$
$\qquad\leqslant\mathbf{E}\left[\bigl{(}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})\right\rvert+\lvert\mathcal{M}_{\beta}(\mu)\rvert\bigr{)}^{p}\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\rvert^{p}\right]$
$\qquad\leqslant\left(\mathbf{E}\left[\bigl{(}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})\right\rvert+\lvert\mathcal{M}_{\beta}(\mu)\rvert\bigr{)}^{2p}\right]\mathbf{E}\left[\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\rvert^{2p}\right]\right)^{\frac{1}{2}}$
$\qquad\leqslant CJ^{-\frac{p}{2}}\left(\mathbf{E}\left[\bigl{(}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})\right\rvert+\lvert\mathcal{M}_{\beta}(\mu)\rvert\bigr{)}^{2p}\right]\right)^{\frac{1}{2}},$
where we used Lemma 3.7. Since $\mathcal{M}_{\beta}(\overline{\mu}^{J})$ converges to $\mathcal{M}_{\beta}(\mu)$ in $L^{2p}(\Omega)$ again by Lemma 3.7, this inequality enables to conclude the proof. ∎
3.4 No collapse in finite time for CBS
In order to deal with the matrix square root in the CBS dynamics, we prove that the weighted covariance stays positive definite for finite times.
Proposition 3.9 (No collapse in finite time).
Let the cost function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ be such that the macroscopic McKean SDE for CBS (1.5) has a solution $(\overline{X}_{t})_{t\in[0,T]}$ over some finite time interval $[0,T]$. Then
$\forall t\in[0,T],\qquad\mathcal{C}(\overline{\rho}_{t})\succcurlyeq\mathcal{C}(\overline{\rho}_{0})\operatorname{e}^{-2t}.$ (3.13)
In particular, if $\mathcal{C}(\overline{\rho}_{0})\succ 0$ and if $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ is continuous, then
$\inf_{t\in[0,T]}\lambda_{\rm min}\bigl{(}\mathcal{C}(\overline{\rho}_{t})\bigr{)}>0$ and $\inf_{t\in[0,T]}\lambda_{\rm min}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$ (3.14)
Proof.
The evolution of the covariance matrix is governed by [11, Equation (2.18)]
$\partial_{t}\left(\mathcal{C}(\overline{\rho}_{t})\right)=-2\mathcal{C}(\overline{\rho}_{t})+2\lambda^{-1}\mathcal{C}_{\beta}(\overline{\rho}_{t}).$
For a given $y\in\mathbf{R}^{d}$ such that $|y|=1$, define $g(t):=y^{\mathsf{T}}\mathcal{C}(\overline{\rho}_{t})y$ for $t\in[0,T]$. Then
$g^{\prime}(t)=-2g(t)+2\lambda^{-1}y^{T}\mathcal{C}_{\beta}(\overline{\rho}_{t})y\geqslant-2g(t),$
since $\mathcal{C}_{\beta}(\overline{\rho}_{t})$ is positive semi-definite. Applying Grönwall’s inequality to $-g$, we deduce that $g(t)\geqslant g(0)\operatorname{e}^{-2t}.$ This shows (3.13) and the first inequality in (3.14). In order to prove the second inequality in (3.14), assume that it was wrong. By continuity of $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$, we have $\lambda_{\rm min}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}=0$ for some $t\in[0,T]$. Then $y^{\mathsf{T}}\mathcal{C}_{\beta}(\overline{\rho}_{t})y=0$ for some $y\in\mathbf{R}^{d}$ such that $|y|=1$. In particular
$\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\operatorname{e}^{-\beta f(x)}\,\overline{\rho}_{t}(\mathrm{d}x)=0.$
Thus we obtain $y^{\mathsf{T}}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\bigr{)}=0$ for $\overline{\rho}_{t}$-almost every $x\in\mathbf{R}^{d}$, and so $y^{\mathsf{T}}\mathcal{M}_{\beta}(\overline{\rho}_{t})=y^{\mathsf{T}}\mathcal{M}(\overline{\rho}_{t})$ by integrating against $\overline{\rho}_{t}$. But then,
$y^{\mathsf{T}}\mathcal{C}(\overline{\rho}_{t})y=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=0.$
which is a contradiction. Thus $C_{\beta}(\overline{\rho}_{t})\succ 0$ for all $t\in[0,T]$, which implies the second inequality in (3.14). ∎
4 Proof of the well-posedness results
4.1 Well-posedness for the microscopic models
For a collection of particles $\mathbf{X}\in\mathbf{R}^{dJ}$, the notation $\lvert\mathbf{X}\rvert$ denotes the usual Euclidean norm; that is
$\lvert\mathbf{X}\rvert:=\sqrt{\sum_{j=1}^{J}\left\lvert X^{(j)}\right\rvert^{2}}.$
For the proof of Theorem 2.2 we need the following stability result, which only requires local Lipschitz continuity of $f$, in contrast to the results in Subsection 3.1.
Lemma 4.1.
For all $\mathbf{X}\in\mathbf{R}^{dJ}$ it holds that
[…]
Moreover, if $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ is locally Lipschitz continuous, then for each $R>0$ and $J\in\mathbf{N}^{+}$ there exists a constant $C_{J,R}>0$ such that
$\lvert\mathcal{M}_{\beta}(\mathbf{X})-\mathcal{M}_{\beta}(\mathbf{Y})\rvert\leqslant C_{J,R}\lvert\mathbf{X}-\mathbf{Y}\rvert$ and $\left\lVert\sqrt{\mathcal{C}_{\beta}(\mathbf{X})}-\sqrt{\mathcal{C}_{\beta}(\mathbf{Y})}\right\rVert_{F}\leqslant C_{J,R}\lvert\mathbf{X}-\mathbf{Y}\rvert$ (4.2)
for all $\mathbf{X},\mathbf{Y}\in\mathbf{R}^{dJ}$ satisfying $\lvert\textbf{X}\rvert,|\mathbf{Y}|\leqslant R$.
Proof.
The bound on $\mathcal{M}_{\beta}(\mathbf{X})$ in (4.1) follows immediately from the definition: We have that
$\Bigl{\lvert}\mathcal{M}_{\beta}(\mathbf{X})\Bigr{\rvert}^{2}=\left\lvert\sum_{j=1}^{J}w_{j}(\mathbf{X})X^{(j)}\right\rvert^{2}\leqslant\sum_{j=1}^{J}w_{j}(\mathbf{X})\left\lvert X^{(j)}\right\rvert^{2}\leqslant\sum_{j=1}^{J}\left\lvert X^{(j)}\right\rvert^{2}=\left\lvert\mathbf{X}\right\rvert^{2},$
where we used convexity of the square function and the fact that the weights, given hereafter, are bounded from above by 1:
$w_{j}(\mathbf{X}):=\frac{\operatorname{e}^{-\beta f\left(X^{(j)}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}\right)}}.$
Similarly, for the weighted covariance, it holds by the triangle inequality that
$\left\lVert\mathcal{C}_{\beta}(\mathbf{X})\right\rVert_{\rm F}$ $=\left\lVert\sum_{j=1}^{J}w_{j}(\mathbf{X})\Bigl{(}X^{(j)}\otimes X^{(j)}\Bigr{)}-\mathcal{M}_{\beta}(\mathbf{X})\otimes\mathcal{M}_{\beta}(\mathbf{X})\right\rVert_{\rm F}$
$\leqslant\left\lVert\sum_{j=1}^{J}w_{j}(\mathbf{X})\Bigl{(}X^{(j)}\otimes X^{(j)}\Bigr{)}\right\rVert_{\rm F}\leqslant\sum_{j=1}^{J}w_{j}(\mathbf{X})\left\lVert X^{(j)}\otimes X^{(j)}\right\rVert_{\rm F}=\sum_{j=1}^{J}w_{j}(\mathbf{X})\left\lvert X^{(j)}\right\rvert^{2}\leqslant\left\lvert\mathbf{X}\right\rvert^{2}.$
The local Lipschitz estimate on $\mathcal{M}_{\beta}(\mathbf{X})$ in (4.2) is [10, Lemma 2.1]. In order to prove the local Lipschitz continuity of $\mathcal{C}_{\beta}(\mathbf{X})$, fix $R>0$ and $J>0$. As in the proof of Proposition 3.2 we obtain
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mathbf{X})}-\sqrt{\mathcal{C}_{\beta}(\mathbf{Y})}\right\rVert_{\rm F}$ $\leqslant\sqrt{2}\left(\sum_{j=1}^{J}\left\lvert\sqrt{w_{j}(\mathbf{X})}\left(X^{(j)}-\mathcal{M}_{\beta}(\mathbf{X})\right)-\sqrt{w_{j}(\mathbf{Y})}\left(Y^{(j)}-\mathcal{M}_{\beta}(\mathbf{Y})\right)\right\rvert^{2}\right)^{\frac{1}{2}}$
$\leqslant\sqrt{2}\sum_{j=1}^{J}\left\lvert\sqrt{w_{j}(\mathbf{X})}\left(X^{(j)}-\mathcal{M}_{\beta}(\mathbf{X})\right)-\sqrt{w_{j}(\mathbf{Y})}\left(Y^{(j)}-\mathcal{M}_{\beta}(\mathbf{Y})\right)\right\rvert.$
Using the triangle inequality, we have that
$\frac{1}{\sqrt{2}}\left\lVert\sqrt{\mathcal{C}_{\beta}(\mathbf{X})}-\sqrt{\mathcal{C}_{\beta}(\mathbf{X})}\right\rVert_{\rm F}$ $\leqslant\sum_{j=1}^{J}\sqrt{w_{j}(\mathbf{X})}\left\lvert X^{(j)}-\mathcal{M}_{\beta}(\mathbf{X})-Y^{(j)}+\mathcal{M}_{\beta}(\mathbf{Y})\right\rvert$
$\qquad+\sum_{j=1}^{J}\left\lvert\left(\sqrt{w_{j}(\mathbf{X})}-\sqrt{w_{j}(\mathbf{Y})}\right)\left(Y^{(j)}-\mathcal{M}_{\beta}(\mathbf{Y})\right)\right\rvert.$
Since $\mathbf{X}\mapsto\mathcal{M}_{\beta}(\mathbf{X})$ is a locally Lipschitz continuous function by [10, Lemma 2.1], it suffices to prove that $\mathbf{X}\mapsto\sqrt{w_{j}(\mathbf{X})}$ is locally Lipschitz as well. To this end, observe that
$\sqrt{w_{j}(\mathbf{X})}-\sqrt{w_{j}(\mathbf{Y})}=\frac{w_{j}(\mathbf{X})-w_{j}(\mathbf{Y})}{\sqrt{w_{j}(\mathbf{X})}+\sqrt{w_{j}(\mathbf{Y})}}.$
The denominator can be controlled by noting that
$\forall\lvert\mathbf{X}\rvert\leqslant R,\qquad w_{j}(\mathbf{X})=\frac{\operatorname{e}^{-\beta f\left(X^{(j)}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}\right)}}\geqslant\frac{1}{J}\exp\left(-\beta\left(\max_{\lvert x\rvert\leqslant R}f(x)-\min_{\lvert x\rvert\leqslant R}f(x)\right)\right).$
Therefore, there exists a constant $c_{R,J}>0$ such that $\sqrt{w_{j}(\mathbf{X})}\geqslant c_{R,J}$ for all $\lvert\mathbf{X}\rvert\leqslant R$. Since $\mathbf{X}\mapsto{w_{j}(\mathbf{X})}$ is locally Lipschitz continuous as a composition of locally Lipschitz continuous functions, the claim follows. ∎
Proof of Theorem 2.2.
We prove only the claim for the CBS dynamics (1.4), since the proof for CBO is similar. We write the system (1.4) in the form
$\mathrm{d}\mathbf{X}_{t}=F(\mathbf{X}_{t})\,\mathrm{d}t+G(\mathbf{X}_{t})\,\mathrm{d}\mathbf{W}_{t}$
where $\mathbf{X}_{t}:=\Bigl{(}X^{(1)}_{t},\dots,X^{(J)}_{t}\Bigr{)}$. The drift and diffusion coefficients are given by
$F(\mathbf{X}_{t}):=-\begin{pmatrix}X^{(1)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\\ \vdots\\ X^{(J)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\end{pmatrix}\in\mathbf{R}^{dJ},$
and
$G(\mathbf{X}_{t}):=$ $\sqrt{2\lambda^{-1}}\operatorname{diag}\left(\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)},\dots,\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\right)\in\mathbf{R}^{dJ\times dJ}.$
By Lemma 4.1, the functions $F$ and $G$ have sublinear growth and are locally Lipschitz continuous. Thus, by the non-explosion criterion from stochastic Lyapunov theory [38, Theorem 3.5], it suffices to find a function $\phi\in C^{2}\left(\mathbf{R}^{dJ},[0,\infty)\right)$ such that
  • The function $\phi$ is coercive: $\phi(\mathbf{X})\to\infty$ as $|\mathbf{X}|\to\infty$ and
  • There exists $c>0$ such that
    $\forall\mathbf{X}\in\mathbf{R}^{dJ},\qquad\mathcal{L}\phi(\mathbf{X}):=F(\mathbf{X})\cdot\nabla\phi(\mathbf{X})+\frac{1}{2}\operatorname*{tr}\left(G^{\mathsf{T}}(\mathbf{X})\operatorname{D}^{2}\phi(\mathbf{X})G(X)\right)\leqslant c\phi(\mathbf{X}).$
Take $\phi(\mathbf{X}):=\tfrac{1}{2}|\mathbf{X}|^{2}$. Then
$\mathcal{L}\phi(\mathbf{X})$ $=-\sum_{j=1}^{J}\left(X^{(j)}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\cdot X^{(j)}+J\lambda^{-1}\operatorname*{tr}\Bigl{(}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}$
$=-\sum_{j=1}^{J}\left(X^{(j)}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\cdot X^{(j)}+\lambda^{-1}\sum_{j=1}^{J}w_{j}(\mathbf{X})\left\lvert X^{(j)}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right\rvert^{2}.$
From (4.1) it follows that this is a valid choice, concluding the proof. ∎
4.2 Well-posedness for the mean-field SDEs
We adapt the argument from [10], using the more general Assumption H and the improved stability estimates from Subsection 3.1. We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3. In the following, we endow the vector space $\mathcal{X}:=\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
$\left\lVert(m,\Gamma)\right\rVert:=\left({|m|^{2}+\left\lVert\Gamma\right\rVert^{2}_{\rm F}}\right)^{\frac{1}{2}}.$
The Hölder space $C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}$ is given by
$C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}:=\left\{\varphi\colon[0,T]\rightarrow\mathcal{X}:\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}<\infty\right\},$
where
$\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}:=\sup_{t\in[0,T]}\bigl{\lVert}\varphi(t)\bigr{\rVert}+[\varphi]_{C^{0,\frac{1}{2}}}<\infty\qquad\text{and}\qquad[\varphi]_{C^{0,\frac{1}{2}}}:=\sup_{0\leqslant s<t\leqslant T}\frac{\bigl{\lVert}\varphi(t)-\varphi(s)\bigr{\rVert}}{|t-s|^{\frac{1}{2}}}.$
In order to ensure that certain matrices are positive semidefinite, we use the map $\mathcal{R}\colon\mathbf{R}^{d\times d}\rightarrow\mathbf{R}^{d\times d}$ defined by $\mathcal{R}(\Gamma)=\sqrt{\Gamma\Gamma^{\mathsf{T}}}$. We will make use of the following auxiliary lemma.
Lemma 4.2.
For $f$ and $\overline{\rho}_{0}$ as in Theorem 2.4, fix $(m,\Gamma)\in C([0,T],\mathcal{X})$ and $y_{0}\sim\overline{\rho}_{0}$. Then the stochastic differential equation
$\mathrm{d}Y_{t}=-(Y_{t}-m_{t})\,\mathrm{d}t+\sqrt{2\lambda^{-1}\mathcal{R}(\Gamma_{t})}\,\mathrm{d}W_{t},\qquad Y_{0}=y_{0}\,,$ (4.3)
admits a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$. Moreover, denoting by $\rho_{t}:=\operatorname{Law}(Y_{t})\in\mathcal{P}(\mathbf{R}^{d})$ the pointwise law of $Y_{t}$, the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ belong to $C([0,T],\mathbf{R}^{d})$ and $C([0,T],\mathbf{R}^{d\times d})$ respectively.
Proof.
The existence of a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$ to (4.3) follows from classical theory of stochastic differential equations [45, Theorem 5.2.1]. It remains to prove the continuity statement for the weighted mean and covariance. For any $0\leqslant r\leqslant t\leqslant T$, it holds by the Burkholder–Davis–Gundy inequality that
$\mathbf{E}\left[\sup_{u\in[r,t]}\left\lvert Y_{u}-Y_{r}\right\rvert^{p}\right]$ $\leqslant\bigl{(}2(t-r)\bigr{)}^{p-1}\int_{r}^{t}\mathbf{E}\bigl{\lvert}Y_{u}-m_{u}\bigr{\rvert}^{p}\,\mathrm{d}u$
[…]
Taking $r=0$ in this inequality and rearranging, we obtain that
$\mathbf{E}\left[\sup_{u\in[0,t]}\left\lvert Y_{u}\right\rvert^{p}\right]\leqslant C\left(\mathbf{E}\left\lvert Y_{0}\right\rvert^{p}+\lVert m\rVert_{L^{\infty}([0,T])}^{p}+\lVert\Gamma\rVert_{L^{\infty}([0,T])}^{\frac{p}{2}}+\int_{0}^{t}\mathbf{E}\lvert Y_{u}\rvert^{p}\,\mathrm{d}u\right),$
for a constant $C$ depending only on $(T,\lambda,p)$. It follows from this inequality, by Grönwall’s lemma, that
$\mathbf{E}\left[\sup_{t\in[0,T]}|Y_{t}|^{p}\right]\leqslant C\left(\mathbf{E}\left\lvert Y_{0}\right\rvert^{p}+\lVert m\rVert_{L^{\infty}([0,T])}^{p}+\lVert\Gamma\rVert_{L^{\infty}([0,T])}^{\frac{p}{2}}\right)\operatorname{e}^{CT}<\infty.$ (4.5)
Therefore, by dominated convergence, we have for any $t\in[0,T]$ that
$\mathcal{M}_{\beta}(\rho_{r})=\frac{\mathbf{E}\left[Y_{r}\operatorname{e}^{-\beta f(Y_{r})}\right]}{\mathbf{E}\left[\operatorname{e}^{-\beta f(Y_{r})}\right]}\xrightarrow[r\to t]{}\frac{\mathbf{E}\left[Y_{t}\operatorname{e}^{-\beta f(Y_{t})}\right]}{\mathbf{E}\left[\operatorname{e}^{-\beta f(Y_{t})}\right]}=\mathcal{M}_{\beta}(\rho_{t}),$
implying that the function $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ belongs to $C([0,T],\mathbf{R}^{d})$. An analogous reasoning shows that the function $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ belongs to $C([0,T],\mathbf{R}^{d\times d})$. ∎
Proof of Theorem 2.4.
Similarly to the proof of [10, Theorem 3.1], we first construct a map
$\mathcal{T}\colon C\left([0,T],\mathcal{X}\right)\rightarrow C\left([0,T],\mathcal{X}\right)$
whose fixed points correspond to solutions of (1.5).
Step 1: Constructing the map $\mathcal{T}$
Consider the map
$\mathcal{T}\colon C([0,T],\mathcal{X})$ $\rightarrow C([0,T],\mathcal{X})$
$(m,\Gamma)$ $\mapsto\bigl{(}\mathcal{M}_{\beta}(\rho_{t}),\mathcal{C}_{\beta}(\rho_{t})\bigr{)},$
where $\rho_{t}:={\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3). Thanks to Lemma 4.2, the map $\mathcal{T}$ is well-defined. Fixed points of $\mathcal{T}$ correspond to solutions of the McKean-Vlasov SDE (1.5). The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [27, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
$\Bigl{\{}(m,\Gamma)\in C\bigl{(}[0,T],\mathcal{X}\bigr{)}:\exists\xi\in[0,1]\text{ such that }(m,\Gamma)=\xi\mathcal{T}\bigl{(}(m,\Gamma)\bigr{)}\Bigr{\}}.$ (4.6)
Step 2: Showing that $\mathcal{T}$ is compact
In order to prove that $\mathcal{T}$ is a compact operator, fix $R>0$ and consider the ball
$B_{R}:=\left\{(m,\Gamma)\in C([0,T],\mathcal{X}):\left\lVert(m,\Gamma)\right\rVert_{L^{\infty}([0,T])}\leqslant R\right\}.$
By the Arzelà–Ascoli theorem, we have a compact embedding
$C^{0,\frac{1}{2}}\left([0,T],\mathcal{X}\right)\hookrightarrow C\left([0,T],\mathcal{X}\right).$
Thus, it suffices to show that $\mathcal{T}(B_{R})$ is bounded in $C^{0,\frac{1}{2}}\left([0,T],\mathcal{X}\right)$. To this end, take $(m,\Gamma)\in B_{R}$ and consider the corresponding solution $Y_{t}$ of (4.3) with pointwise law $\rho_{t}$, and let $\varphi=\mathcal{T}(m,\Gamma)$. We first observe that
$\bigl{\lVert}\varphi(0)\bigr{\rVert}^{2}=\bigl{\lvert}\mathcal{M}_{\beta}(\overline{\rho}_{0})\bigr{\rvert}^{2}+\bigl{\lVert}\mathcal{C}_{\beta}(\overline{\rho}_{0})\bigr{\rVert}_{\rm F}^{2}<\infty.$
The right-hand side is independent of $m$ and $\Gamma$, and it is indeed finite by Proposition A.3. Given that
$\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}\leqslant\left\lVert\varphi(0)\right\rVert+[\varphi]_{C^{0,\frac{1}{2}}}\bigl{(}\sqrt{T}+1\bigr{)},$
it only remains to show that $[\varphi]_{C^{0,\frac{1}{2}}}$ is bounded by a finite constant depending on $(m,\Gamma)$ only through $R$. In view of (4.4) and (4.5), there is a constant $L=L(R)$ such that
$\mathbf{E}\left\lvert Y_{t}-Y_{r}\right\rvert^{p}\leqslant L\lvert t-r\rvert^{\frac{p}{2}}.$
It follows that $W_{p}(\rho_{t},\rho_{r})\leqslant L^{\frac{1}{p}}\lvert t-r\rvert^{\frac{1}{2}}$, which by Proposition 3.1 implies the Hölder continuity with exponent $\frac{1}{2}$ of the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$. This completes the proof that $\mathcal{T}$ is compact.
Step 3: Showing that the set (4.6) is bounded
To this end, let $(m,\Gamma)\in C([0,T],\mathcal{X})$ be such that
$(m,\Gamma)=\xi\mathcal{T}(m,\Gamma)$ (4.7)
for some $\xi\in[0,1]$, and let $(Y_{t})$ denote the corresponding solution to (4.3). By (4.7), the stochastic process $(Y_{t})$ is also a solution to
$\mathrm{d}Y_{t}=-\Bigl{(}Y_{t}-\xi\mathcal{M}_{\beta}(\rho_{t})\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\xi\mathcal{C}_{\beta}(\rho_{t})}\,\mathrm{d}W_{t},\qquad\rho_{t}:={\rm Law}(Y_{t}).$
But then, reasoning as in Subsection 3.2 using a Grönwall-type argument and Proposition A.3 with $u=\ell$, we find that
$\mathbf{E}\left[\sup_{t\in[0,T]}\bigl{\lvert}Y_{t}\bigr{\rvert}^{p}\right]\leqslant C\mathbf{E}\lvert Y_{0}\rvert^{p}=C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\overline{\rho}_{0}(\mathrm{d}x).$
By Proposition A.3, it follows that $\mathcal{M}_{\beta}(\rho_{t})$ and $\mathcal{C}_{\beta}(\rho_{t})$ can be bounded uniformly in $[0,T]$ in terms of $\overline{\rho}_{0}$, implying that the set (4.6) is indeed bounded. Thus, we established the existence of a fixed point of $\mathcal{T}$, and we obtain as a byproduct the moment bounds in (2.6).
Step 4: Showing uniqueness
For uniqueness, let $(m_{t},\Gamma_{t})$ and $(\widehat{m}_{t},\widehat{\Gamma}_{t})$ be two fixed points of $\mathcal{T}$ with corresponding solutions $Y_{t},\widehat{Y}_{t}\in C([0,T],\mathcal{X})$ of (4.3) and let $z_{t}:=Y_{t}-\widehat{Y}_{t}$. By definition of $\mathcal{T}$ and since $\mathcal{C}_{\beta}(\mu)$ is symmetric and positive semidefinite for all $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d})$, we have $\mathcal{R}\left(\Gamma_{t}\right)=\Gamma_{t}$ and $\mathcal{R}\bigl{(}\widehat{\Gamma}_{t}\bigr{)}=\widehat{\Gamma}_{t}$ for all $t\in[0,T]$. Let $\rho_{t},\widehat{\rho}_{t}\in\mathcal{P}(\mathbf{R}^{d})$ denote the marginal laws of $Y_{t}$ and $\widehat{Y}_{t}$, respectively. By the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$
$\frac{1}{2^{p-1}}\mathbf{E}\left[\sup_{r\in[0,t]}|z_{r}|^{p}\right]$ $\leqslant T^{p-1}\int_{0}^{t}\mathbf{E}\left\lvert Y_{r}-\widehat{Y}_{r}-\mathcal{M}_{\beta}(\rho_{r})+\mathcal{M}_{\beta}(\widehat{\rho}_{r})\right\rvert^{p}\,\mathrm{d}r$
$\qquad+C_{\rm BDG}T^{\frac{p}{2}-1}(2\lambda^{-1})^{\frac{p}{2}}\int_{0}^{t}\left\lVert\sqrt{\mathcal{C}_{\beta}(\rho_{r})}-\sqrt{\mathcal{C}_{\beta}(\widehat{\rho}_{r})}\right\rVert_{\rm F}^{p}\,\mathrm{d}r.$
By Proposition 3.9, we have that
$\eta:=\inf\left\{\lambda_{\rm min}\Bigl{(}\mathcal{C}_{\beta}(\rho_{t})\Bigr{)}:t\in[0,T]\right\}>0.$
By virtue of the van Hemmen–Ando inequality [50, Proposition 3.2], it holds for any unitarily invariant norm, thus in particular for the Frobenius norm, that
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\rho_{r})}-\sqrt{\mathcal{C}_{\beta}(\widehat{\rho}_{r})}\right\rVert_{\rm F}\leqslant\eta^{-\frac{1}{2}}\left\lVert\mathcal{C}_{\beta}(\rho_{r})-\mathcal{C}_{\beta}(\widehat{\rho}_{r})\right\rVert_{\rm F}\,.$
Using Proposition 3.1 and noting that $W_{p}(\rho_{r},\widehat{\rho}_{r})^{p}\leqslant\mathbf{E}\lvert z_{r}\rvert^{p}$ by definition of the Wasserstein distance, we conclude that
$\mathbf{E}\left[\sup_{r\in[0,t]}\bigl{\lvert}z_{r}\bigr{\rvert}^{p}\right]\leqslant C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,r]}\bigl{\lvert}z_{u}\bigr{\rvert}^{p}\right]\,\mathrm{d}r.$
By Grönwall’s lemma, and given that $z_{0}=0$, it follows that
$\mathbf{E}\left[\sup_{t\in[0,T]}\lvert z_{t}\rvert^{p}\right]=0,$
which concludes the proof of uniqueness.
Step 5: Showing that $\overline{\rho}_{t}$ solves the nonlocal Fokker–Planck equation (2.7)
This part of the proof is standard, and we present only a formal derivation. Let $(m,\Gamma)\in C([0,T],\mathcal{X})$ denote the fixed point of $\mathcal{T}$, in which case $\mathcal{R}(\Gamma_{t})=\Gamma_{t}$, and let $\overline{Y}_{t}$ denote the corresponding solution to (4.3). Denote also by $\overline{\rho}_{t}$ the law of $\overline{Y}_{t}$. For a smooth function $\phi\colon\mathbf{R}^{d}\to\mathbf{R}$ with compact support, it holds by Itô’s lemma that
$\mathbf{E}\phi\left(\overline{Y}_{t}\right)-\mathbf{E}\phi\left(\overline{Y}_{r}\right)=\mathbf{E}\left[\int_{r}^{t}\mathcal{L}_{u}\phi\left(\overline{Y}_{u}\right)\,\mathrm{d}u\right],$ (4.8)
where $\mathcal{L}_{u}\phi(y):=-\bigl{(}y-\mathcal{M}_{\beta}\left(\overline{\rho}_{u}\right)\bigr{)}\cdot\nabla\phi(y)+\lambda^{-1}\mathcal{C}_{\beta}\left(\overline{\rho}_{u}\right):\operatorname{D}^{2}\phi(y).$ It follows from (4.8) that $\overline{\rho}_{t}$ satisfies, for all smooth compactly supported $\phi\colon\mathbf{R}^{d}\to\mathbf{R}$,
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbf{R}^{d}}\phi(x)\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\mathcal{L}_{t}\phi\,\overline{\rho}_{t}(\mathrm{d}x),$ (4.9)
which is indeed a weak formulation of the Fokker–Planck equation (2.5). For rigorous uniqueness and existence results for the Fokker–Planck equation (4.9) as well as a more precise discussion of its connection with (4.3), see for example [25]. ∎
Remark 4.3 (Proof of Theorem 2.3).
Observe that the drift coefficients of CBO and CBS are identical and that the diffusion coefficient of CBO is a globally Lipschitz continuous function of its drift coefficient. By, among other things, replacing $\sqrt{2\lambda^{-1}\mathcal{R}(\Gamma_{t})}$ in (4.3) by $S\left(\overline{X}_{t}-\mathcal{M}_{\beta}(\overline{\rho}_{t})\right)$, the above proof can be easily adapted to provide a proof for Theorem 2.3. See also [10], where a weaker version of Theorem 2.3 was already proven in the case where $p=4$ and $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$.
Remark 4.4.
Note that in case of $p\geqslant 2p_{\mathcal{M}}$, we can drop the requirement of $\mathcal{C}(\overline{\rho}_{0})$ to be positive definite by using Proposition 3.2 instead of Proposition 3.1.
5 Extensions and perspectives for future work
This section is organized as follows. First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.12. Then, in the following subsections, we discuss three natural extensions of our results. In Subsection 5.2, we show how our mean-field limit result for CBS (Theorem 2.12) can be improved when the initial condition $\overline{\rho}_{0}$ is Gaussian. In Subsection 5.3, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid. Finally, in Subsection 5.4, we show how Theorems 2.6 and 2.12 can be generalized to smaller values of $p\geqslant 2$.
5.1 Proving mean-field results using stopping times
In this section, we present an alternative approach for proving Theorem 2.6 and Theorem 2.12. This approach is based on stopping times, and inspired by a method employed in [33] to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method. A similar approach was employed in [36] to obtain a non-quantitative mean-field estimate. Before presenting the proof, we briefly discuss the main differences between this approach and the indicator set approach presented in Section 2.
  • The main advantage of the approach used in Section 2 is that the event $\Omega_{J,t}$ for which our stability estimates do not apply can be time-dependent. For the proof of Theorem 2.6, this flexibility is in fact not necessary, as we could have employed instead of $\Omega_{J,t}$ in (2.15) the event
    $\Omega_{J}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\}.$ (5.1)
    In certain contexts, however, such as the one discussed in Subsection 5.2, being able to exclude a time-dependent event is very useful.
  • The stopping time approach presented hereafter does not enjoy this flexibility, but is convenient in settings where moments of both $\mu^{J}$ and $\overline{\mu}^{J}$ need to be controlled in order to apply the stability estimates obtained in Proposition 3.1 for the weighted moments. This is illustrated in Subsections 5.3 and 5.4.
We illustrate the stopping times approach only for Theorem 2.12, noting that Theorem 2.6 can be proved in exactly the same manner.
Alternative proof of Theorem 2.12.
Let $R>0$ be such that the stopping time
$\theta_{J}(R):=\inf\Bigl{\{}t\geqslant 0:\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R\Bigr{\}},$ (5.2)
satisfies $\mathbf{P}[\theta_{J}\leqslant T]=\mathcal{O}(J^{-2}).$ Note that the event $\{\theta_{J}\leqslant T\}$ coincides with the event $\Omega_{J}$ given in (5.1). From Hölder’s inequality it follows that
$\frac{1}{2^{p-1}}\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]$
$\qquad\qquad=\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\mathsf{1}_{\theta_{J}(R)>T}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\mathsf{1}_{\theta_{J}(R)\leqslant T}\right]$
$\qquad\qquad\leqslant\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\Big{\rvert}^{p}\right]+\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{q}\right]\right)^{\frac{p}{q}}\Bigl{(}\mathbf{P}[\theta_{J}(R)\leqslant T]\Bigr{)}^{\frac{q-p}{q}}.$ (5.3)
Bounding the first term in (5.3)
We again employ an adapted version of Sznitman’s argument. We have
$\left\lvert X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\right\rvert^{p}\leqslant$ $\left\lvert\int_{0}^{t\wedge\theta_{J}}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
$+\left\lvert\int_{0}^{t\wedge\theta_{J}}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$
Let us introduce the martingale
$M_{t}=\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}.$
By Doob’s optional stopping theorem [33, Theorem 3.3], see also [20, Equation 2.29, p.285], the stopped process $(M_{t\wedge\theta_{J}})_{t\geqslant 0}$ is a martingale, with a quadratic variation process given by the stopped quadratic variation process of $M_{t}$, i.e. by $(\langle M\rangle_{t\wedge\theta_{J}})_{t\geqslant 0}$. Therefore, by the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$ that
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad+C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.4)
From the triangle inequality we obtain that
$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lvert b\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-b\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\mathbf{E}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s.$ (5.5)
Similarly, for the diffusion term, we have
$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.6)
Next, we need to bound the terms on the right-hand side of (5.5) and (5.6). Since the drift terms and the diffusion terms are bounded in a similar manner, we only present the bounds for the diffusion terms.
  • For the first diffusion term, we have that
    $\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}$ $=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\biggl{\lVert}\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{s\wedge\theta_{J}}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)}\biggr{\rVert}_{\rm F}^{p}.$
    Using Corollary 3.3, together with the fact
    $\mathbf{E}\left[W_{p}\left(\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\right\rvert^{p},$
    we deduce that
    $\mathbf{E}\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p}.$
  • For the other diffusion term, we use Lemma 3.8. The assumptions of this lemma are satisfied since, on the one hand, $\overline{\rho}_{s}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ by Theorem 2.4 and $2p<q$, and on the other hand, it holds that $\mathcal{C}_{\beta}(\overline{\rho}_{s})\succ 0$ by virtue of Proposition 3.9. We deduce that
    $\mathbf{E}\left\lVert\sigma\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\rho}_{s}\right)}\right\rVert_{\rm F}^{p}\leqslant CJ^{-\frac{p}{2}}.$
Combining the bounds on both terms, we deduce that
$\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u\wedge\theta_{J}}-\overline{X}^{(j)}_{u\wedge\theta_{J}}\right\rvert^{p}\right]\,\mathrm{d}s.$
Applying a similar reasoning to the drift terms, we finally obtain from (5.4) that
$\mathbf{E}\left[\sup_{s\in[0,t]}\Big{\lvert}X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\bigl{\lvert}X_{u\wedge\theta_{J}}^{(j)}-\overline{X}_{u\wedge\theta_{J}}^{(j)}\bigr{\rvert}^{p}\right]\,\mathrm{d}s.$
By Grönwall’s inequality, this implies that
$\mathbf{E}\left[\sup_{t\in[0,T]}\Bigl{\lvert}X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\Bigr{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
Bounding the second term in (5.3)
We have by Lemma 3.5 and (2.6) that
$\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\big{\rvert}^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}\big{\rvert}^{q}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}\overline{X}^{(t)}_{j}\big{\rvert}^{q}\right]\right)\leqslant 2^{q}\kappa.$
In addition, by Lemma 2.5 it holds that
$\mathbf{P}[\theta_{J}(R)\leqslant T]=\mathbf{P}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\right]\leqslant CJ^{-2}.$
Conclusion
We finally obtain from (5.3) that
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
which concludes the proof. ∎
5.2 Improving the rate of convergence for Gaussian initial conditions.
It is proved in [11] that, if the initial condition to the mean-field equation 1.5 associated with CBS is Gaussian, then at each time $t$ the solution $\overline{\rho}_{t}$ to the equation is also Gaussian. In this setting, the probability measure $\overline{\rho}_{0}$ has moments of all orders. Using a concentration inequality for Gaussian measures (see [43] and [42, Section 2]), it is possible to prove that the probability of the event
$\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\lvert\overline{X}^{(j)}_{t}(\omega)\rvert^{p}\geqslant R\right\}$
decreases in fact as $\exp\bigl{(}-CJ^{\frac{p}{2}}\bigr{)}$, for a constant $C$ depending on $R$ and on the parameters of the Gaussian distribution. Carrying out the proof of Theorem 2.6 with the sharper bound $\exp\bigl{(}-CJ^{\frac{p}{2}}\bigr{)}$ instead of $CJ^{-2}$ for $\mathbf{P}(\Omega_{J,t})$ as in Lemma 2.5, we obtain for all sufficiently large $p$ a mean-field estimate of the form
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
5.3 Handling the case where $u>\ell>0$ in Assumption H
Our main results stand only when $f\in\mathcal{A}(s,\ell,u)$ with $u=\ell$. This is a convenient setting because it then holds by Proposition A.3 that the $p$-th moment of the reweighted measure $L_{\beta}\mu$ is controlled by the $p$-th moment of the $\mu$, which is key in order to prove well-posedness of the mean-field system, to obtain moment estimates for empirical measures, and, most importantly, to improve the Wasserstein stability estimates for the weighted mean and weighted covariance, see Corollary 3.3. In the interest of simplicity, we refrained from stating precise results for the cases $u>\ell=0$ and $u>\ell>0$. The former remains open, but we note that a quantitative mean-field estimate can still be obtained in the latter setting. The main ingredients to this end are the following:
  • Moment estimates for the empirical measure and well-posedness of the mean-field equation can be obtained by using Proposition A.3 together with the Bihari–LaSalle inequality instead of Grönwall’s inequality, using the same reasoning as in Lemma 3.5 and Section 4, respectively. In contrast with the setting $\ell=u$, these results hold only up to some finite time $T<T_{+}$, where $T_{+}\in(0,\infty)$ is the time at which the bounds through the Bihari–LaSalle inequality become singular.
  • In order to obtain a mean-field result up to some time $T<T_{+}$, the approach presented in Subsection 5.1 can be used, this time with the stopping time $\theta_{J}:=\tau_{J}\wedge\overline{\tau}_{J}$, where
    $\tau_{J}(R)$ $=\inf\Bigl{\{}t\geqslant 0:\mu_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}},$
    $\overline{\tau}_{J}(R)$ $=\inf\Bigl{\{}t\geqslant 0:\overline{\mu}_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}}.$
    Employing this stopping time instead of (5.2) allows for using the basic stability estimates for the weighted moments of Proposition 3.1, which are valid even when $\ell<u$, instead of the improved stability estimates in Corollary 3.3, which require that $\ell=u$. In order to complete the proof as in Subsection 5.1, it remains to bound the probability $\mathbf{P}\left[\theta_{J}\leqslant T\right]$, which can be achieved by noticing that
    $\mathbf{P}[\theta_{J}(R)\leqslant T]=\mathbf{P}\left[\tau_{J}(R)\leqslant T<\overline{\tau}_{J}(R)\right]+\mathbf{P}\left[\overline{\tau}_{J}(R)\leqslant T\right].$
    Using the almost sure continuity of the solution to the interacting particle system by Theorem 2.2, together with the triangle inequality, we bound the first probability as follows:
    $\mathbf{P}[\tau_{J}\leqslant T<\overline{\tau}_{J}]$ $\leqslant\mathbf{P}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}X^{(j)}_{t\wedge\theta_{J}}\Bigr{\rvert}^{p}=R\right]$
    $\leqslant\mathbf{P}\left[\frac{2^{p-1}}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}+\frac{2^{p-1}}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t\wedge\theta_{J}}-\overline{X}^{(j)}_{t\wedge\theta_{J}}\Bigr{\rvert}^{p}\geqslant R\right]$
    $\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t\wedge\theta_{J}}-\overline{X}^{(j)}_{t\wedge\theta_{J}}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right],$
    where we used that $\mathbf{P}[A+B\geqslant k]\leqslant\mathbf{P}[A\geqslant k/2]+\mathbf{P}[B\geqslant k/2]$ for any two real-valued random variables $A$ and $B$, because $\{A+B\geqslant k\}\subset\{A\geqslant k/2\}\cup\{B\geqslant k/2\}$. A similar trick, of bounding the probability of an event associated with a stochastic system in terms of a similar event associated with a simpler system and the discrepancy between the two systems, is employed in [19, Theorem 2.10]. The probability $\mathbf{P}[\theta_{J}\leqslant T]$ can then be bounded as follows:
    $\mathbf{P}[\theta_{J}(R)\leqslant T]$ $=\mathbf{P}\left[\tau_{J}(R)\leqslant T<\overline{\tau}_{J}\right]+\mathbf{P}\left[\overline{\tau}_{J}(R)\leqslant T\right]$
    $\leqslant 2\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t\wedge\theta_{J}}-\overline{X}^{(j)}_{t\wedge\theta_{J}}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right].$
    Using Markov’s inequality for the second term, we then obtain
    $\mathbf{P}[\theta_{J}(R)\leqslant T]\leqslant 2\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\frac{2^{p}}{R}\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t\wedge\theta_{J}}-\overline{X}^{(j)}_{t\wedge\theta_{J}}\Bigr{\rvert}^{p}\right].$ (5.7)
    The first term is bounded from above by $CJ^{-2}$ by Lemma 2.5, whereas the second term can be bounded by $CJ^{-\frac{p}{2}}$ as in the proof in Subsection 5.1. Thus, the probability $\mathbf{P}[\theta_{J}(R)\leqslant T]$ scales as $J^{-\min\{2,\frac{p}{2}\}}$, instead of $J^{-2}$ in the case where $u=\ell$. As a consequence, we obtain a mean-field result of the type
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{2,\frac{p}{2}\right\}\frac{q-p}{q}}.$
5.4 Removing the lower bound on $p$
In Theorems 2.6 and 2.12, we assume that $p$ is bounded from below by $p_{\mathcal{M}}$ and $2p_{\mathcal{M}}$, respectively. It is possible to replace this restriction by the less restrictive inequality $p\geqslant 2$ in the case of Theorem 2.6, and $p>2$ in the case of Theorem 2.12, provided that the initial condition $\overline{\rho}_{0}$ has sufficiently many moments. We motivate this statement hereafter for the specific case of CBO. Suppose for simplicity that $\overline{\rho}_{0}$ has finite moments of all orders. Then by Theorem 2.6, for $f\in\mathcal{A}(s,\ell,\ell)$ there exists an exponent $p_{\star}\geqslant 2$ such that for all $p\geqslant p_{\star}$ and all $\varepsilon>0$, it holds for some constant $C$ independent of $J$ that
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-(2-\varepsilon)}.$ (5.8)
  • First, note that since
    $\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{r}\right]\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{r}{p}}$
    for any $1\leqslant r\leqslant p$, the existence of a mean-field result of the type (5.8) with $p=p_{\star}$ immediately implies a similar result for all $r\in[2,p_{\star}]$, but with a rate that is degraded by a factor $\frac{r}{p_{\star}}$. In the next bullet point, we present a better approach.
  • Fix $r\in[2,p_{\star}]$. From the proof of Proposition 3.1, it is easily observed that there is $\varrho$ sufficiently large to ensure that for all $R>0$, it holds that
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\times\mathcal{P}_{\varrho,R},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert+\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L\,W_{r}(\mu,\nu),$ (5.9)
    for some constant $L=L(R,r)>0$, see Lemma A.6. In order to obtain a mean-field result for the exponent $r$, we can employ the same approach as in the proof in Subsection 5.1 with $p$ replaced by $r$, but with the stopping time
    $\theta_{J}:=\overline{\tau}_{J}\wedge\tau_{J},\qquad\left\{\begin{aligned} \tau_{J}(R)&=\inf\Bigl{\{}t\geqslant 0:\mu_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}},\\ \overline{\tau}_{J}(R)&=\inf\Bigl{\{}t\geqslant 0:\overline{\mu}_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}}.\end{aligned}\right.$
    for some $p\geqslant p_{\star}\vee\varrho$. The proof then goes through as before, except that we use (5.9) instead of Corollary 3.3, and that the probability $\mathbf{P}[\theta_{J}\leqslant T]$ is bounded from above differently. Reasoning as in (5.7), it is simple to show that
    $\mathbf{P}[\theta_{J}\leqslant T]\leqslant 2\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\frac{2^{p}}{R}\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\right].$
    The first term on the right-hand side decreases as $\mathcal{O}(J^{-2})$ by Lemma 2.5, whereas the second term decreases as $\mathcal{O}(J^{-(2-\varepsilon)})$ by (5.8). This enables to prove that
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{r}\right]\leqslant CJ^{-\min\left\{\frac{r}{2},2-\varepsilon\right\}}.$ (5.10)
Appendix A Further technical results
A.1 Stability estimates for ratios of integrals
Lemma A.1.
[…]
$\forall x,y\in\mathbf{R}^{d},\qquad\bigl{\lVert}g(x)-g(y)\bigr{\rVert}\vee\bigl{\lvert}h(x)-h(y)\bigr{\rvert}\leqslant L\Bigl{(}1+|x|+|y|\Bigr{)}^{\varsigma}\lvert x-y\rvert.$ (A.1)
Then for all $p\geqslant\varsigma+1$ and all $R>0$, there exist a constant $C=C(p,R)$ such that
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert\leqslant CW_{p}(\mu,\nu).\qquad$
Proof.
Fix $p\geqslant\varsigma+1$ and $R>0$.
Step 1
Let us first generalize [10, Lemma 3.1] by showing that there exists a constant $\eta=\eta(p,R)>0$ such that
$\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\int_{\mathbf{R}^{d}}h\,\mathrm{d}\mu\geqslant\eta.$ (A.2)
Define $B_{K}:=\{x\in\mathbf{R}^{d}:\lvert x\rvert\leqslant K\}$ where $K:=2^{1/p}R$. By Markov’s inequality, we have
$\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\mu\left(B_{K}\right)\geqslant 1-\mu\left(\mathbf{R}^{d}\setminus B_{K}\right)\geqslant 1-\frac{R^{p}}{K^{p}}=\frac{1}{2}.$
Therefore we obtain
$\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\int_{\mathbf{R}^{d}}h\,\mathrm{d}\mu\geqslant\int_{B_{K}}h\,\mathrm{d}\mu\geqslant\frac{1}{2}\min\bigl{\{}h(x):x\in B_{K}\bigr{\}}=:\eta>0,$
since $h$ is continuous by A.1. This proves (A.2).
Step 2
Fix probability measures $\mu,\nu\in\mathcal{P}_{p,R}$. We omit the integration domains for conciseness. For any coupling $\pi\in\Pi(\mu,\nu)$, we have by (A.1) and (A.2) that
$\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert$ $\leqslant\left\lVert\frac{\iint g(x)-g(y)\,\pi(\mathrm{d}x\,\mathrm{d}y)}{\mu[h]}\right\rVert+\left\lVert\int g\,\mathrm{d}\nu\right\rVert\cdot\left\lvert\frac{\iint h(x)-h(y)\,\pi(\mathrm{d}x\,\mathrm{d}y)}{\mu[h]\nu[h]}\right\rvert$
$\leqslant L\left(\eta^{-1}+\eta^{-2}\int\left\lVert g\right\rVert\,\mathrm{d}\nu\right)\iint\Bigl{(}1+|x|+|y|\Bigr{)}^{\varsigma}\lvert x-y\rvert\,\pi(\mathrm{d}x\,\mathrm{d}y).$
By (A.1), there exists $C>0$ such that $\left\lVert g(x)\right\rVert\leqslant C\bigl{(}1+|x|\bigr{)}^{\varsigma+1}$ for all $x\in\mathbf{R}^{d}$. The claim then follows by first applying Hölder’s inequality to obtain the bound
$\iint\Bigl{(}1+|x|+|y|\Bigr{)}^{\varsigma}\lvert x-y\rvert\,\pi(\mathrm{d}x\,\mathrm{d}y)\leqslant\left(\iint\Bigl{(}1+|x|+|y|\Bigr{)}^{\frac{\varsigma p}{p-1}}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{p-1}{p}}\left(\iint\lvert x-y\rvert^{p}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{1}{p}},$
noting that $\frac{\varsigma p}{p-1}\leqslant p$ under the assumption that $p\geqslant\varsigma+1$, and then infimizing over $\pi\in\Pi(\mu,\nu)$. ∎
Lemma A.2.
Let $g\colon\mathbf{R}^{d}\rightarrow\mathbf{R}^{m\times n}$, $h\colon\mathbf{R}^{d}\rightarrow(0,\infty)$ and $\mathcal{S}\colon\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{n}$ be functions such that the following conditions hold.
  1. (a)
    There is $\varsigma\geqslant 0$ and $L>0$ such that
    $\forall x,y\in\mathbf{R}^{d},\qquad\bigl{\lVert}g(x)-g(y)\bigr{\rVert}\vee\bigl{\lvert}h(x)-h(y)\bigr{\rvert}\leqslant L\Bigl{(}1+|x|+|y|\Bigr{)}^{\varsigma}\lvert x-y\rvert.$
  2. (b)
    The map $\mathcal{S}\colon\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{n}$ satisfies
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\lvert\mathcal{S}(\mu)-\mathcal{S}(\nu)\rvert\leqslant L_{\mathcal{S}}W_{p}(\mu,\nu).$
Then, for all $R>0$ and $p\geqslant 2\varsigma+2$, there exist a constant $C>0$ such that for all $\mu,\nu\in\mathcal{P}_{p,R}$ and all couplings $\pi\in\Pi(\mu,\nu)$, it holds that
$\left(\iint_{\mathbf{R}^{d}\times\mathbf{R}^{d}}\left\lVert\frac{g(x)\mathcal{S}(\mu)}{\sqrt{\mu[h]}}-\frac{g(y)\mathcal{S}(\nu)}{\sqrt{\nu[h]}}\right\rVert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/2}\leqslant C\,\left(\int\lvert x-y\rvert^{p}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/p}.$ (A.3)
Proof.
Fix $p\geqslant 2\varsigma+2$ and $R>0$. Again, there is a constant $\eta=\eta(R,p)>0$ such that (A.2) holds. From the triangle inequality we obtain
$\left\lVert\frac{g(x)\mathcal{S}(\mu)}{\sqrt{\mu[h]}}-\frac{g(y)\mathcal{S}(\nu)}{\sqrt{\nu[h]}}\right\rVert_{L^{2}(\pi)}$ $\leqslant\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\Bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\Bigr{)}\right\rVert_{L^{2}(\pi)}+\left\lVert\Bigl{(}g(x)-g(y)\Bigr{)}\frac{\mathcal{S}(\nu)}{\sqrt{\mu[h]}}\right\rVert_{L^{2}(\pi)}$
$\quad+\left\lVert{g(y)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)}.$
where, by a slight abuse of notation, we wrote $g(x)$ and $g(y)$ in order to refer to the functions $(x,y)\mapsto g(x)$ and $(x,y)\mapsto g(y)$, respectively. In the rest of the proof, we denote by $C$ any constant depending only on the parameters $(R,L,\varsigma,\eta,p)$.
Bounding the first term
It follows from (a) that $\left\lVert g(x)\right\rVert^{2}\leqslant C(1+|x|)^{2\varsigma+2}$ for all $x\in\mathbf{R}^{d}$, for a constant $C$. Using (b), we obtain the bound
$\left(\iint\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{)}\right\rVert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/2}\leqslant\sqrt{\frac{\mu[\lVert g(x)\rVert^{2}]}{\mu[h]}}\bigl{\lvert}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{\rvert}\leqslant\sqrt{\frac{C}{\eta}}L_{\mathcal{S}}W_{p}(\mu,\nu).$
Bounding the second term
For the second term, we use (a) to find
$\forall x,y\in\mathbf{R}^{d},\qquad\lvert g(x)-g(y)\rvert^{2}\leqslant L^{2}\left(1+|x|+|y|\right)^{2\varsigma}\lvert x-y\rvert^{2}.$
Using Hölder’s inequality, we therefore obtain
$\bigl{\lVert}g(x)-g(y)\bigr{\rVert}_{L^{2}(\pi)}$ $\leqslant L\left(\int\Bigl{(}1+|x|+|y|\Bigr{)}^{\frac{2\varsigma p}{p-2}}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{p-2}{2p}}\left(\int\lvert x-y\rvert^{p}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{1}{p}}$
$\leqslant C\left(\int\lvert x-y\rvert^{p}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{\frac{1}{p}},$
where we used that $\frac{2\varsigma p}{p-2}\leqslant p$ since $p\geqslant 2+2\varsigma$. In addition, from (b) we have
$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R.$ (A.4)
and so we conclude that
$\left\lVert\Bigl{(}g(x)-g(y)\Bigr{)}\frac{\mathcal{S}(\nu)}{\sqrt{\mu[h]}}\right\rVert_{L^{2}(\pi)}\leqslant C\left(\int\lvert x-y\rvert^{p}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/p}.$ (A.5)
Bounding the third term
Using again (A.4), we have
$\left\lVert{g(y)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)}=\lVert g\rVert_{L^{2}(\mu)}\left\lvert\mathcal{S}(\nu)\right\rvert\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\leqslant C\left\lvert\frac{\sqrt{\mu[h]}-\sqrt{\nu[h]}}{\sqrt{\mu[h]}\sqrt{\nu[h]}}\right\rvert.$
Since $\sqrt{a}-\sqrt{b}=(a-b)/(\sqrt{a}+\sqrt{b})$ for all $a,b>0$, we deduce that
$\left\lVert{g(x)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)}$ $\leqslant C\eta^{-{\frac{3}{2}}}\iint\bigl{\lvert}h(x)-h(y)\bigr{\rvert}\,\pi(\mathrm{d}x\,\mathrm{d}y)$
$\leqslant C\iint\bigl{(}1+|x|+|y|\bigr{)}^{\varsigma}|x-y|\,\pi(\mathrm{d}x\,\mathrm{d}y).$
By employing Hölder’s inequality, the integral can be bounded by the right-hand side of (A.5) up to multiplication by a constant.
Conclusion
The claim follows then by combining the bounds on the three terms, and noting, for the first term, that $W_{p}(\mu,\nu)$ is bounded from above by the bracket on the right-hand side of (A.3) by definition of the Wasserstein distance. ∎
A.2 Bound on weighted moments
If the function $f$ is bounded from below and above, it is trivial to show that
$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x).$
for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved. The following is a generalization of [10, Lemma 3.3].
Proposition A.3.
Suppose that $f$ is bounded from below by $f_{\star}$ and that there are real numbers $u\geqslant\ell>0$ and positive constants $c_{u},c_{\ell}$ such that
$\forall x\in\mathbf{R}^{d},\qquad\left\{\begin{aligned} f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)}\\ f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.\end{aligned}\right.$ (A.6)
Then for all $p,q>0$ there are constants $C_{1}$ and $C_{2}$ depending on $(p,q,u,\ell,c_{u},c_{\ell})$ such that
$\forall\mu\in\mathcal{P}_{q}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant\left(C_{1}+C_{2}\int_{\mathbf{R}^{d}}\lvert x\rvert^{q}\,\mu(\mathrm{d}x)\right)^{\frac{pu}{q\ell}}$ (A.7)
Furthermore, if $u=\ell$ and $p\leqslant q$, then this statement holds with $C_{1}=0$.
The proof relies on the following lemma.
Lemma A.4.
For all $p,q,\beta>0$, there is $C=C(p,q,\beta)$ such that
$\forall\nu\in\mathcal{P}_{q}(\mathbf{R}^{+}),\qquad\frac{\int_{\mathbf{R}^{+}}y^{p}\,\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}\leqslant C\left(1+\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{p}{q}},$ (A.8)
where $\mathbf{R}^{+}:=[0,\infty)$. The constant is given by
$C(q,p,\beta)=\left(1-\frac{1}{2^{q}}\right)^{-1}(p\beta^{-1})^{p}\operatorname{e}^{2\beta-p}.$
Proof.
Fix $\nu\in\mathcal{P}(\mathbf{R}^{+})$. If the right-hand side of (A.8) is infinite, then the inequality (A.8) holds with any constant $C$, so we assume from now on that
$\lambda:=\left(\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{1}{q}}<\infty.$
Step 1
Suppose first that $\lambda\leqslant 1$. Then, by Markov’s inequality, it holds that
$\nu\bigl{(}[0,2]\bigr{)}=1-\int_{\mathbf{R}^{+}\setminus[0,2]}\,\nu(\mathrm{d}y)\geqslant 1-\frac{1}{2^{q}}\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\geqslant 1-\frac{1}{2^{q}}.$
Using this inequality and the bound $y^{p}\operatorname{e}^{-\beta y}\leqslant(p\beta^{-1})^{p}\operatorname{e}^{-p}$, which is valid for all $y\in\mathbf{R}^{+}$ and obtained by maximization of $y\mapsto y^{p}\operatorname{e}^{-\beta y}$, we have that
$\frac{\int_{\mathbf{R}^{+}}y^{p}\,\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}\leqslant\frac{\int_{\mathbf{R}^{+}}(p\beta^{-1})^{p}\operatorname{e}^{-p}\,\nu(\mathrm{d}y)}{\int_{[0,2]}\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}\leqslant\left(1-\frac{1}{2^{q}}\right)^{-1}(p\beta^{-1})^{p}\operatorname{e}^{2\beta-p}.$
It follows that (A.8) is satisfied in this case.
Step 2
Assume now that $\lambda\geqslant 1$, and let $\pi$ denote the pushforward of $\nu$ under the contraction $x\mapsto\lambda^{-1}x$. Then, applying a change of variable, using Step 3 and using Step 1 step, we have that
$\frac{\int_{\mathbf{R}^{+}}y^{p}\,\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}$ $=\lambda^{p}\frac{\int_{\mathbf{R}^{+}}z^{p}\,\operatorname{e}^{-\beta\lambda z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta\lambda z}\,\pi(\mathrm{d}z)}$
$\leqslant\lambda^{p}\frac{\int_{\mathbf{R}^{+}}z^{p}\,\operatorname{e}^{-\beta z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta z}\,\pi(\mathrm{d}z)}\leqslant\lambda^{p}C(p,q,\beta),$ (A.9)
which is the required inequality.
Step 3
In the first inequality of (A.9), we used that, for any increasing function $g\colon\mathbf{R}^{+}\to\mathbf{R}$, the function
$F(\gamma):=\frac{\int_{\mathbf{R}^{+}}g(z)\,\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}$
is a non-increasing function of $\gamma$. This follows from [11, Lemma A.3], and could also be proved by showing that $F^{\prime}(\gamma)$ is non-positive for all $\gamma\geqslant 0$. A simple calculation indeed gives that
$F^{\prime}(\gamma)=-\int_{\mathbf{R}^{+}}g(z)\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z),\qquad L_{\gamma}\pi(\mathrm{d}z):=\frac{\operatorname{e}^{-\gamma z}\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)},\qquad\overline{z}:=\int_{\mathbf{R}^{+}}z\,L_{\gamma}\pi(\mathrm{d}z).$
By decomposing the integration domain into the union $[0,\overline{z})\cup\{\overline{z}\}\cup(\overline{z},\infty)$, and using the fact that $g$ is non-decreasing, we have
$F^{\prime}(\gamma)$ $=-\int_{[0,\overline{z})}g(z)\left(z-\overline{z}\right)\,\pi(\mathrm{d}z)-\int_{(\overline{z},\infty)}g(z)\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z)$
$\leqslant-g(\overline{z})\int_{[0,\overline{z})\cup(\overline{z},\infty)}\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z)=0,$
which completes the proof. ∎
Proof of Proposition A.3.
We assume without loss of generality that $f_{\star}=0$. By the second inequality in (A.6), we have that
$\lvert x\rvert^{p}=\left(\lvert x\rvert^{\ell}\right)^{\frac{p}{\ell}}\leqslant\left(\frac{f(x)}{c_{\ell}}+1\right)^{\frac{p}{\ell}}.$ (A.10)
For convenience, we denote by $C$ any constant depending only on $(p,q,\ell,u,c_{\ell},c_{u})$. From (A.10) we deduce that
$\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}$ $\leqslant C\frac{\int_{\mathbf{R}^{d}}\bigl{(}f(x)+1\bigr{)}^{\frac{p}{\ell}}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}$
$=C\frac{\int_{\mathbf{R}^{+}}\bigl{(}y+1\bigr{)}^{\frac{p}{\ell}}\,\operatorname{e}^{-\beta y}\,f_{\sharp}\mu(\mathrm{d}y)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta y}\,f_{\sharp}\mu(\mathrm{d}y)},$
where $f_{\sharp}\mu\in\mathcal{P}(\mathbf{R}^{+})$ is the pushforward of $\mu$ by $f$. Using Lemma A.4 and then the first inequality in (A.6), we obtain
$\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}$ $\leqslant C\left(1+\int_{\mathbf{R}^{+}}\lvert y\rvert^{\frac{q}{u}}\,f_{\sharp}\mu(\mathrm{d}y)\right)^{\frac{pu}{q\ell}}$
$=C\left(1+\int_{\mathbf{R}^{d}}\bigl{\lvert}f(x)\bigr{\rvert}^{\frac{q}{u}}\,\mu(\mathrm{d}x)\right)^{\frac{pu}{q\ell}}\leqslant C\left(1+\int_{\mathbf{R}^{d}}\bigl{\lvert}x\bigr{\rvert}^{q}\,\mu(\mathrm{d}x)\right)^{\frac{pu}{q\ell}},$
which completes the proof of (A.7).
Proof that (A.7) holds with $C_{1}=0$ when $\ell=u$ and $q\geqslant p$
The case $q>p$ follows from the case $q=p$, so we assume the latter. Suppose for contradiction that this was not true. Then there would exist a sequence $\{\mu_{i}\}_{i\in\mathbf{N}}$ of probability measures such that
$\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}\geqslant i\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu_{i}(\mathrm{d}x).$ (A.11)
By the first part of the proof, this inequality implies that
$\frac{C_{1}+C_{2}\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu_{i}(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu_{i}(\mathrm{d}x)}\geqslant i,$
and so the integral in this expression necessarily converges to 0 in the limit as $i\to\infty$. In turn, this implies that the sequence $\{\mu_{i}\}_{i\in\mathbf{N}}$ converges in distribution to the Dirac measure at 00 when $i\to\infty$. But then it holds that
$\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}\left(\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu_{i}(\mathrm{d}x)\right)^{-1}\leqslant\frac{\operatorname{e}^{-\beta f_{\star}}}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}\xrightarrow[i\to\infty]{}\operatorname{e}^{-\beta\bigl{(}f_{\star}-f(0)\bigr{)}},$
which contradicts (A.11). ∎
Remark A.5.
When $p=q>0$, Proposition A.3 implies that
$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant\left(C_{1}+C_{2}\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x)\right)^{\frac{u}{\ell}}.$ (A.12)
We motivate in this remark that the exponent $\frac{u}{\ell}$ is optimal. Indeed, let $p=1$ and consider the function $f\colon\mathbf{R}^{2}\to\mathbf{R}$ given by $f(x_{1},x_{2})=\lvert x_{1}\rvert^{\ell}+\lvert x_{2}\rvert^{u+\varepsilon}$, for any $u\geqslant\ell>0$ and $\varepsilon>0$. Then for the probability measures
$\mu_{i}=\frac{1}{i^{u}}\delta_{\left(i^{u},0\right)}+\left(1-\frac{1}{i^{u}}\right)\delta_{\left(0,i^{\ell}\right)},$
it holds that
$\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert\,\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu_{i}(\mathrm{d}x)}\sim i^{u}\qquad\text{ and }\qquad\int_{\mathbf{R}^{2}}\lvert x\rvert\,\mu_{i}(\mathrm{d}x)\sim i^{\ell}\qquad\text{ as $i\to\infty$},$
where $\sim$ signifies asymptotic equivalence. This example suggests that Lemma A.4 is sharp.
A.3 Generalized Wasserstein stability estimates
By adapting the proof of Proposition 3.1, it is simple to prove Wasserstein stability estimates with different values of $p$ on the left and right sides of the inequality. More precisely, we have the following statement.
Lemma A.6.
Suppose that $f\in\mathcal{A}(s,0,u)$. Then
  • for all $R>0$ and $p>1$, there is $C=C(p,R)$
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant C\,W_{p}(\mu,\nu),\qquad\varrho:=\frac{(s+1)p}{p-1}$ (A.13)
  • for all $R>0$ and $p>1$, there is $C=C(p,R)$
    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\rvert\leqslant C\,W_{p}(\mu,\nu),\qquad\varrho:=\frac{(s+2)p}{p-1}$ (A.14)
Remark A.7.
For $p=s+2$, the bound (A.13) coincides with (3.1), and for $p=s+3$, the bound (A.14) coincides with (3.2).
Acknowledgments
We are grateful to Yunan Yang, Arnaud Guyader and Louis-Pierre Chaintron for useful discussions, and to the Lorentz Center (Leiden, the Netherlands) for hosting the workshop “Purpose-driven particle systems” in March 2023, which enabled discussions about this work. FH is supported by start-up funds at the California Institute of Technology. FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM.

### v2

@@
-latexOverwriting file
+Mean-field limits for Consensus-Based Optimization and Sampling
 Abstract
@@
 In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling.
-Our approach extends Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems.
+Our approach requires a generalization of Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems.
+In addition, we present new results on the well-posedness of the particle systems and their mean-field limit, and provide novel stability estimates for the weighted mean and the weighted covariance.
 Keywords. Mean-field limits, Interacting particle systems, Consensus-Based Optimization, Consensus-Based Sampling, Coupling methods, Wasserstein stability estimates.
@@
 This work focuses on the rigorous derivation of quantitative mean-field limits for recently proposed algorithms known as Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants.
-The former, initially proposed in [46], is an algorithm for global optimization tasks, while the latter, proposed in [11], can be employed for both sampling and optimization purposes.
+The former, initially proposed in [CBO], is an algorithm for global optimization tasks, while the latter, proposed in [CBS-Carrillo2021], can be employed for both sampling and optimization purposes.
 Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels.
@@
 Given an objective function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, we consider the following version of Consensus-Based Optimization (CBO):
-$\mathrm{d}X^{(j)}_{t}$ $=-\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}t+S\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}W^{(j)}_{t},\qquad t\geqslant 0,\qquad j\in,$ (1.1)
-where $J>1$ is a natural number, :={$:=\{1,\dots,J\}$ and $\mu^{J}_{t}:=\frac{1}{J}\sum_{i=1}^{J}\delta_{X^{(i)}_{t}}$ denotes the empirical measure associated with the particle system.
+$\mathrm{d}X^{(j)}_{t}$ $=-\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}t+S\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}W^{(j)}_{t},\qquad t\geqslant 0,\qquad j\in\llbracket 1,J\rrbracket,$ (1.1)
+where $J>1$ is a natural number, $\llbracket 1,J\rrbracket:=\{1,\dots,J\}$ and $\mu^{J}_{t}:=\frac{1}{J}\sum_{i=1}^{J}\delta_{X^{(i)}_{t}}$ denotes the empirical measure associated with the particle system.
 For a probability measure $\mu\in\mathcal{P}(\mathbf{R}^{d})$, the notation $\mathcal{M}_{\beta}(\mu)$ denotes the mean of $\mu$ after reweighting by $\operatorname{e}^{-\beta f}$.
@@
 The weighted mean $\mathcal{M}_{\beta}(\mu^{J}_{t})$ may be viewed as an approximation of the particle position $X^{(j)}_{t}$ such that $f(X^{(j)}_{t})$ is minimal.
-In (1.1), the processes $\bigl{(}W^{(j)}_{t}\bigr{)}_{j\in}$ are independent $\mathbf{R}^{d}$-valued Brownian motions and $S\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ is a globally Lipschitz continuous function.
-The two diffusion types commonly studied in the literature are isotropic diffusion [46, 10, 23], with $S(u)=\sqrt{2\theta}\lvert u\rvert\operatorname{I}_{d}$, and anisotropic or component-wise [13, 24] diffusion with $S(u)=\sqrt{2\theta}\operatorname{diag}(u)$.
+In (1.1), the processes $\bigl{(}W^{(j)}_{t}\bigr{)}_{j\in\llbracket 1,J\rrbracket}$ are independent $\mathbf{R}^{d}$-valued Brownian motions and $S\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ is a globally Lipschitz continuous function.
+The two diffusion types commonly studied in the literature are isotropic diffusion [CBO, carrillo2018analytical, fornasier2021consensusbased], with $S(u)=\sqrt{2\theta}\lvert u\rvert\operatorname{I}_{d}$, and anisotropic or component-wise [CBO-high-dim, FornasierKlockRiedl2022-anisotropic] diffusion with $S(u)=\sqrt{2\theta}\operatorname{diag}(u)$.
 Here $\operatorname{I}_{d}\in\mathbf{R}^{d\times d}$ is the identity matrix and $\operatorname{diag}\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ the operator mapping a vector onto a diagonal matrix with the vector as its diagonal.
-It was shown in [13] that CBO with anisotropic diffusion is well-suited for high-dimensional global optimization problems, for example those arising in machine learning.
+It was shown in [CBO-high-dim] that CBO with anisotropic diffusion is well-suited for high-dimensional global optimization problems, for example those arising in machine learning.
 Commonly a constant coefficient is included in front of the drift term in (1.1), but since this coefficient can be removed by rescaling time upon redefining the diffusion parameter $\theta$, we choose to omit it.
 By design, CBO allows passage to the corresponding mean-field description.
-Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [37], it is possible to formally write down the corresponding mean-field equation:
+Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [PSO], it is possible to formally write down the corresponding mean-field equation:
 […]
-where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [29].
-The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [34, 10, 23].
-In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [10].
+where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [CBO-PSO-zero-inertia-1].
+The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased].
+In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical].
 The distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large exponential parameter $\beta$.
-The proof in [10] covers also non-convex cost functions $f$.
-The convergence, in an appropriate sense, of the empirical measure associated with the CBO finite-size particle system (1.1) to the solution of the mean-field equation (1.3) in the large particle limit $J\to\infty$ was first proved rigorously in a recent paper [34] using a compactness argument, as we discuss further in Subsection 1.3.
-However, the method of proof employed in [34] does not provide an explicit rate of convergence with respect to the number of particles.
+The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
+The convergence, in an appropriate sense, of the empirical measure associated with the CBO finite-size particle system (1.1) to the solution of the mean-field equation (1.3) in the large particle limit $J\to\infty$ was first proved rigorously in a recent paper [CBO-mfl-Huang2021] using a compactness argument, as we discuss further in Subsection 1.3.
+However, the method of proof employed in [CBO-mfl-Huang2021] does not provide an explicit rate of convergence with respect to the number of particles.
 Obtaining such an explicit rate of convergence is highly desirable, as it would enable to transfer convergence results concerning the mean-field equation (1.3) to the finite ensemble size setting (1.1) by means of a triangle inequality argument, modulo appropriate interfacing of the metrics.
-This is the approach proposed in [23, Theorem 13], where the authors obtain a partial mean-field result in a set of high probability, enabling them to prove convergence with high probability to the global minimizer in the finite ensemble size setting.
-Here, we provide a stronger version of the mean-field limit result given in [23, Proposition 15], extending to more general diffusion terms and providing scope for global convergence results in a stronger metric than convergence in probability.
+This is the approach proposed in [fornasier2021consensusbased, Theorem 13], where the authors obtain a partial mean-field result in a set of high probability, enabling them to prove convergence with high probability to the global minimizer in the finite ensemble size setting.
+Here, we provide a stronger version of the mean-field limit result given in [fornasier2021consensusbased, Proposition 15], extending to more general diffusion terms and providing scope for global convergence results in a stronger metric than convergence in probability.
 A number of extensions of CBO have been proposed in the literature.
-As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [13].
-Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [30, 31, 41].
-(On this subject, we point out that the results in [31] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [35], memory effects [49], and Lévy noise [36], as well as extensions to multi-objective optimization [40, 7] and global optimization with constraints [21, 22, 12, 8, 32, 4].
-Moreover, a variant tailored for objective functions with several minima was proposed in [9], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [29, 17].
-Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [48].
+As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [CBO-high-dim].
+Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [Ha2019, Ha2020ConvergenceAE, HA_Discrete_CBO_hetero_noise].
+(On this subject, we point out that the results in [Ha2020ConvergenceAE] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [jin2018random], memory effects [MR4160256], and Lévy noise [MR4553241], as well as extensions to multi-objective optimization [klamroth2022consensusbased, borghi2022adaptive] and global optimization with constraints [CBO-hypersurface-Fornasier_2020, MR4329816, CTV21, MR4538901, MR4404221, MR4337731].
+Moreover, a variant tailored for objective functions with several minima was proposed in [bungert], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [CBO-PSO-zero-inertia-1, CBO-PSO-zero-inertia-2].
+Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [CBO-survey-Totzeck2022].
+Furthermore, a free software implementation of CBO and some of its extensions was recently published [Bailo2024].
 1.2 Consensus-Based Sampling (CBS)
-Consensus-Based Sampling is a sampling method inspired by the Ensemble Kalman Sampler (EKS) [26] and by consensus-based approaches such as CBO.
+Consensus-Based Sampling is a sampling method inspired by the Ensemble Kalman Sampler (EKS) [EKS-1-Hofmann-Stuart-Li-GI-2019] and by consensus-based approaches such as CBO.
 It is based on an interacting particle system with a noise term preconditioned by the square root of the empirical covariance.
 The method in the finite ensemble size and continuous time setting reads
-$\mathrm{d}X^{(j)}_{t}=-\Bigl{(}X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\,\mathrm{d}W^{(j)}_{t},\qquad j\in,$ (1.4)
+$\mathrm{d}X^{(j)}_{t}=-\Bigl{(}X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\,\mathrm{d}W^{(j)}_{t},\qquad j\in\llbracket 1,J\rrbracket,$ (1.4)
 with parameters $\beta,\lambda>0$.
@@
 $\rho_{f}:=\frac{\operatorname{e}^{-f}}{\mathcal{Z}},\qquad\mathcal{Z}:=\int_{\mathbf{R}^{d}}\operatorname{e}^{-f(x)}\,\mathrm{d}x.$
-More precisely, the authors of [11] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$.
-In addition, it was noticed in [11] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [46].
-Since the convergence analysis in [11] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$.
+More precisely, the authors of [CBS-Carrillo2021] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$.
+In addition, it was noticed in [CBS-Carrillo2021] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [CBO].
+Since the convergence analysis in [CBS-Carrillo2021] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$.
 This has so far been left as an open problem.
-Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [9], and for rare event estimation in [2].
+Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [bungert], and for rare event estimation in [2023arXiv230409077A].
 1.3 Propagation of chaos
@@
 Although the consensus-based dynamics can be written in the form (1.6), it is not possible to do so in such a way that $\widetilde{b},\widetilde{\sigma},K_{b}$ and $K_{\sigma}$ are all globally Lipschitz continuous and at the same time $K_{b}$ and $K_{\sigma}$ are globally bounded.
-Therefore, the classical McKean’s theorem [15, Theorem 3.1] is not directly applicable, and a different approach is needed.
+Therefore, the classical McKean’s theorem [ReviewChaintronII, Theorem 3.1] is not directly applicable, and a different approach is needed.
 We briefly discuss hereafter the existing mean-field results for CBO, noting that no results have been obtained so far for CBS.
-  • In [34], propagation of chaos for CBO is proved through a compactness argument based on Prokhorov’s theorem.
+  • In [CBO-mfl-Huang2021], propagation of chaos for CBO is proved through a compactness argument based on Prokhorov’s theorem.
 However, this method comes without convergence rates in terms of the number of particles $J$.
-  • Later, propagation of chaos by means of synchronous coupling was shown with optimal rates in [21] for a variant of CBO constrained to hypersurfaces.
-  • Recent results in [23] also show a probabilistic mean-field approximation of the form
+  • Later, propagation of chaos by means of synchronous coupling was shown with optimal rates in [CBO-hypersurface-Fornasier_2020] for a variant of CBO constrained to hypersurfaces.
+  • Recent results in [fornasier2021consensusbased] also show a probabilistic mean-field approximation of the form
     $\sup_{t\in[0,T]}\mathbf{E}\left[\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{2}\mathsf{1}_{\Omega\setminus\Omega_{M}}\right]\leqslant\frac{C}{J},\qquad\Omega_{M}:=\left\{\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\max\left\{\left\lvert X^{(j)}_{t}\right\rvert^{4},\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{4}\right\}\geqslant M\right\}.$
-    It is additionally proved in [23], by application of Markov’s inequality, that the set of realizations $\Omega_{M}$ has probability close to 0 for large $M$.
+    It is additionally proved in [fornasier2021consensusbased], by application of Markov’s inequality, that the set of realizations $\Omega_{M}$ has probability close to 0 for large $M$.
 Specifically, it holds that $\mathbf{P}(\Omega_{M})=\mathcal{O}(M^{-1})$.
-  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [36].
+  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [MR4553241].
 In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1.
-There are, however, several differences between our results and those presented in [36].
-Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [1, 18] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3).
-Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [36].
+There are, however, several differences between our results and those presented in [MR4553241].
+Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [MR3696003, MR2597592] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3).
+Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [MR4553241].
 1.4 Our contributions
@@
 One of these approaches is based on the careful use of indicator sets, whereas the other is based on stopping times.
-The latter approach is inspired by a method initially proposed in [33] in order to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
+The latter approach is inspired by a method initially proposed in [MR1949404] in order to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
   • In order to prove mean-field results, we generalize existing theorems and prove new results on the well-posedness of the particle systems and their mean-field limit.
@@
     $\forall x,y\in\mathbf{R}^{d},\qquad|f(x)-f(y)|\leqslant L_{f}\bigl{(}1+|x|+|y|\bigr{)}^{s}|x-y|,$ (2.1)
-    for some constants $L_{f}>0$ and $s\geqslant-1$.
+    for some constants $L_{f}>0$ and $s\geqslant 0$.
   2. (H2)
     There exist constants $c_{\ell},c_{u}>0$ and $u\geqslant\ell\geqslant 0$ such that
-    $\forall x\in\mathbf{R}^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
-    $\forall x\in\mathbf{R}^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
+    $\forall x\in^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
+    $\forall x\in^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
 Remark 2.1.
@@
   • Observe that (H1) implies (2.2a) with $u=s+1$.
-  • In the particular case where $s=1$ and either $u=\ell=0$ or $u=\ell=2$, the assumptions in Assumption H coincide with those in [10].
+  • In the particular case where $s=1$ and either $u=\ell=0$ or $u=\ell=2$, the assumptions in Assumption H coincide with those in [carrillo2018analytical].
 The well-posedness and mean-field results proved in this paper generalize beyond this setting, as explained below.
@@
 In this section, we state well-posedness results for CBO and CBS, referring for the proofs to Section 4.
-While for CBO, the following theorems are only generalizations of already existing results [10], the well-posedness proofs for the CBS particle and mean-field systems are new.
+While for CBO, the following theorems are only generalizations of already existing results [carrillo2018analytical], the well-posedness proofs for the CBS particle and mean-field systems are new.
 Theorem 2.2 (Existence and uniqueness for the particle models).
@@
 The solutions are almost surely continuous.
-In the case of CBO, Theorem 2.2 was already proved in [10, Theorem 2.1].
+In the case of CBO, Theorem 2.2 was already proved in [carrillo2018analytical, Theorem 2.1].
 We now present well-posedness results for the mean-field SDEs, beginning with CBO.
-For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [10, Theorems 3.1 and 3.2].
+For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [carrillo2018analytical, Theorems 3.1 and 3.2].
 Theorem 2.3 (CBO: Existence and uniqueness for the mean-field SDE).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant-1$, $\ell\geqslant 0$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$.
 Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.3) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$.
 Furthermore, it holds that
-[…]
+$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty\qquad\text{and}\qquad\sup_{t\in[0,T]}\Bigl{\lvert}\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{\rvert}<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.4)
 and the function $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$.
@@
 Theorem 2.4 (CBS: Existence and uniqueness for the mean-field SDE).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant-1$, $\ell\geqslant 0$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$.
 Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$.
@@
 2.3 Key preparatory lemma
-We first show a key preparatory lemma which is needed for the mean-field results.
-This result is a simple corollary of a standard non-asymptotic version of the central limit theorem known as Bikelis’ theorem, see for example [28, Theorem 3.7].
+We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of Markov’s inequality,
 Lemma 2.5 (Bound on the probability of large excursions).
 Let $(Z_{j})_{j\in\mathbf{N}^{+}}$ be a family of i.i.d.
-$\mathbf{R}$-valued random variables such that $\mathbf{E}\left[\left\lvert Z_{1}\right\rvert^{3}\right]<\infty$.
+-valued random variables such that $\mathbf{E}\left[\left\lvert Z_{1}\right\rvert^{r}\right]<\infty$ for some $r\geqslant 2$.
 Then for all $R>\mathbf{E}\bigl{[}\left\lvert Z_{1}\right\rvert\bigr{]}$, there exists a constant $C>0$ such that
-$\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant\frac{C}{J^{2}}.$
+$\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant CJ^{-\frac{r}{2}}.$
 Proof.
-Denote by $m$ and $\sigma^{2}$ the mean and variance of $Z_{j}$, respectively. Bikelis’ theorem provides the following non-asymptotic error estimate:
-$\forall z\in\mathbf{R},\qquad\bigl{\lvert}\phi_{J}(z)-\phi(z)\bigr{\rvert}\leqslant\frac{1}{\sqrt{J}}\,\left(\frac{\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]}{\sigma^{3}(1+\lvert z\rvert)^{3}}\right),\qquad\phi_{J}(z):=\mathbf{P}\left[\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\frac{Z_{j}-m}{\sigma}\leqslant z\right],$
-where $\phi$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ Gaussian law. Thus, if $R>m$, we obtain
-$\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]$ $=\mathbf{P}\left[\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\frac{Z_{j}-m}{\sigma}\geqslant\sqrt{J}\left(\frac{R-m}{\sigma}\right)\right]$
-$=1-\phi_{J}\left(\sqrt{J}k\right),\qquad k:=\frac{R-m}{\sigma}$
-$\leqslant 1-\phi\left(\sqrt{J}k\right)+\frac{1}{\sqrt{J}}\,\left(\frac{\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]}{\sigma^{3}(1+\sqrt{J}k)^{3}}\right)=\mathcal{O}(J^{-2}),$
-which completes the proof. ∎
+This follows from a generalization of Chebychev’s inequality [MR3674428, Exercise 3.21]. Let
+$X=\frac{1}{J}\sum_{j=1}^{J}Z_{j}.$
+By the classical Marcinkiewicz–Zygmund inequality, it holds that
+$\mathbf{E}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}$ $\leqslant J^{-r}C_{\rm MZ}(r)\mathbf{E}\left[\left(\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{2}\right)^{\frac{r}{2}}\right]$
+$\leqslant J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{r}\right]=J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\Bigl{\lvert}Z_{1}-\mathbf{E}[Z_{1}]\Bigr{\rvert}^{r}\right],$
+where we used the convexity of the function $x\mapsto x^{\frac{r}{2}}$ in the second inequality. Therefore, using the Markov inequality, we deduce that
+$\mathbf{P}\left[X\geqslant R\right]\leqslant\mathbf{P}\Bigl{[}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}\geqslant\bigl{(}R-\mathbf{E}[X]\bigr{)}^{r}\Bigr{]}$ $\leqslant\mathbf{E}\left[\frac{\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}}{\bigl{(}R-\mathbf{E}[X]\bigr{)}^{r}}\right]\leqslant\frac{CJ^{-\frac{r}{2}}}{\bigl{(}R-\mathbf{E}[X]\bigr{)}^{r}},$
+which concludes the proof. ∎
 2.4 Quantitative mean-field limit for CBO
 We now present and prove the mean-field results, starting with CBO.
-Synchronous coupling for McKean-Vlasov diffusions with drift $b\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d}$ and diffusion $\sigma\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d\times d}$ assumes the following setting [14, 15].
+Synchronous coupling for McKean-Vlasov diffusions with drift $b\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d}$ and diffusion $\sigma\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d\times d}$ assumes the following setting [ReviewChaintronI, ReviewChaintronII].
 Given i.i.d.
@@
 from some $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$, we consider for each $J\in\mathbf{N}^{+}$ the interacting particle system
-$\forall j\in,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
+$\forall j\in\llbracket 1,J\rrbracket,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
 to which we couple the system of i.i.d. mean-field McKean-Vlasov diffusions
-$\forall N\in\mathbf{N}^{+},\forall j\in,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
+$\forall N\in\mathbf{N}^{+},\forall j\in\llbracket 1,J\rrbracket,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
 driven by the same Brownian motions and initial conditions as (2.8). In the case of CBO, the drift and diffusion coefficients for (2.8) and (2.9) are given by
@@
 Theorem 2.6 (Mean-field limit for CBO).
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ has moments of sufficiently high order.
+Consider the systems (2.8) and (2.9) with the coefficients given by (2.10).
+Then, for each $p>0$, there exists a constant $C>0$ independent of $J$ such that
+$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\frac{1}{2}}.$ (2.11)
+Remark 2.7.
+We prove below that the following sharper statement holds.
 Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$, and consider the systems (2.8) and (2.9) with the coefficients given by (2.10).
-Then for all $p\geqslant 2$ such that
-$\frac{q}{3}\geqslant p\geqslant p_{\mathcal{M}}(s,\ell),$ (2.11)
-where $p_{\mathcal{M}}$ is given in (2.3), there is $C>0$ independent of $J$ such that
-$\forall J\in\mathbf{N}^{+},\qquad\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\min\left\{\frac{p}{2},2-\frac{2p}{q}\right\}}.$
-Remark 2.7.
-In particular, we obtain a decay bound of order $\mathcal{O}\left(\frac{1}{J}\right)$ if $p=2$.
-If $p\geqslant 4$, then for each $\varepsilon>0$, decay of order $\mathcal{O}\left(\frac{1}{J^{2-\varepsilon}}\right)$ holds provided that $\overline{\rho}_{0}\in\mathcal{P}_{q}(^{d})$ with $q\geqslant\frac{2p}{\varepsilon}\vee 3p$.
+If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
+$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-p}{2p^{2}},\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}.$ (2.12)
 Remark 2.8.
 Theorem 2.6 implies pathwise Wasserstein-$p$ chaos. More precisely, it holds that
-$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{2}{p}-\frac{2}{q}\right\}}.$
+$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\frac{1}{2}}.$
 where $\rho^{J}_{[0,T]}$ is the pathwise law of the microscopic particle system over $[0,T]$ and $\overline{\rho}_{[0,T]}$ is the pathwise mean-field law over $[0,T]$. That is to say,
@@
 $d(\mathbf{X},\mathbf{Y})=\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}-Y^{(j)}_{t}\right\rvert,\qquad\mathbf{X},\mathbf{Y}\in\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}.$
-See [14, Definitions 3.1 and 3.5] for more details.
+See [ReviewChaintronI, Definitions 3.1 and 3.5] for more details.
 Before presenting the proof, we summarize below a number of useful auxiliary results which are valid for $f\in\mathcal{A}(s,\ell,\ell)$. These results are stated precisely and proved in Section 3.
@@
 Proof of Theorem 2.6.
-Fix $j\in$. As in Sznitman’s argument [15, Theorem 3.1], we begin by writing
+As mentioned in Remark 2.7, we prove the sharp result (2.12).
+It is sufficient to prove this statement for all $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ (by the assumption on $q$, this interval is nonempty).
+Indeed, if the statement holds true in this case, then by Jensen’s inequality, it holds for all $\mathfrak{p}\in(0,2\vee p_{\mathcal{M}}]$ that
+$\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{\mathfrak{p}}\right]\right)^{\frac{1}{\mathfrak{p}}}$ $\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2\vee p_{\mathcal{M}}}\right]\right)^{\frac{1}{2\vee p_{\mathcal{M}}}}$
+$\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}=CJ^{-\min\left\{\frac{1}{2},\frac{q-\mathfrak{p}}{2\mathfrak{p}^{2}},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}},$
+where we used that the function $z\mapsto(q-z)/z^{2}$ is decreasing.
+Take $p\in[2\vee p_{\mathcal{M}},q/2]$ and fix $j\in\llbracket 1,J\rrbracket$.
+As in Sznitman’s argument [ReviewChaintronII, Theorem 3.1], we begin by writing
 $\bigl{\lvert}X_{t}^{(j)}-\overline{X}_{t}^{(j)}\bigr{\rvert}^{p}\leqslant$  $\,2^{p-1}\left\lvert\int_{0}^{t}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
 $+2^{p-1}\left\lvert\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$
-By the Burkholder–Davis–Gundy inequality, see [44, Theorem 7.3], it holds for any $t\in[0,T]$ that
+By the Burkholder–Davis–Gundy inequality, see [MR2380366, Theorem 7.3], it holds for any $t\in[0,T]$ that
 $\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant 2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert\,\mathrm{d}s\right)^{p}$
@@
 $\leqslant 2^{p-1}\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+2^{p-1}\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}.$ (2.15)
-In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [15, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
+In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [ReviewChaintronII, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
 $\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\},$ (2.16)
 where $R>\mathbf{E}[Z_{j}]$ is fixed. Define
-$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}$
-Since $q\geqslant 3p$, Theorem 2.3 implies that $\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]<\infty.$ Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
-$\forall t\in[0,T],\qquad\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\Omega_{J,t}\right]\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant\frac{C}{J^{2}}.$
-We can split the second term on the right-hand of (2.14) as follows:
+$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}.$
+By Theorem 2.3 it holds that $\mathbf{E}\left[\lvert Z_{j}\rvert^{q/p}\right]<\infty.$ Note that $q\geqslant 2p$. Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
+$\forall t\in[0,T],\qquad\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\Omega_{J,t}\right]\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant CJ^{-\frac{q}{2p}}.$
+We split the second term on the right-hand of (2.15) as follows:
 $\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]$
 $\qquad+\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right].$ (2.17)
-Let us now bound each term on the right-hand side of 2.16 separately.
+Let us now bound each term on the right-hand side of (2.17) separately.
   • Since the $p$-th moment of $\overline{\mu}_{s}^{J}$ is bounded from above by $R$ for all $\omega\in\Omega\setminus\Omega_{J,s}$, we can use the stability estimate on the weighted mean given (see Item (a) and Corollary 3.3) to obtain
@@
     Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{2(q-p)}{q}}.$
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
 We have thus shown that
-$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{2(q-p)}{q}},$
+$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}},$
 which in turn by (2.15) implies that
-$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{2(q-p)}{q}}.$
+$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
 Bounding the second term in (2.14)
@@
 Combining the bounds on both terms in (2.14), we deduce
-$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
+$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
 Using that $S\colon\mathbf{R}\rightarrow\mathbf{R}^{d\times d}$ is globally Lipschitz continuous, we obtain analogously
-$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
+$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
 Substituting in (2.13), we finally obtain
-$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
-The conclusion then follows from Grönwall’s lemma. ∎
+$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
+The conclusion then follows from Grönwall’s lemma, since $\frac{q-p}{2p^{2}}\leqslant\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}$ for the value of $p$ considered. ∎
 Remark 2.9.
-The proof is similar to that of McKean’s theorem [15, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
+The proof is similar to that of McKean’s theorem [ReviewChaintronII, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
 Remark 2.10.
@@
 Remark 2.11.
-In his famous lecture notes [47], Sznitman shows for a laboratory example that
-$\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
-More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [16].
+In his famous lecture notes [MR1108185], Sznitman shows for a laboratory example that
+$\forall j\in\llbracket 1,J\rrbracket,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
+More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [MR1476912].
 Although CBO does not fit within Sznitman’s simplified framework, Item (c) enables to show a bound with the same rate for the second term on the right-hand side of (2.14),
@@
 Theorem 2.12 (Mean-field limit for CBS).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s\geqslant-1$, $\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$, and consider the systems (2.8) and (2.9) with the coefficients given by
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$ has moments of sufficiently high order.
+Consider the systems (2.8) and (2.9) with the coefficients given by
 $b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}(\mu)}.$ (2.18)
-Then for all $p\geqslant 2$ such that
-$\frac{q}{3}\geqslant p\geqslant 2p_{\mathcal{M}}(s,\ell)=\begin{cases}2s+4,\,&\text{if $\ell=0$},\\ 2,\,&\text{if $\ell>0$},\end{cases}$
-there is $C>0$ independent of $J$ such that
-$\forall J\in\mathbf{N}^{+},\qquad\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{\frac{p}{2},2-\frac{2p}{q}\right\}}.$
+Then for all $p>0$, there is $C>0$ independent of $J$ such that
+$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\frac{1}{2}}.$ (2.19)
+Remark 2.13.
+In the same way as for CBO, it is possible to prove the following sharper statement.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$.
+Consider the system (2.8) and (2.9) with the coefficients given by (2.18).
+If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
+$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-p}{2p^{2}},\frac{q-2p_{\mathcal{M}}}{2(2p_{\mathcal{M}})^{2}}\right\}}.$ (2.20)
 We summarize below the additional auxiliary results necessary for the proof, which are valid for an objective function $f\in\mathcal{A}(s,\ell,\ell)$.
@@
 3.1 Wasserstein stability estimates for the weighted moments
-In this subsection, we generalize [10, Lemma 3.2] to the more general conditions in Assumption H and, more importantly, to the weighted covariance and its square root.
+In this subsection, we generalize [carrillo2018analytical, Lemma 3.2] to the more general conditions in Assumption H and, more importantly, to the weighted covariance and its square root.
 Proposition 3.1 (Basic stability estimates for the weighted mean and the weighted covariance).
-Suppose that $f\in\mathcal{A}(s,\ell,u)$ with $s\geqslant-1$ and $u\geqslant\ell\geqslant 0$. Then:
+Suppose that $f\in\mathcal{A}(s,\ell,u)$ with $s\geqslant 0$ and $u\geqslant\ell\geqslant 0$. Then:
   1. (a)
@@
 Proof.
-We prove first (a) and then (b). Both proofs rely on the technical result Lemma A.1 in the appendix.
+The proof for both (a) and then (b) relies on the technical result Lemma A.1 in the appendix.
 Proof of (a)
@@
 Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$.
-By [51, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$.
-Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [51, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$.
+By [MR2459454, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$.
+Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [MR2459454, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$.
 Then
@@
 Proof of the statement for empirical measures
-Fix $R>0$. By [52, p.5], the Wasserstein distance between empirical measures $\mu^{J}$ and $\nu^{J}$ of the form (3.6) is equal to
+Fix $R>0$. By [MR1964483, p.5], the Wasserstein distance between empirical measures $\mu^{J}$ and $\nu^{J}$ of the form (3.6) is equal to
 $W_{p}(\mu^{J},\nu^{J})=\min_{\sigma\in\mathcal{S}_{J}}\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{\sigma(j)}\right\rvert^{p}\right)^{\frac{1}{p}},$
@@
 Proceeding in the same manner, we construct a matrix $M_{\mathbf{Y}}\in\mathbf{R}^{d\times J}$ such that $\mathcal{M}_{\beta}(\nu)=M_{\mathbf{Y}}M_{\mathbf{Y}}^{\mathsf{T}}$.
-A result by Araki and Yamagami [3], later generalized by Kittaneh [39] and Bhatia [6], states for any two matrices $A$ and $B$ with the same shape, it holds that
+A result by Araki and Yamagami [Araki1981], later generalized by Kittaneh [MR0787884] and Bhatia [MR1275617], states for any two matrices $A$ and $B$ with the same shape, it holds that
 $\left\lVert\sqrt{A^{\mathsf{T}}A}-\sqrt{B^{\mathsf{T}}B}\right\rVert_{\rm F}\leqslant\sqrt{2}\left\lVert A-B\right\rVert_{\rm F},$
-see also [5, Theorem VII.5.7] for a textbook presentation. This yields
+see also [MR1477662, Theorem VII.5.7] for a textbook presentation. This yields
 $\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}$
@@
     $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)\bigr{\rVert}\leqslant KW_{p}(\mu,\delta_{0}).$ (3.9)
-    When $\mathcal{T}=\mathcal{M}_{\beta}$, this follows immediately from the inequality $\operatorname{e}^{-\beta f^{*}}\leqslant\operatorname{e}^{-\beta f(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})}\leqslant\operatorname{e}^{-\beta f_{*}}$ if $\ell=0$ and from Proposition A.3 if $\ell>0$, where $f^{*}$ the supremum of $f$ over ${}^{d}$.
+    When $\mathcal{T}=\mathcal{M}_{\beta}$, this follows immediately from the inequality $\operatorname{e}^{-\beta f^{*}}\leqslant\operatorname{e}^{-\beta f(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})}\leqslant\operatorname{e}^{-\beta f_{*}}$ if $\ell=0$ and from Proposition A.3 if $\ell>0$, where $f^{*}$ the supremum of $f$ over d.
 When $\mathcal{T}=\sqrt{\mathcal{C}_{\beta}}$, the inequality (3.9) follows in the same way after noticing that
@@
 Proof.
-The result can be proved using the same reasoning as in [10, Lemma 3.4] or [23, Lemma 15].
+The result can be proved using the same reasoning as in [carrillo2018analytical, Lemma 3.4] or [fornasier2021consensusbased, Lemma 15].
 For convenience, we denote by $C$ any constant depending only on $(f,\beta,\lambda,p,T)$.
@@
 $\forall x\in\mathbf{R}^{d},\qquad\forall\mu\in\mathcal{P}(\mathbf{R}^{d}),\qquad\bigl{\lvert}b(x,\mu)\bigr{\rvert}\vee\bigl{\lVert}\sigma(x,\mu)\bigr{\rVert}_{\rm F}\leqslant C\left(\lvert x\rvert^{p}+\int_{\mathbf{R}^{d}}|y|^{p}\,\mu(\mathrm{d}y)\right)^{\frac{1}{p}}.$
-Arguing in parallel for (a) and (b), we fix $j\in$ and apply the Burkholder–Davis–Gundy inequality to find for all $t\in[0,T]$ that
+Arguing in parallel for (a) and (b), we fix $j\in\llbracket 1,J\rrbracket$ and apply the Burkholder–Davis–Gundy inequality to find for all $t\in[0,T]$ that
 $\frac{1}{3^{p-1}}\mathbf{E}\left[\sup_{s\in[0,t]}\left\lvert X_{s}^{(j)}\right\rvert^{p}\right]$ $\leqslant\mathbf{E}\left\lvert X_{0}^{(j)}\right\rvert^{p}+T^{p-1}\int_{0}^{t}\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
@@
 3.3 Convergence of weighted moments for i.i.d. samples
-For the proof of the lemmas in this section, we will need [18, Theorem 1], which we recall for the reader’s convenience in Theorem 3.6 below, slightly adapted to the i.i.d.
+For the proof of the lemmas in this section, we will need [MR2597592, Theorem 1], which we recall for the reader’s convenience in Theorem 3.6 below, slightly adapted to the i.i.d.
 setting.
-We also refer to [1, Theorem 2.3], where similar results for $p=1$ and $p=2$ are proved in the context of importance sampling, with explicit constant prefactors.
+We also refer to [MR3696003, Theorem 2.3], where similar results for $p=1$ and $p=2$ are proved in the context of importance sampling, with explicit constant prefactors.
 Theorem 3.6 ($L^{p}$ convergence of weighted sums).
@@
 Consequently, the assumptions of Theorem 3.6 are satisfied, and we obtain the statement.
+∎
 Lemma 3.8 (Convergence of the weighted covariance for i.i.d. samples).
@@
 By assumption it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$.
-By an inequality due to van Hemmen and Ando [50], see also [5, Problem X.5.5], it holds that
+By an inequality due to van Hemmen and Ando [vanHemmen1980], see also [MR1477662, Problem X.5.5], it holds that
 $\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert_{\rm F}\leqslant\frac{1}{\eta}\Bigl{\lVert}\mathcal{C}_{\beta}(\overline{\mu}^{J})-\mathcal{C}_{\beta}(\mu)\Bigr{\rVert}_{\rm F}.$
@@
 Proof.
-The evolution of the covariance matrix is governed by [11, Equation (2.18)]
+The evolution of the covariance matrix is governed by [CBS-Carrillo2021, Equation (2.18)]
 $\partial_{t}\left(\mathcal{C}(\overline{\rho}_{t})\right)=-2\mathcal{C}(\overline{\rho}_{t})+2\lambda^{-1}\mathcal{C}_{\beta}(\overline{\rho}_{t}).$
@@
 For all $\mathbf{X}\in\mathbf{R}^{dJ}$ it holds that
-[…]
+$\lvert\mathcal{M}_{\beta}(\mathbf{X})\rvert\leqslant\lvert\mathbf{X}\rvert\qquad\text{and}\qquad\left\lVert\mathcal{C}_{\beta}(\mathbf{X})\right\rVert\leqslant\lvert\mathbf{X}\rvert^{2}.$ (4.1)
 Moreover, if $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ is locally Lipschitz continuous, then for each $R>0$ and $J\in\mathbf{N}^{+}$ there exists a constant $C_{J,R}>0$ such that
@@
 $\leqslant\left\lVert\sum_{j=1}^{J}w_{j}(\mathbf{X})\Bigl{(}X^{(j)}\otimes X^{(j)}\Bigr{)}\right\rVert_{\rm F}\leqslant\sum_{j=1}^{J}w_{j}(\mathbf{X})\left\lVert X^{(j)}\otimes X^{(j)}\right\rVert_{\rm F}=\sum_{j=1}^{J}w_{j}(\mathbf{X})\left\lvert X^{(j)}\right\rvert^{2}\leqslant\left\lvert\mathbf{X}\right\rvert^{2}.$
-The local Lipschitz estimate on $\mathcal{M}_{\beta}(\mathbf{X})$ in (4.2) is [10, Lemma 2.1].
+The local Lipschitz estimate on $\mathcal{M}_{\beta}(\mathbf{X})$ in (4.2) is [carrillo2018analytical, Lemma 2.1].
 In order to prove the local Lipschitz continuity of $\mathcal{C}_{\beta}(\mathbf{X})$, fix $R>0$ and $J>0$.
@@
 $\qquad+\sum_{j=1}^{J}\left\lvert\left(\sqrt{w_{j}(\mathbf{X})}-\sqrt{w_{j}(\mathbf{Y})}\right)\left(Y^{(j)}-\mathcal{M}_{\beta}(\mathbf{Y})\right)\right\rvert.$
-Since $\mathbf{X}\mapsto\mathcal{M}_{\beta}(\mathbf{X})$ is a locally Lipschitz continuous function by [10, Lemma 2.1], it suffices to prove that $\mathbf{X}\mapsto\sqrt{w_{j}(\mathbf{X})}$ is locally Lipschitz as well.
+Since $\mathbf{X}\mapsto\mathcal{M}_{\beta}(\mathbf{X})$ is a locally Lipschitz continuous function by [carrillo2018analytical, Lemma 2.1], it suffices to prove that $\mathbf{X}\mapsto\sqrt{w_{j}(\mathbf{X})}$ is locally Lipschitz as well.
 To this end, observe that
@@
 By Lemma 4.1, the functions $F$ and $G$ have sublinear growth and are locally Lipschitz continuous.
-Thus, by the non-explosion criterion from stochastic Lyapunov theory [38, Theorem 3.5], it suffices to find a function $\phi\in C^{2}\left(\mathbf{R}^{dJ},[0,\infty)\right)$ such that
+Thus, by the non-explosion criterion from stochastic Lyapunov theory [khasminskii2013stochastic, Theorem 3.5], it suffices to find a function $\phi\in C^{2}\left(\mathbf{R}^{dJ},[0,\infty)\right)$ such that
   • The function $\phi$ is coercive: $\phi(\mathbf{X})\to\infty$ as $|\mathbf{X}|\to\infty$ and
@@
 4.2 Well-posedness for the mean-field SDEs
-We adapt the argument from [10], using the more general Assumption H and the improved stability estimates from Subsection 3.1.
+We adapt the argument from [carrillo2018analytical], using the more general Assumption H and the improved stability estimates from Subsection 3.1.
 We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3.
@@
 Proof.
-The existence of a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$ to (4.3) follows from classical theory of stochastic differential equations [45, Theorem 5.2.1].
+The existence of a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$ to (4.3) follows from classical theory of stochastic differential equations [MR2001996, Theorem 5.2.1].
 It remains to prove the continuity statement for the weighted mean and covariance.
@@
 Proof of Theorem 2.4.
-Similarly to the proof of [10, Theorem 3.1], we first construct a map
+Similarly to the proof of [carrillo2018analytical, Theorem 3.1], we first construct a map
 $\mathcal{T}\colon C\left([0,T],\mathcal{X}\right)\rightarrow C\left([0,T],\mathcal{X}\right)$
@@
 Fixed points of $\mathcal{T}$ correspond to solutions of the McKean-Vlasov SDE (1.5).
-The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [27, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
+The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [gilbarg1977elliptic, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
 $\Bigl{\{}(m,\Gamma)\in C\bigl{(}[0,T],\mathcal{X}\bigr{)}:\exists\xi\in[0,1]\text{ such that }(m,\Gamma)=\xi\mathcal{T}\bigl{(}(m,\Gamma)\bigr{)}\Bigr{\}}.$ (4.6)
@@
 $\eta:=\inf\left\{\lambda_{\rm min}\Bigl{(}\mathcal{C}_{\beta}(\rho_{t})\Bigr{)}:t\in[0,T]\right\}>0.$
-By virtue of the van Hemmen–Ando inequality [50, Proposition 3.2], it holds for any unitarily invariant norm, thus in particular for the Frobenius norm, that
+By virtue of the van Hemmen–Ando inequality [vanHemmen1980, Proposition 3.2], it holds for any unitarily invariant norm, thus in particular for the Frobenius norm, that
 $\left\lVert\sqrt{\mathcal{C}_{\beta}(\rho_{r})}-\sqrt{\mathcal{C}_{\beta}(\widehat{\rho}_{r})}\right\rVert_{\rm F}\leqslant\eta^{-\frac{1}{2}}\left\lVert\mathcal{C}_{\beta}(\rho_{r})-\mathcal{C}_{\beta}(\widehat{\rho}_{r})\right\rVert_{\rm F}\,.$
@@
 which is indeed a weak formulation of the Fokker–Planck equation (2.5).
-For rigorous uniqueness and existence results for the Fokker–Planck equation (4.9) as well as a more precise discussion of its connection with (4.3), see for example [25].
+For rigorous uniqueness and existence results for the Fokker–Planck equation (4.9) as well as a more precise discussion of its connection with (4.3), see for example [MR2830391].
 ∎
@@
 By, among other things, replacing $\sqrt{2\lambda^{-1}\mathcal{R}(\Gamma_{t})}$ in (4.3) by $S\left(\overline{X}_{t}-\mathcal{M}_{\beta}(\overline{\rho}_{t})\right)$, the above proof can be easily adapted to provide a proof for Theorem 2.3.
-See also [10], where a weaker version of Theorem 2.3 was already proven in the case where $p=4$ and $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$.
+See also [carrillo2018analytical], where a weaker version of Theorem 2.3 was already proven in the case where $p=4$ and $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$.
 Remark 4.4.
@@
 First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.12.
-Then, in the following subsections, we discuss three natural extensions of our results.
-In Subsection 5.2, we show how our mean-field limit result for CBS (Theorem 2.12) can be improved when the initial condition $\overline{\rho}_{0}$ is Gaussian.
-In Subsection 5.3, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid.
-Finally, in Subsection 5.4, we show how Theorems 2.6 and 2.12 can be generalized to smaller values of $p\geqslant 2$.
+Then, in Subsection 5.2, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid.
 5.1 Proving mean-field results using stopping times
 In this section, we present an alternative approach for proving Theorem 2.6 and Theorem 2.12.
-This approach is based on stopping times, and inspired by a method employed in [33] to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
-A similar approach was employed in [36] to obtain a non-quantitative mean-field estimate.
+This approach is based on stopping times, and inspired by a method employed in [MR1949404] to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
+A similar approach was employed in [MR4553241] to obtain a non-quantitative mean-field estimate, and more recently in [vaes2024sharppropagationchaosensemble] to prove sharp propagation of chaos estimates for an ensemble sampler related to the ensemble Kalman filter [EKS-1-Hofmann-Stuart-Li-GI-2019, MR4123680].
 Before presenting the proof, we briefly discuss the main differences between this approach and the indicator set approach presented in Section 2.
@@
     $\Omega_{J}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\}.$ (5.1)
-    In certain contexts, however, such as the one discussed in Subsection 5.2, being able to exclude a time-dependent event is very useful.
+    In certain contexts, however, being able to exclude a time-dependent event could be very useful.
   • The stopping time approach presented hereafter does not enjoy this flexibility, but is convenient in settings where moments of both $\mu^{J}$ and $\overline{\mu}^{J}$ need to be controlled in order to apply the stability estimates obtained in Proposition 3.1 for the weighted moments.
-This is illustrated in Subsections 5.3 and 5.4.
+This is illustrated in Subsection 5.2.
 We illustrate the stopping times approach only for Theorem 2.12, noting that Theorem 2.6 can be proved in exactly the same manner.
@@
 $M_{t}=\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}.$
-By Doob’s optional stopping theorem [33, Theorem 3.3], see also [20, Equation 2.29, p.285], the stopped process $(M_{t\wedge\theta_{J}})_{t\geqslant 0}$ is a martingale, with a quadratic variation process given by the stopped quadratic variation process of $M_{t}$, i.e.
+By Doob’s optional stopping theorem [MR1949404, Theorem 3.3], see also [MR0838085, Equation 2.29, p.285], the stopped process $(M_{t\wedge\theta_{J}})_{t\geqslant 0}$ is a martingale, with a quadratic variation process given by the stopped quadratic variation process of $M_{t}$, i.e.
 by $(\langle M\rangle_{t\wedge\theta_{J}})_{t\geqslant 0}$.
@@
 We have by Lemma 3.5 and (2.6) that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\big{\rvert}^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}\big{\rvert}^{q}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}\overline{X}^{(t)}_{j}\big{\rvert}^{q}\right]\right)\leqslant 2^{q}\kappa.$
+$\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\big{\rvert}^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}\big{\rvert}^{q}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}\overline{X}^{(j)}_{t}\big{\rvert}^{q}\right]\right)\leqslant 2^{q}\kappa.$
 In addition, by Lemma 2.5 it holds that
@@
 We finally obtain from (5.3) that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
+$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
 which concludes the proof. ∎
-5.2 Improving the rate of convergence for Gaussian initial conditions.
-It is proved in [11] that, if the initial condition to the mean-field equation 1.5 associated with CBS is Gaussian, then at each time $t$ the solution $\overline{\rho}_{t}$ to the equation is also Gaussian.
-In this setting, the probability measure $\overline{\rho}_{0}$ has moments of all orders.
-Using a concentration inequality for Gaussian measures (see [43] and [42, Section 2]), it is possible to prove that the probability of the event
-$\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\lvert\overline{X}^{(j)}_{t}(\omega)\rvert^{p}\geqslant R\right\}$
-decreases in fact as $\exp\bigl{(}-CJ^{\frac{p}{2}}\bigr{)}$, for a constant $C$ depending on $R$ and on the parameters of the Gaussian distribution.
-Carrying out the proof of Theorem 2.6 with the sharper bound $\exp\bigl{(}-CJ^{\frac{p}{2}}\bigr{)}$ instead of $CJ^{-2}$ for $\mathbf{P}(\Omega_{J,t})$ as in Lemma 2.5, we obtain for all sufficiently large $p$ a mean-field estimate of the form
-$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
 5.2 Handling the case where $u>\ell>0$ in Assumption H
@@
     where we used that $\mathbf{P}[A+B\geqslant k]\leqslant\mathbf{P}[A\geqslant k/2]+\mathbf{P}[B\geqslant k/2]$ for any two real-valued random variables $A$ and $B$, because $\{A+B\geqslant k\}\subset\{A\geqslant k/2\}\cup\{B\geqslant k/2\}$.
-A similar trick, of bounding the probability of an event associated with a stochastic system in terms of a similar event associated with a simpler system and the discrepancy between the two systems, is employed in [19, Theorem 2.10].
+A similar trick, of bounding the probability of an event associated with a stochastic system in terms of a similar event associated with a simpler system and the discrepancy between the two systems, is employed in [2023arXiv230503637D, Theorem 2.10].
 The probability $\mathbf{P}[\theta_{J}\leqslant T]$ can then be bounded as follows:
@@
     $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{2,\frac{p}{2}\right\}\frac{q-p}{q}}.$
-5.4 Removing the lower bound on $p$
-In Theorems 2.6 and 2.12, we assume that $p$ is bounded from below by $p_{\mathcal{M}}$ and $2p_{\mathcal{M}}$, respectively.
-It is possible to replace this restriction by the less restrictive inequality $p\geqslant 2$ in the case of Theorem 2.6, and $p>2$ in the case of Theorem 2.12, provided that the initial condition $\overline{\rho}_{0}$ has sufficiently many moments.
-We motivate this statement hereafter for the specific case of CBO.
-Suppose for simplicity that $\overline{\rho}_{0}$ has finite moments of all orders.
-Then by Theorem 2.6, for $f\in\mathcal{A}(s,\ell,\ell)$ there exists an exponent $p_{\star}\geqslant 2$ such that for all $p\geqslant p_{\star}$ and all $\varepsilon>0$, it holds for some constant $C$ independent of $J$ that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-(2-\varepsilon)}.$ (5.8)
-  • First, note that since
-    $\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{r}\right]\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{r}{p}}$
-    for any $1\leqslant r\leqslant p$, the existence of a mean-field result of the type (5.8) with $p=p_{\star}$ immediately implies a similar result for all $r\in[2,p_{\star}]$, but with a rate that is degraded by a factor $\frac{r}{p_{\star}}$.
-In the next bullet point, we present a better approach.
-  • Fix $r\in[2,p_{\star}]$. From the proof of Proposition 3.1, it is easily observed that there is $\varrho$ sufficiently large to ensure that for all $R>0$, it holds that
-    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\times\mathcal{P}_{\varrho,R},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert+\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L\,W_{r}(\mu,\nu),$ (5.9)
-    for some constant $L=L(R,r)>0$, see Lemma A.6.
-In order to obtain a mean-field result for the exponent $r$, we can employ the same approach as in the proof in Subsection 5.1 with $p$ replaced by $r$, but with the stopping time
-    $\theta_{J}:=\overline{\tau}_{J}\wedge\tau_{J},\qquad\left\{\begin{aligned} \tau_{J}(R)&=\inf\Bigl{\{}t\geqslant 0:\mu_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}},\\ \overline{\tau}_{J}(R)&=\inf\Bigl{\{}t\geqslant 0:\overline{\mu}_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}}.\end{aligned}\right.$
-    for some $p\geqslant p_{\star}\vee\varrho$.
-The proof then goes through as before, except that we use (5.9) instead of Corollary 3.3, and that the probability $\mathbf{P}[\theta_{J}\leqslant T]$ is bounded from above differently.
-Reasoning as in (5.7), it is simple to show that
-    $\mathbf{P}[\theta_{J}\leqslant T]\leqslant 2\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\frac{2^{p}}{R}\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\right].$
-    The first term on the right-hand side decreases as $\mathcal{O}(J^{-2})$ by Lemma 2.5, whereas the second term decreases as $\mathcal{O}(J^{-(2-\varepsilon)})$ by (5.8). This enables to prove that
-    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(t)}_{j}\Big{\rvert}^{r}\right]\leqslant CJ^{-\min\left\{\frac{r}{2},2-\varepsilon\right\}}.$ (5.10)
 Appendix A Further technical results
@@
 Step 1
-Let us first generalize [10, Lemma 3.1] by showing that there exists a constant $\eta=\eta(p,R)>0$ such that
+Let us first generalize [carrillo2018analytical, Lemma 3.1] by showing that there exists a constant $\eta=\eta(p,R)>0$ such that
 $\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\int_{\mathbf{R}^{d}}h\,\mathrm{d}\mu\geqslant\eta.$ (A.2)
@@
 $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x).$
-for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved. The following is a generalization of [10, Lemma 3.3].
+for a constant $C$ depending only on $\beta$ and $f$.
+When $f$ is unbounded, proving an inequality of this type is more involved.
+The following is a generalization of [carrillo2018analytical, Lemma 3.3].
 Proposition A.3.
@@
 is a non-increasing function of $\gamma$.
-This follows from [11, Lemma A.3], and could also be proved by showing that $F^{\prime}(\gamma)$ is non-positive for all $\gamma\geqslant 0$.
+This follows from [CBS-Carrillo2021, Lemma A.3], and could also be proved by showing that $F^{\prime}(\gamma)$ is non-positive for all $\gamma\geqslant 0$.
 A simple calculation indeed gives that
@@
 FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM.
+[…]

### v3

@@
 For algorithms based on interacting particle systems that admit a mean-field description, convergence analysis is often more accessible at the mean-field level.
-In order to transpose convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics.
+In order to transfer convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics.
 In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling.
@@
 Keywords. Mean-field limits, Interacting particle systems, Consensus-Based Optimization, Consensus-Based Sampling, Coupling methods, Wasserstein stability estimates.
-AMS subject classification. 35Q93, 65C35, 70F45, 35K55.
+MSC codes. 35Q93, 65C35, 70F45, 35K55.
 1 Introduction
-This work focuses on the rigorous derivation of quantitative mean-field limits for recently proposed algorithms known as Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants.
+Particle methods provide a powerful paradigm for solving complex global optimization problems leading to highly parallelizable algorithms.
+Despite widespread and growing adoption, theory underpinning their behavior has been mainly based on meta-heuristics.
+In application settings involving black-box procedures, or where gradients are too costly to obtain, one relies on derivative-free approaches instead.
+This work focuses on the rigorous derivation of quantitative mean-field limits for the recently proposed derivative-free algorithms Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants.
 The former, initially proposed in [CBO], is an algorithm for global optimization tasks, while the latter, proposed in [CBS-Carrillo2021], can be employed for both sampling and optimization purposes.
-Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels.
+Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels, which makes it challenging to rigorously prove propagation of chaos.
+CBO and CBS methods are a novel family of algorithms providing an alternative to the popular particle swarm optimization (PSO) algorithm [Kennedy_Eberhart, Kennedy1997ThePS] that is found in many optimization toolboxes.
+Recent studies have shown that CBO is competitive with particle swarm optimization [totzeck2018numerical] for a number of benchmark functions, and able to solve high-dimensional optimization problems arising from machine learning [CBO-high-dim, MR4329816].
+By now, these methods have also been adapted for various other tasks and application settings, including finance [MR4337731], rare event simulation [2023arXiv230409077A] and federated learning [CBO-federal-carrillo2023], to mention just a few.
+Compared to PSO however, this family of algorithms is amenable to theoretical analysis.
+Since current theoretical guarantees are only available on the level of the mean-field description, providing quantitative estimates quantifying the distance between the evolution of the interacting particles and the corresponding mean-field probability distribution is essential.
 This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature.
@@
 In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical].
-The distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large exponential parameter $\beta$.
+While the consensus point does not necessarily coincide with the global minimizer of the cost function $f$, [carrillo2018analytical, Theorem 4.2] shows that the distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large parameter $\beta$.
 The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
@@
 As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [CBO-high-dim].
-Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [Ha2019, Ha2020ConvergenceAE, HA_Discrete_CBO_hetero_noise].
-(On this subject, we point out that the results in [Ha2020ConvergenceAE] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [jin2018random], memory effects [MR4160256], and Lévy noise [MR4553241], as well as extensions to multi-objective optimization [klamroth2022consensusbased, borghi2022adaptive] and global optimization with constraints [CBO-hypersurface-Fornasier_2020, MR4329816, CTV21, MR4538901, MR4404221, MR4337731].
+Using componentwise geometric Brownian motion in this way allows for proving consensus formation directly at the particle level, which was achieved in [Ha2019, Ha2020ConvergenceAE, HA_Discrete_CBO_hetero_noise].
+(On this subject, we point out that the results in [Ha2020ConvergenceAE] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [jin2018random], memory effects [MR4160256], and Lévy noise [MR4553241], as well as extensions to multi-objective optimization [klamroth2022consensusbased, borghi2022adaptive] and global optimization with constraints [MR4538901, carrillo2024interacting, CTV21, CBO-hypersurface-Fornasier_2020, beddrich2024constrained, MR4329816, CBO_sphere2, MR4337731].
 Moreover, a variant tailored for objective functions with several minima was proposed in [bungert], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [CBO-PSO-zero-inertia-1, CBO-PSO-zero-inertia-2].
@@
 In addition, it was noticed in [CBS-Carrillo2021] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [CBO].
-Since the convergence analysis in [CBS-Carrillo2021] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$.
+Since the convergence analysis in [CBS-Carrillo2021] is confined to the mean-field level, it is of interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$.
 This has so far been left as an open problem.
 Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [bungert], and for rare event estimation in [2023arXiv230409077A].
+It has also been shown that CBS, together with other sampling methods based on interacting particle systems, can be metropolized to correct the sampling error [sprungk2023metropolisadjustedinteractingparticlesampling].
 1.3 Propagation of chaos
 The CBO (1.1) and CBS (1.4) dynamics can both be recast as interacting particle systems of the form
-$\mathrm{d}X^{(j)}_{t}=b\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}t+\sigma\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}W^{(j)},\qquad j\in\llbracket 1,J\rrbracket.$ (1.6)
+$\mathrm{d}X^{(j)}_{t}=b\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}t+\sigma\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}W^{(j)},\qquad j\in\llbracket 1,J\rrbracket,$ (1.6)
 with drift coefficient $b$ and diffusion coefficient $\sigma$ depending on $\mu$ through a convolution with appropriate kernels $K_{b}$ and $K_{\sigma}$:
@@
 Specifically, it holds that $\mathbf{P}(\Omega_{M})=\mathcal{O}(M^{-1})$.
-  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [MR4553241].
-In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1.
+  • Shortly before completing this work, we found out that propagation of chaos for a variant of CBO had been proved using a coupling approach very recently in [MR4553241].
+In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1, which was developed independently.
 There are, however, several differences between our results and those presented in [MR4553241].
-Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [MR3696003, MR2597592] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3).
+Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [MR3696003, MR2597592] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a concentration inequality to bound the probability that empirical measures exit a compact set (Subsection 2.3).
 Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [MR4553241].
+  • Finally, we mention the recent preprint [koß2024meanfieldlimitconsensus], which appeared a few months after this one and proves propagation of chaos for consensus-based sampling using a compactness argument, similar to the approach used in [CBO-mfl-Huang2021].
 1.4 Our contributions
@@
 More precisely, we prove results with an explicit rate of convergence with respect to the number $J$ of particles.
-Investigating the roles of dimension and other problem parameters on the value of the constant prefactor is left for future work.
+Investigating the roles of the dimension and other problem parameters on the value of the constant prefactor is left for future work.
   • We present two approaches for dealing with the weighted mean (in the case of CBO) and the weighted covariance (in the case of CBS).
@@
 Theorem 2.3 (CBO: Existence and uniqueness for the mean-field SDE).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$.
-Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.3) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$.
-Furthermore, it holds that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty\qquad\text{and}\qquad\sup_{t\in[0,T]}\Bigl{\lvert}\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{\rvert}<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.4)
-and the function $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$.
+Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, and fix $x_{0}\sim\overline{\rho}_{0}$.
+Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.3) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$.
+Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.3) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$, and it holds that
+$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.4)
 Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C\left([0,T],\mathcal{P}_{p}(\mathbf{R}^{d})\right)$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
-$\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}\Bigr{)}+\frac{1}{2}\nabla\cdot\nabla\cdot\Bigl{(}D(\overline{\rho},x)\overline{\rho}\Bigr{)}.$ (2.5)
+$\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}\Bigr{)}+\frac{1}{2}\nabla\cdot\nabla\cdot\Bigl{(}D(\overline{\rho},x)\overline{\rho}\Bigr{)},$ (2.5)
 where $D(\overline{\rho},x):=S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}^{\mathsf{T}}$.
 Theorem 2.4 (CBS: Existence and uniqueness for the mean-field SDE).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$.
 Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$.
-Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.5) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$.
-Furthermore, it holds that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,\qquad\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\overline{\rho}_{t})\right\rvert<\infty,\qquad\sup_{t\in[0,T]}\left\lVert\mathcal{C}_{\beta}(\overline{\rho}_{t})\right\rVert<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.6)
-and the functions $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$, $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$.
+Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.5) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.5) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, and it holds that
+$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.6)
 Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C([0,T],\mathcal{P}_{p}(\mathbf{R}^{d}))$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
@@
 By the classical Marcinkiewicz–Zygmund inequality, it holds that
-$\mathbf{E}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}$ $\leqslant J^{-r}C_{\rm MZ}(r)\mathbf{E}\left[\left(\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{2}\right)^{\frac{r}{2}}\right]$
-$\leqslant J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{r}\right]=J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\Bigl{\lvert}Z_{1}-\mathbf{E}[Z_{1}]\Bigr{\rvert}^{r}\right],$
+$\mathbf{E}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}$ $\leqslant J^{-r}C_{\rm MZ}(r)\mathbf{E}\left[\left(\sum_{j=1}^{J}\Bigl{\lvert}Z_{j}-\mathbf{E}[Z_{j}]\Bigr{\rvert}^{2}\right)^{\frac{r}{2}}\right]$
+$\leqslant J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}Z_{j}-\mathbf{E}[Z_{j}]\Bigr{\rvert}^{r}\right]=J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\Bigl{\lvert}Z_{1}-\mathbf{E}[Z_{1}]\Bigr{\rvert}^{r}\right],$
 where we used the convexity of the function $x\mapsto x^{\frac{r}{2}}$ in the second inequality. Therefore, using the Markov inequality, we deduce that
@@
 from some $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$, we consider for each $J\in\mathbf{N}^{+}$ the interacting particle system
-$\forall j\in\llbracket 1,J\rrbracket,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
+$\forall j\in\llbracket 1,J\rrbracket,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}\,\mathrm{d}s+\int_{0}^{t}\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}\mathrm{d}W^{(j)}_{s},$ (2.8)
 to which we couple the system of i.i.d. mean-field McKean-Vlasov diffusions
-$\forall N\in\mathbf{N}^{+},\forall j\in\llbracket 1,J\rrbracket,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
+$\forall N\in\mathbf{N}^{+},\forall j\in\llbracket 1,J\rrbracket,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}s+\int_{0}^{t}\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
 driven by the same Brownian motions and initial conditions as (2.8). In the case of CBO, the drift and diffusion coefficients for (2.8) and (2.9) are given by
@@
 Remark 2.7.
-We prove below that the following sharper statement holds.
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$, and consider the systems (2.8) and (2.9) with the coefficients given by (2.10).
+We prove below the following more general statement: suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$.
 If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
@@
 Theorem 2.6 implies pathwise Wasserstein-$p$ chaos. More precisely, it holds that
-$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\frac{1}{2}}.$
+$W_{p}\Bigl{(}\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\Bigr{)}\leqslant CJ^{-\frac{1}{2}},$
 where $\rho^{J}_{[0,T]}$ is the pathwise law of the microscopic particle system over $[0,T]$ and $\overline{\rho}_{[0,T]}$ is the pathwise mean-field law over $[0,T]$. That is to say,
@@
 Proof of Theorem 2.6.
-As mentioned in Remark 2.7, we prove the sharp result (2.12).
+As mentioned in Remark 2.7, we prove the general result (2.12).
 It is sufficient to prove this statement for all $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ (by the assumption on $q$, this interval is nonempty).
-Indeed, if the statement holds true in this case, then by Jensen’s inequality, it holds for all $\mathfrak{p}\in(0,2\vee p_{\mathcal{M}}]$ that
-$\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{\mathfrak{p}}\right]\right)^{\frac{1}{\mathfrak{p}}}$ $\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2\vee p_{\mathcal{M}}}\right]\right)^{\frac{1}{2\vee p_{\mathcal{M}}}}$
-$\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}=CJ^{-\min\left\{\frac{1}{2},\frac{q-\mathfrak{p}}{2\mathfrak{p}^{2}},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}},$
+Indeed, if the statement holds true in this case, then by Jensen’s inequality, it holds for all $\bar{p}\in(0,2\vee p_{\mathcal{M}}]$ that
+$\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{\bar{p}}\right]\right)^{\frac{1}{\bar{p}}}$ $\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2\vee p_{\mathcal{M}}}\right]\right)^{\frac{1}{2\vee p_{\mathcal{M}}}}$
+$\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}=CJ^{-\min\left\{\frac{1}{2},\frac{q-\bar{p}}{2\bar{p}^{2}},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}},$
 where we used that the function $z\mapsto(q-z)/z^{2}$ is decreasing.
@@
 As in Sznitman’s argument [ReviewChaintronII, Theorem 3.1], we begin by writing
-$\bigl{\lvert}X_{t}^{(j)}-\overline{X}_{t}^{(j)}\bigr{\rvert}^{p}\leqslant$  $\,2^{p-1}\left\lvert\int_{0}^{t}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
-$+2^{p-1}\left\lvert\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$
+$\bigl{\lvert}X_{t}^{(j)}-\overline{X}_{t}^{(j)}\bigr{\rvert}^{p}\leqslant$  $\,2^{p-1}\left\lvert\int_{0}^{t}b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}s\right\rvert^{p}$
+$+2^{p-1}\left\lvert\int_{0}^{t}\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W_{s}\right\rvert^{p}.$
 By the Burkholder–Davis–Gundy inequality, see [MR2380366, Theorem 7.3], it holds for any $t\in[0,T]$ that
-$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant 2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert\,\mathrm{d}s\right)^{p}$
-$\qquad+C_{\rm BDG}2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
-$\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
-$\qquad+\,C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (2.13)
+$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant 2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert\,\mathrm{d}s\right)^{p}$
+$\qquad+C_{\rm BDG}2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lVert\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
+$\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$
+$\qquad+\,C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t}\left\lVert\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (2.13)
 By the triangle inequality, the first expectation on the right-hand side can be bounded as follows:
-$\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$ $\leqslant 2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\,\right\rvert^{p}\mathrm{d}s$
-$\qquad+2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s.$ (2.14)
+$\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$ $\leqslant 2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\,\right\rvert^{p}\mathrm{d}s$
+$\qquad+2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s.$ (2.14)
 Bounding the first term in (2.14)
 From (2.10) we have
-$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left\lvert\Bigl{(}\overline{X}^{(j)}_{s}-X^{(j)}_{s}+\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\Bigr{)}\right\rvert^{p}$
-$\leqslant 2^{p-1}\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+2^{p-1}\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}.$ (2.15)
+$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}$ $=\mathbf{E}\left\lvert\Bigl{(}\overline{X}^{(j)}_{s}-X^{(j)}_{s}+\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\Bigr{)}\right\rvert^{p}$
+$\leqslant 2^{p-1}\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+2^{p-1}\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}.$ (2.15)
 In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [ReviewChaintronII, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
@@
 We split the second term on the right-hand of (2.15) as follows:
-$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]$
-$\qquad+\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right].$ (2.17)
+$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}$ $=\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]$
+$\qquad+\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right].$ (2.17)
 Let us now bound each term on the right-hand side of (2.17) separately.
   • Since the $p$-th moment of $\overline{\mu}_{s}^{J}$ is bounded from above by $R$ for all $\omega\in\Omega\setminus\Omega_{J,s}$, we can use the stability estimate on the weighted mean given (see Item (a) and Corollary 3.3) to obtain
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\left(\mu^{J}_{s},\overline{\mu}^{J}_{s}\right)^{p}\right]$
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]$
     for the first term in (2.17). Since in addition
-    $\mathbf{E}\left[W_{p}\left(\mu^{J}_{s},\overline{\mu}^{J}_{s}\right)^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
+    $\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
     it follows that
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}.$
-  • We now bound the second term in (2.17). From Item (b) (see Lemma 3.5) and (2.4) we have
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right)\leqslant 2^{q}\kappa.$
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}.$
+  • We now bound the second term in (2.17). From Item (b) (see Lemma 3.5) as well as (2.4) and Proposition A.3 we have
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right)\leqslant 2^{q-1}\left(\kappa+\widetilde{\kappa}\right)$
+    where $\widetilde{\kappa}:=C_{2}\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{q}\right]<\infty$.
+Here, $C_{2}>0$ is the constant from Proposition A.3.
 Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{\frac{(q-1)p}{q}}\left(\kappa+\widetilde{\kappa}\right)^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
 We have thus shown that
-$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}},$
+$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}},$
 which in turn by (2.15) implies that
-$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
+$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
 Bounding the second term in (2.14)
 For the second term on the right-hand side of (2.14), we have by Item (c) (see Lemma 3.7) that
-$\mathbf{E}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}=\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\rho}_{s}\right)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}}.$
+$\mathbf{E}\left\lvert b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}=\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\rho}_{s}\bigr{)}\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}}.$
 Conclusion
 Combining the bounds on both terms in (2.14), we deduce
-$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
+$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
 Using that $S\colon\mathbf{R}\rightarrow\mathbf{R}^{d\times d}$ is globally Lipschitz continuous, we obtain analogously
-$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
+$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
 Substituting in (2.13), we finally obtain
 $\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
-The conclusion then follows from Grönwall’s lemma, since $\frac{q-p}{2p^{2}}\leqslant\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}$ for the value of $p$ considered. ∎
+The conclusion follows from Grönwall’s lemma, since $\frac{q-p}{2p^{2}}\leqslant\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}$ for the value of $p$ considered. ∎
 Remark 2.9.
 The proof is similar to that of McKean’s theorem [ReviewChaintronII, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
-Remark 2.10.
-A crucial detail in the proof is that Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below.
+A crucial detail here is the following: Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below.
 This allows us to define $\Omega_{J,t}$ in (2.16) only in terms of the i.i.d.
-mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ need not be included in the definition of $\Omega_{J,t}$
-Remark 2.11.
-In his famous lecture notes [MR1108185], Sznitman shows for a laboratory example that
-$\forall j\in\llbracket 1,J\rrbracket,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
-More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [MR1476912].
-Although CBO does not fit within Sznitman’s simplified framework, Item (c) enables to show a bound with the same rate for the second term on the right-hand side of (2.14),
-$\mathbf{E}\biggl{[}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,t}}\biggr{]}\leqslant\frac{C}{J^{\frac{p}{2}}}.$
-The limiting factor, for the overall rate of convergence in Theorem 2.6, is the bound on the probability $\mathbf{P}[\Omega_{J,t}]$.
+mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ does not need to be included in the definition of $\Omega_{J,t}$.
+This trick allows to show the strong mean-field limit result (2.12), improving on the previous results from [CBO-mfl-Huang2021, CBO-hypersurface-Fornasier_2020, fornasier2021consensusbased].
 2.5 Quantitative mean-field limit for CBS
-Using the same reasoning as that in the proof of Theorem 2.6, we can prove a mean-field result for the CBS dynamics (1.4).
-We state the theorem and, given the similarity of the proof with that of Theorem 2.6, only list the key ingredients needed for the proof.
+Proving a mean-field limit for the CBS dynamics (1.4) is more challenging than for CBO because of the square root of the weighted covariance in the diffusion coefficient.
+Thanks to our novel stability estimate for the square root of the weighted covariance (see Corollary 3.3 below), we can prove a mean-field result for the CBS dynamics (1.4) in the same way as for CBO.
 Theorem 2.10 (Mean-field limit for CBS).
@@
 Remark 2.11.
-In the same way as for CBO, it is possible to prove the following sharper statement.
+In the same way as for CBO, it is possible to prove the following more general statement.
 Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$.
-Consider the system (2.8) and (2.9) with the coefficients given by (2.18).
 If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
@@
   • Wasserstein stability estimate for the weighted covariance.
+The crucial ingredient for the proof of Theorem 2.10 is a stability estimate for the square root of the weighted covariance.
 We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant 2p_{\mathcal{M}}(\ell,s)$, there exists $\widetilde{L}_{\sqrt{\mathcal{C}}}(R,p)>0$ such that
@@
     $\eta:=\inf_{t\in[0,T]}\lambda_{\text{min}}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$
+With these ingredients at hand, we can prove Theorem 2.10 in the same way as for CBO, so we skip the proof here.
+Let us mention, however, that a different proof based on a technique using stopping times is presented in Subsection 5.1.
 3 Auxiliary results
@@
 In Subsection 3.1, we improve on existing Wasserstein stability estimates for the weighted mean and prove novel ones for the weighted covariance and the square root of the weighted covariance.
-In Subsection 3.2, we present moment estimates for the empirical measures associated with the CBO (1.1) and CBS (1.4) dynamics.
+Then, in Subsection 3.2, we present moment estimates for the empirical measures associated with the CBO (1.1) and CBS (1.4) dynamics.
 In Subsection 3.3, we prove convergence estimates for the weighted moments when particle positions are drawn independently from a given probability distribution, using results from the statistics literature.
@@
 In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of 3.11 would be bounded from above uniformly in $i$ by (3.8).
-Therefore $W_{p}(\nu_{i},\delta_{0})\to\infty$.
+Therefore, $W_{p}(\nu_{i},\delta_{0})\to\infty$.
 On the other hand, by the triangle inequality,
@@
 Proof.
-By assumption it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$.
+By assumption, it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$.
 By an inequality due to van Hemmen and Ando [vanHemmen1980], see also [MR1477662, Problem X.5.5], it holds that
@@
 In order to deal with the matrix square root in the CBS dynamics, we prove that the weighted covariance stays positive definite for finite times.
+This is particularly relevant when CBS is run in optimization mode with $\lambda=1$, since then $\mathcal{C}(\overline{\rho}_{t})\to 0$ as $t\to\infty$ under appropriate assumptions [CBS-Carrillo2021].
 Proposition 3.9 (No collapse in finite time).
@@
 $\forall t\in[0,T],\qquad\mathcal{C}(\overline{\rho}_{t})\succcurlyeq\mathcal{C}(\overline{\rho}_{0})\operatorname{e}^{-2t}.$ (3.13)
-In particular, if $\mathcal{C}(\overline{\rho}_{0})\succ 0$ and if $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ is continuous, then
+In particular, if $\mathcal{C}(\overline{\rho}_{0})\succ 0$ and if $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ is continuous (which is true under the assumptions of Theorem 2.4), then
 $\inf_{t\in[0,T]}\lambda_{\rm min}\bigl{(}\mathcal{C}(\overline{\rho}_{t})\bigr{)}>0$ and $\inf_{t\in[0,T]}\lambda_{\rm min}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$ (3.14)
@@
 $\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\operatorname{e}^{-\beta f(x)}\,\overline{\rho}_{t}(\mathrm{d}x)=0.$
-Thus we obtain $y^{\mathsf{T}}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\bigr{)}=0$ for $\overline{\rho}_{t}$-almost every $x\in\mathbf{R}^{d}$, and so $y^{\mathsf{T}}\mathcal{M}_{\beta}(\overline{\rho}_{t})=y^{\mathsf{T}}\mathcal{M}(\overline{\rho}_{t})$ by integrating against $\overline{\rho}_{t}$.
+Thus, we obtain $y^{\mathsf{T}}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\bigr{)}=0$ for $\overline{\rho}_{t}$-almost every $x\in\mathbf{R}^{d}$, and so $y^{\mathsf{T}}\mathcal{M}_{\beta}(\overline{\rho}_{t})=y^{\mathsf{T}}\mathcal{M}(\overline{\rho}_{t})$ by integrating against $\overline{\rho}_{t}$.
 But then,
-$y^{\mathsf{T}}\mathcal{C}(\overline{\rho}_{t})y=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=0.$
-which is a contradiction. Thus $C_{\beta}(\overline{\rho}_{t})\succ 0$ for all $t\in[0,T]$, which implies the second inequality in (3.14). ∎
+$y^{\mathsf{T}}\mathcal{C}(\overline{\rho}_{t})y=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=0,$
+which is a contradiction. Thus, $C_{\beta}(\overline{\rho}_{t})\succ 0$ for all $t\in[0,T]$, which implies the second inequality in (3.14). ∎
 4 Proof of the well-posedness results
@@
 This section is organized as follows.
-First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.12.
-Then, in Subsection 5.2, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid.
+First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.10.
+Then, in Subsection 5.2, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, in which the improved stability estimates in Corollary 3.3 are no longer valid.
 5.1 Proving mean-field results using stopping times
@@
 We illustrate the stopping times approach only for Theorem 2.10, noting that Theorem 2.6 can be proved in exactly the same manner.
+Furthermore, we assume for simplicity that $\overline{\rho}_{0}$ has infinitely many moments.
 Alternative proof of Theorem 2.10.
-Let $R>0$ be such that the stopping time
-$\theta_{J}(R):=\inf\Bigl{\{}t\geqslant 0:\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R\Bigr{\}},$ (5.2)
-satisfies $\mathbf{P}[\theta_{J}\leqslant T]=\mathcal{O}(J^{-2}).$ Note that the event $\{\theta_{J}\leqslant T\}$ coincides with the event $\Omega_{J}$ given in (5.1).
-From Hölder’s inequality it follows that
+For $R>0$, let $\theta_{J}(R)$ be the stopping time
+$\theta_{J}(R):=\inf\Bigl{\{}t\geqslant 0:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\Bigr{\}},$ (5.2)
+Note that the event $\{\theta_{J}\leqslant T\}$ coincides with the event $\Omega_{J}$ given in (5.1). By Hölder’s inequality, it holds for all $q>p$ that
 $\frac{1}{2^{p-1}}\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]$
@@
 We again employ an adapted version of Sznitman’s argument. We have
-$\left\lvert X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\right\rvert^{p}\leqslant$ $\left\lvert\int_{0}^{t\wedge\theta_{J}}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
-$+\left\lvert\int_{0}^{t\wedge\theta_{J}}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$
+$\frac{1}{2^{p-1}}\left\lvert X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\right\rvert^{p}\leqslant$ $\left\lvert\int_{0}^{t\wedge\theta_{J}}b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}s\right\rvert^{p}$
+$+\left\lvert\int_{0}^{t\wedge\theta_{J}}\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W_{s}\right\rvert^{p}.$
 Let us introduce the martingale
-$M_{t}=\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}.$
+$M_{t}=\int_{0}^{t}\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W_{s}.$
 By Doob’s optional stopping theorem [MR1949404, Theorem 3.3], see also [MR0838085, Equation 2.29, p.285], the stopped process $(M_{t\wedge\theta_{J}})_{t\geqslant 0}$ is a martingale, with a quadratic variation process given by the stopped quadratic variation process of $M_{t}$, i.e.
@@
 Therefore, by the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$ that
-$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
-$\qquad+C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.4)
-From the triangle inequality we obtain that
-$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lvert b\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-b\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rvert^{p}\,\mathrm{d}s$
-$\qquad+\int_{0}^{t}\mathbf{E}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s.$ (5.5)
+$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\bigr{\rvert}^{p}\right]$ $\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$
+$\qquad+C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.4)
+Applying the triangle inequality gives that
+$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\bigl{\lvert}b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\bigr{\rvert}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lvert b\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$
+$\qquad+\int_{0}^{t}\mathbf{E}\left\lvert b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s.$ (5.5)
 Similarly, for the diffusion term, we have
-$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$
-$\qquad+\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.6)
+$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lVert\sigma\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s$
+$\qquad+\int_{0}^{t}\mathbf{E}\left\lVert\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.6)
 Next, we need to bound the terms on the right-hand side of (5.5) and (5.6).
@@
   • For the first diffusion term, we have that
-    $\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}$ $=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\biggl{\lVert}\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{s\wedge\theta_{J}}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)}\biggr{\rVert}_{\rm F}^{p}.$
+    $\left\lVert\sigma\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rVert_{\rm F}^{p}$ $=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\biggl{\lVert}\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{s\wedge\theta_{J}}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)}\biggr{\rVert}_{\rm F}^{p}.$
     Using Corollary 3.3, together with the fact
-    $\mathbf{E}\left[W_{p}\left(\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\right\rvert^{p},$
+    $\mathbf{E}\Bigl{[}W_{p}\bigl{(}\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\bigr{)}^{p}\Bigr{]}\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{k=1}^{J}\left\lvert X^{(k)}_{s\wedge\theta_{J}}-\overline{X}^{(k)}_{s\wedge\theta_{J}}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p},$
     we deduce that
-    $\mathbf{E}\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p}.$
+    $\mathbf{E}\left\lVert\sigma\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rVert_{\rm F}^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p}.$
   • For the other diffusion term, we use Lemma 3.8.
-The assumptions of this lemma are satisfied since, on the one hand, $\overline{\rho}_{s}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ by Theorem 2.4 and $2p<q$, and on the other hand, it holds that $\mathcal{C}_{\beta}(\overline{\rho}_{s})\succ 0$ by virtue of Proposition 3.9.
+The assumptions of this lemma are satisfied, since it holds that $\overline{\rho}_{s}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ for all $q>0$ by Theorem 2.4, and additionally $\mathcal{C}_{\beta}(\overline{\rho}_{s})\succ 0$ by Proposition 3.9.
 We deduce that
-    $\mathbf{E}\left\lVert\sigma\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\rho}_{s}\right)}\right\rVert_{\rm F}^{p}\leqslant CJ^{-\frac{p}{2}}.$
+    $\mathbf{E}\left\lVert\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\rho}_{s}\right)}\right\rVert_{\rm F}^{p}\leqslant CJ^{-\frac{p}{2}}.$
 Combining the bounds on both terms, we deduce that
-$\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u\wedge\theta_{J}}-\overline{X}^{(j)}_{u\wedge\theta_{J}}\right\rvert^{p}\right]\,\mathrm{d}s.$
-Applying a similar reasoning to the drift terms, we finally obtain from (5.4) that
-$\mathbf{E}\left[\sup_{s\in[0,t]}\Big{\lvert}X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\bigl{\lvert}X_{u\wedge\theta_{J}}^{(j)}-\overline{X}_{u\wedge\theta_{J}}^{(j)}\bigr{\rvert}^{p}\right]\,\mathrm{d}s.$
+$\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u\wedge\theta_{J}}-\overline{X}^{(j)}_{u\wedge\theta_{J}}\right\rvert^{p}\right]\,\mathrm{d}s.$
+Applying a similar reasoning for the drift terms, we finally obtain from (5.4) that
+$\mathbf{E}\left[\sup_{s\in[0,t]}\Big{\lvert}X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\Bigl{\lvert}X_{u\wedge\theta_{J}}^{(j)}-\overline{X}_{u\wedge\theta_{J}}^{(j)}\Bigr{\rvert}^{p}\right]\,\mathrm{d}s.$
 By Grönwall’s inequality, this implies that
@@
 $\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\big{\rvert}^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}X^{(j)}_{t}\big{\rvert}^{q}\right]+\mathbf{E}\left[\sup_{t\in[0,T]}\big{\lvert}\overline{X}^{(j)}_{t}\big{\rvert}^{q}\right]\right)\leqslant 2^{q}\kappa.$
-In addition, by Lemma 2.5 it holds that
-$\mathbf{P}[\theta_{J}(R)\leqslant T]=\mathbf{P}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\right]\leqslant CJ^{-2}.$
+In addition, by Lemma 2.5, it holds for any sufficiently large fixed $R$, that for all $\alpha>0$ there is $C_{\alpha}$ such that
+$\mathbf{P}[\theta_{J}(R)\leqslant T]=\mathbf{P}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\right]\leqslant C_{\alpha}J^{-\alpha}.$
 Conclusion
 We finally obtain from (5.3) that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
-which concludes the proof. ∎
+$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{\alpha(q-p)}{q}}\right),$
+which concludes the proof since $q$ and $\alpha$ can be taken arbitrarily large. ∎
 5.2 Handling the case where $u>\ell>0$ in Assumption H
 Our main results stand only when $f\in\mathcal{A}(s,\ell,u)$ with $u=\ell$.
-This is a convenient setting because it then holds by Proposition A.3 that the $p$-th moment of the reweighted measure $L_{\beta}\mu$ is controlled by the $p$-th moment of the $\mu$, which is key in order to prove well-posedness of the mean-field system, to obtain moment estimates for empirical measures, and, most importantly, to improve the Wasserstein stability estimates for the weighted mean and weighted covariance, see Corollary 3.3.
+This is a convenient setting because it then holds by Proposition A.3 that the $p$-th moment of the reweighted measure $L_{\beta}\mu$ is controlled by the $p$-th moment of $\mu$, which is key in order to prove well-posedness of the mean-field system, to obtain moment estimates for empirical measures, and, most importantly, to improve the Wasserstein stability estimates for the weighted mean and weighted covariance, see Corollary 3.3.
 In the interest of simplicity, we refrained from stating precise results for the cases $u>\ell=0$ and $u>\ell>0$.
-The former remains open, but we note that a quantitative mean-field estimate can still be obtained in the latter setting.
+We do not address the former setting, but note that a quantitative mean-field estimate can still be obtained in the latter.
 The main ingredients to this end are the following:
   • Moment estimates for the empirical measure and well-posedness of the mean-field equation can be obtained by using Proposition A.3 together with the Bihari–LaSalle inequality instead of Grönwall’s inequality, using the same reasoning as in Lemma 3.5 and Section 4, respectively.
-In contrast with the setting $\ell=u$, these results hold only up to some finite time $T<T_{+}$, where $T_{+}\in(0,\infty)$ is the time at which the bounds through the Bihari–LaSalle inequality become singular.
-  • In order to obtain a mean-field result up to some time $T<T_{+}$, the approach presented in Subsection 5.1 can be used, this time with the stopping time $\theta_{J}:=\tau_{J}\wedge\overline{\tau}_{J}$, where
+In contrast with the setting $\ell=u$, these results hold only up to some finite time $T<T_{+}$, where $T_{+}\in(0,\infty)$ is the time at which the bounds given by the Bihari–LaSalle inequality become singular.
+  • In order to obtain a mean-field result up to some time $T<T_{+}$, the approach presented in Subsection 5.1 can be used, but with the stopping time $\theta_{J}:=\tau_{J}\wedge\overline{\tau}_{J}$, where
     $\tau_{J}(R)$ $=\inf\Bigl{\{}t\geqslant 0:\mu_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}},$
     $\overline{\tau}_{J}(R)$ $=\inf\Bigl{\{}t\geqslant 0:\overline{\mu}_{t}^{J}\bigl{[}\lvert x\rvert^{p}\bigr{]}\geqslant R\Bigr{\}}.$
-    Employing this stopping time instead of (5.2) allows for using the basic stability estimates for the weighted moments of Proposition 3.1, which are valid even when $\ell<u$, instead of the improved stability estimates in Corollary 3.3, which require that $\ell=u$.
+    Employing this stopping time instead of (5.2) enables to use the basic stability estimates for the weighted moments of Proposition 3.1, which are valid even when $\ell<u$, instead of the improved stability estimates in Corollary 3.3, which require that $\ell=u$.
+It is then simple to obtain a propagation of chaos estimate for the stopped particle system, of the form
+    $\mathbf{E}\left[\sup_{t\in[0,T]}\Bigl{\lvert}X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\Bigr{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
     In order to complete the proof as in Subsection 5.1, it remains to bound the probability $\mathbf{P}\left[\theta_{J}\leqslant T\right]$, which can be achieved by noticing that
@@
     $\mathbf{P}[\theta_{J}(R)\leqslant T]\leqslant 2\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}\overline{X}^{(j)}_{t}\Bigr{\rvert}^{p}\geqslant\frac{R}{2^{p}}\right]+\frac{2^{p}}{R}\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Bigl{\lvert}X^{(j)}_{t\wedge\theta_{J}}-\overline{X}^{(j)}_{t\wedge\theta_{J}}\Bigr{\rvert}^{p}\right].$ (5.7)
-    The first term is bounded from above by $CJ^{-2}$ by Lemma 2.5, whereas the second term can be bounded by $CJ^{-\frac{p}{2}}$ as in the proof in Subsection 5.1.
-Thus, the probability $\mathbf{P}[\theta_{J}(R)\leqslant T]$ scales as $J^{-\min\{2,\frac{p}{2}\}}$, instead of $J^{-2}$ in the case where $u=\ell$.
+    By Lemma 2.5, For any $\alpha>0$ there exists $C_{\alpha}>0$ such that the first term is bounded from above by $C_{\alpha}J^{-\alpha}$ for all $J\in\mathbf{N}^{+}$, whereas the second term can be bounded by $CJ^{-\frac{p}{2}}$ as in the proof presented in Subsection 5.1.
+Thus, the probability $\mathbf{P}[\theta_{J}(R)\leqslant T]$ scales as $J^{-\min\left\{\alpha,\frac{p}{2}\right\}}$, instead of $J^{-\alpha}$ in the case where $u=\ell$.
 As a consequence, we obtain a mean-field result of the type
-    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{2,\frac{p}{2}\right\}\frac{q-p}{q}}.$
+    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\frac{p}{2}\left(\frac{q-p}{q}\right)}.$
+    Since $q$ was arbitrary given the assumption that $\overline{\rho}_{0}$ has infinitely many moments, it follows that for every $\varepsilon>0$, there exists $C_{\varepsilon}$ such that
+    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\frac{p}{2}\left(1-\varepsilon\right)}.$
+    Hence, we recover almost the classical Monte Carlo rate.
+By using the same method as in [vaes2024sharppropagationchaosensemble], it is in fact possible to prove that this statement holds also for $\varepsilon=0$, but for conciseness we do not present the details here.
 Appendix A Further technical results
@@
 $\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\mu\left(B_{K}\right)\geqslant 1-\mu\left(\mathbf{R}^{d}\setminus B_{K}\right)\geqslant 1-\frac{R^{p}}{K^{p}}=\frac{1}{2}.$
-Therefore we obtain
+Therefore, we obtain
 $\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\int_{\mathbf{R}^{d}}h\,\mathrm{d}\mu\geqslant\int_{B_{K}}h\,\mathrm{d}\mu\geqslant\frac{1}{2}\min\bigl{\{}h(x):x\in B_{K}\bigr{\}}=:\eta>0,$
@@
 $\left\lVert\frac{g(x)\mathcal{S}(\mu)}{\sqrt{\mu[h]}}-\frac{g(y)\mathcal{S}(\nu)}{\sqrt{\nu[h]}}\right\rVert_{L^{2}(\pi)}$ $\leqslant\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\Bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\Bigr{)}\right\rVert_{L^{2}(\pi)}+\left\lVert\Bigl{(}g(x)-g(y)\Bigr{)}\frac{\mathcal{S}(\nu)}{\sqrt{\mu[h]}}\right\rVert_{L^{2}(\pi)}$
-$\quad+\left\lVert{g(y)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)}.$
+$\quad+\left\lVert{g(y)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)},$
 where, by a slight abuse of notation, we wrote $g(x)$ and $g(y)$ in order to refer to the functions $(x,y)\mapsto g(x)$ and $(x,y)\mapsto g(y)$, respectively.
@@
 where we used that $\frac{2\varsigma p}{p-2}\leqslant p$ since $p\geqslant 2+2\varsigma$. In addition, from (b) we have
-$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R.$ (A.4)
+$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R,$ (A.4)
 and so we conclude that
@@
 If the function $f$ is bounded from below and above, it is trivial to show that
-$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x).$
+$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x)$
 for a constant $C$ depending only on $\beta$ and $f$.

### v4

@@
 CBO and CBS methods are a novel family of algorithms providing an alternative to the popular particle swarm optimization (PSO) algorithm [Kennedy_Eberhart, Kennedy1997ThePS] that is found in many optimization toolboxes.
-Recent studies have shown that CBO is competitive with particle swarm optimization [totzeck2018numerical] for a number of benchmark functions, and able to solve high-dimensional optimization problems arising from machine learning [CBO-high-dim, MR4329816].
-By now, these methods have also been adapted for various other tasks and application settings, including finance [MR4337731], rare event simulation [2023arXiv230409077A] and federated learning [CBO-federal-carrillo2023], to mention just a few.
-Compared to PSO however, this family of algorithms is amenable to theoretical analysis.
-Since current theoretical guarantees are only available on the level of the mean-field description, providing quantitative estimates quantifying the distance between the evolution of the interacting particles and the corresponding mean-field probability distribution is essential.
-This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature.
+Recent studies have shown that CBO is competitive with PSO for a number of benchmark functions [totzeck2018numerical], and able to solve high-dimensional optimization problems arising from machine learning [CBO-high-dim, MR4329816].
+For a detailed comparison between CBO and PSO, and additional context on the connections between the two methods, we refer the reader to the articles [CBO-PSO-zero-inertia-1, grassi2023mean, huang2023global].
+By now, CBO and CBS have also been adapted for various other tasks and application settings, including finance [MR4337731], rare event simulation [2023arXiv230409077A] and federated learning [CBO-federal-carrillo2023], to mention just a few.
+Compared to PSO however, this family of algorithms is more amenable to theoretical analysis.
+Since current theoretical guarantees for CBO and CBS are only available at the level of the mean-field description, providing quantitative estimates for the distance between the evolution of the interacting particles and the corresponding mean-field probability distribution is essential.
+The rest of this introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature.
 In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients.
@@
 1.1 Consensus-Based Optimization (CBO)
-Given an objective function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, we consider the following version of Consensus-Based Optimization (CBO):
+Given an objective function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ to be minimized, we consider the following version of Consensus-Based Optimization (CBO):
 $\mathrm{d}X^{(j)}_{t}$ $=-\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}t+S\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}W^{(j)}_{t},\qquad t\geqslant 0,\qquad j\in\llbracket 1,J\rrbracket,$ (1.1)
@@
 Its precise definition, together with its expression when $\mu=\mu^{J}_{t}$, is given by
-$\mathcal{M}_{\beta}(\mu):=\frac{\int_{\mathbf{R}^{d}}x\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)},\qquad\mathcal{M}_{\beta}(\mu^{J}_{t}):=\frac{\sum_{j=1}^{J}X^{(j)}\operatorname{e}^{-\beta f\left(X^{(j)}_{t}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}_{t}\right)}}.$ (1.2)
+$\mathcal{M}_{\beta}(\mu):=\frac{\int_{\mathbf{R}^{d}}x\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)},\qquad\mathcal{M}_{\beta}(\mu^{J}_{t}):=\frac{\sum_{j=1}^{J}X^{(j)}\operatorname{e}^{-\beta f\bigl{(}X^{(j)}_{t}\bigr{)}}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\bigl{(}X^{(k)}_{t}\bigr{)}}}.$ (1.2)
 The weighted mean $\mathcal{M}_{\beta}(\mu^{J}_{t})$ may be viewed as an approximation of the particle position $X^{(j)}_{t}$ such that $f(X^{(j)}_{t})$ is minimal.
@@
 Commonly a constant coefficient is included in front of the drift term in (1.1), but since this coefficient can be removed by rescaling time upon redefining the diffusion parameter $\theta$, we choose to omit it.
-By design, CBO allows passage to the corresponding mean-field description.
+By design, CBO allows passage to the corresponding mean-field description, which is not the case for the PSO algorithm.
 Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [PSO], it is possible to formally write down the corresponding mean-field equation:
 […]
-where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [CBO-PSO-zero-inertia-1].
+where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ‘current best’ that appears in PSO [CBO-PSO-zero-inertia-1].
 The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased].
@@
 We then prove quantitative mean-field results for Consensus-Based Optimization and Sampling in Subsections 2.4 and 2.5, respectively.
+Finally, a numerical experiment illustrating the convergence as $J\to\infty$ for CBO is presented in Subsection 2.6.
 2.1 Main assumptions
@@
   2. (H2)
-    There exist constants $c_{\ell},c_{u}>0$ and $u\geqslant\ell\geqslant 0$ such that
-    $\forall x\in^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
-    $\forall x\in^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
+    There exist constants $c_{\ell},C_{\ell},c_{u},C_{u}>0$ and $u\geqslant\ell\geqslant 0$ such that
+    $\forall x\in^{d},\qquad f(x)-f_{\star}\leqslant c_{u}|x|^{u}+C_{u},$ (2.2a)
+    $\forall x\in^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}|x|^{\ell}-C_{\ell}.$ (2.2b)
+In the rest of this paper, we denote by $\mathcal{A}(s,\ell,u)$ the set of objective functions that satisfy Assumption H with these parameters.
 Remark 2.1.
 A few remarks are in order.
-  • When $\ell=u=0$, assumption (H2) implies that $f$ is bounded from above.
-  • Observe that (H1) implies (2.2a) with $u=s+1$.
+  • The case $\ell=u=0$ corresponds to bounded $f$.
+  • If $f\in\mathcal{A}(s,\ell,u)$, then necessarily $\ell\leqslant s+1$, because the set $\mathcal{A}(s,\ell,u)$ is empty for $\ell>s+1$.
   • In the particular case where $s=1$ and either $u=\ell=0$ or $u=\ell=2$, the assumptions in Assumption H coincide with those in [carrillo2018analytical].
 The well-posedness and mean-field results proved in this paper generalize beyond this setting, as explained below.
-In the rest of this paper, we denote by $\mathcal{A}(s,\ell,u)$ the set of objective functions that satisfy Assumption H with these parameters.
-Our main results hold only for $f\in\mathcal{A}(s,\ell,\ell)$, i.e.
+Our main results hold for $f\in\mathcal{A}(s,\ell,\ell)$, i.e.
 under Assumption H with $\ell=u$.
+The problem of proving mean-field limits for CBO/CBS in the case $\ell<u$ is discussed in Subsection 5.2.
 Many of the results we present hold only for values of $p$ in a range dictated by the interval of validity of the Wasserstein stability estimates presented in Subsection 3.1.
 In order to avoid repetitions, we introduce the associated notation
-$p_{\mathcal{M}}(s,\ell):=\begin{cases}s+2,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$},\end{cases}\qquad p_{\mathcal{C}}(s,\ell):=\begin{cases}s+3,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$}.\end{cases}$ (2.3)
+$p_{\mathcal{M}}(s,\ell)\coloneq\begin{cases}s+2,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$},\end{cases}\qquad p_{\mathcal{C}}(s,\ell)\coloneq\begin{cases}s+3,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$}.\end{cases}$ (2.3)
 2.2 Well-posedness for CBO and CBS
@@
 We now present well-posedness results for the mean-field SDEs, beginning with CBO.
-For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [carrillo2018analytical, Theorems 3.1 and 3.2].
+For $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$, the following result generalizes [carrillo2018analytical, Theorems 3.1 and 3.2].
 Theorem 2.3 (CBO: Existence and uniqueness for the mean-field SDE).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and fix a final time $T>0$.
 Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, and fix $x_{0}\sim\overline{\rho}_{0}$.
-Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.3) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$.
+Then, there exists a strong solution $\overline{X}\colon\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.3) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$.
 Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.3) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$, and it holds that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.4)
+$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}\coloneq\operatorname{Law}(\overline{X}_{t}).$ (2.4)
 Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C\left([0,T],\mathcal{P}_{p}(\mathbf{R}^{d})\right)$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
 $\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}\Bigr{)}+\frac{1}{2}\nabla\cdot\nabla\cdot\Bigl{(}D(\overline{\rho},x)\overline{\rho}\Bigr{)},$ (2.5)
-where $D(\overline{\rho},x):=S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}^{\mathsf{T}}$.
+where $D(\overline{\rho},x)\coloneq S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}^{\mathsf{T}}$.
 Theorem 2.4 (CBS: Existence and uniqueness for the mean-field SDE).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and fix a final time $T>0$.
 Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$.
-Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.5) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.5) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, and it holds that
-$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.6)
+Then, there exists a strong solution $\overline{X}\colon\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.5) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.5) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, and it holds that
+$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}\coloneq\operatorname{Law}(\overline{X}_{t}).$ (2.6)
 Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C([0,T],\mathcal{P}_{p}(\mathbf{R}^{d}))$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
@@
 2.3 Key preparatory lemma
-We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of Markov’s inequality,
+We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of Markov’s inequality.
 Lemma 2.5 (Bound on the probability of large excursions).
@@
 driven by the same Brownian motions and initial conditions as (2.8). In the case of CBO, the drift and diffusion coefficients for (2.8) and (2.9) are given by
-$b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=S\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)},$ (2.10)
+$b(x,\mu)\coloneq-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu)\coloneq S\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)},$ (2.10)
 where the diffusion operator $S\colon\mathbf{R}^{d}\rightarrow\mathbf{R}^{d\times d}$ is, as always in this work, a globally Lipschitz continuous function.
 Theorem 2.6 (Mean-field limit for CBO).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ has moments of sufficiently high order.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ has bounded moments of all orders.
 Consider the systems (2.8) and (2.9) with the coefficients given by (2.10).
@@
 Remark 2.7.
-We prove below the following more general statement: suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$.
+We prove below the following more general statement: Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$.
 If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
@@
 Theorem 2.6 implies pathwise Wasserstein-$p$ chaos. More precisely, it holds that
-$W_{p}\Bigl{(}\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\Bigr{)}\leqslant CJ^{-\frac{1}{2}},$
+$\mathcal{W}_{p}\Bigl{(}\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\Bigr{)}\leqslant CJ^{-\frac{1}{2}},$
 where $\rho^{J}_{[0,T]}$ is the pathwise law of the microscopic particle system over $[0,T]$ and $\overline{\rho}_{[0,T]}$ is the pathwise mean-field law over $[0,T]$. That is to say,
 […]
-Here, $W_{p}$ denotes the pathwise Wasserstein-$p$ distance, associated with the following metric on $\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}$:
+Here, $\mathcal{W}_{p}$ denotes the pathwise Wasserstein-$p$ distance, associated with the following metric on $\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}$:
 $d(\mathbf{X},\mathbf{Y})=\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}-Y^{(j)}_{t}\right\rvert,\qquad\mathbf{X},\mathbf{Y}\in\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}.$
@@
 We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant p_{\mathcal{M}}(s,\ell)$, where $p_{\mathcal{M}}(s,\ell)\in^{+}$ is given in (2.3), there exists $\widetilde{L}_{\mathcal{M}}=\widetilde{L}_{\mathcal{M}}(R,p)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,W_{p}(\mu,\nu).$
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,\mathcal{W}_{p}(\mu,\nu).$
   2. (b)
@@
     Convergence of the weighted mean for i.i.d. samples. We prove in Lemma 3.7 that, for any reals $r>p>0$ and any $\mu\in\mathcal{P}_{r}(\mathbf{R}^{d})$, there is $C$ such that
-    $\mathbf{E}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
+    $\mathbf{E}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}\coloneq\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
 Proof of Theorem 2.6.
@@
 where we used that the function $z\mapsto(q-z)/z^{2}$ is decreasing.
-Take $p\in[2\vee p_{\mathcal{M}},q/2]$ and fix $j\in\llbracket 1,J\rrbracket$.
+Take $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ and fix $j\in\llbracket 1,J\rrbracket$.
 As in Sznitman’s argument [ReviewChaintronII, Theorem 3.1], we begin by writing
@@
 In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [ReviewChaintronII, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
-$\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\},$ (2.16)
-where $R>\mathbf{E}[Z_{j}]$ is fixed. Define
-$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}.$
-By Theorem 2.3 it holds that $\mathbf{E}\left[\lvert Z_{j}\rvert^{q/p}\right]<\infty.$ Note that $q\geqslant 2p$. Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
+$\Omega_{J,t}\coloneq\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\},$ (2.16)
+where
+$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p},$
+and $R>\mathbf{E}[Z_{j}]$ is fixed.
+By Theorem 2.3 it holds that $\mathbf{E}\left[\lvert Z_{j}\rvert^{q/p}\right]<\infty.$ Note that $q\geqslant 2p$.
+Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
 $\forall t\in[0,T],\qquad\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\Omega_{J,t}\right]\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant CJ^{-\frac{q}{2p}}.$
@@
   • Since the $p$-th moment of $\overline{\mu}_{s}^{J}$ is bounded from above by $R$ for all $\omega\in\Omega\setminus\Omega_{J,s}$, we can use the stability estimate on the weighted mean given (see Item (a) and Corollary 3.3) to obtain
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]$
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[\mathcal{W}_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]$
     for the first term in (2.17). Since in addition
-    $\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
+    $\mathbf{E}\left[\mathcal{W}_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
     it follows that
@@
   • We now bound the second term in (2.17). From Item (b) (see Lemma 3.5) as well as (2.4) and Proposition A.4 we have
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right)\leqslant 2^{q-1}\left(\kappa+\widetilde{\kappa}\right)$
-    where $\widetilde{\kappa}:=C_{2}\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{q}\right]<\infty$.
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right)\leqslant 2^{q-1}\left(\kappa+\overline{\kappa}\right)$
+    where $\overline{\kappa}:=C_{2}\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{q}\right]<\infty$.
 Here, $C_{2}>0$ is the constant from Proposition A.4.
 Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
-    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{\frac{(q-1)p}{q}}\left(\kappa+\widetilde{\kappa}\right)^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
+    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{\frac{(q-1)p}{q}}\left(\kappa+\overline{\kappa}\right)^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
 We have thus shown that
@@
 which in turn by (2.15) implies that
-$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
+$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
 Bounding the second term in (2.14)
@@
 The proof is similar to that of McKean’s theorem [ReviewChaintronII, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
-A crucial detail here is the following: Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below.
+A crucial detail here is that Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below.
 This allows us to define $\Omega_{J,t}$ in (2.16) only in terms of the i.i.d.
@@
 Theorem 2.10 (Mean-field limit for CBS).
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$ has moments of sufficiently high order.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$ has moments of all orders.
 Consider the systems (2.8) and (2.9) with the coefficients given by
-$b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}(\mu)}.$ (2.18)
+$b(x,\mu)\coloneq-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu)\coloneq\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}(\mu)}.$ (2.18)
 Then for all $p>0$, there is $C>0$ independent of $J$ such that
@@
 In the same way as for CBO, it is possible to prove the following more general statement.
-Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$.
+Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$.
+Assume that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$.
 If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
@@
 We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant 2p_{\mathcal{M}}(\ell,s)$, there exists $\widetilde{L}_{\sqrt{\mathcal{C}}}(R,p)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,\mathcal{W}_{p}(\mu,\nu).$
   • Moment estimate for the particle system.
@@
 We prove in Lemma 3.8 that, for any reals $r>2p>0$ and any $\mu\in\mathcal{P}_{r}(\mathbf{R}^{d})$ such that $\mathcal{C}_{\beta}(\mu)\succ 0$, then there is $C$ such that
-    $\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
+    $\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}\coloneq\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
   • No collapse in finite time. In order to apply the third item with $\mu=\overline{\rho}_{t}$, we prove in Proposition 3.9 that the assumption $\mathcal{C}(\overline{\rho}_{0})\succ 0$ implies that
-    $\eta:=\inf_{t\in[0,T]}\lambda_{\text{min}}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$
+    $\eta\coloneq\inf_{t\in[0,T]}\lambda_{\text{min}}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$
 With these ingredients at hand, we can prove Theorem 2.10 in the same way as for CBO, so we skip the proof here.
 Let us mention, however, that a different proof based on a technique using stopping times is presented in Subsection 5.1.
+2.6 Numerical experiment
+In this section, we present a numerical experiment illustrating our first main result, on the quantitative mean field limit for CBO; see Theorem 2.6.
+To this end, we apply CBO to minimize the Ackley function, a standard benchmark function in optimization given by
+[…]
+where $d$ is the dimension.
+This function, which is illustrated in Figure 2, is globally Lipschitz continuous and bounded from above and below uniformly over d, and so $f_{\rm Ack}\in\mathcal{A}(0,0,0)$.
+To illustrate the theorem, we first approximate the mean field dynamics by simulating the interacting particle system (1.1) using a large number $J_{\infty}$ of particles.
+Then, for various smaller values of $J\ll J_{\infty}$, we simulate the CBO evolution of a system of $J$ particles, initialized at the same positions and driven by the same Brownian motions as the first $J$ particles of the large system of size $J_{\infty}$.
+In all cases, we use CBO with isotropic noise, discretized using the Euler–Maruyama method with time step $\Delta t$.
+All the parameters used in this experiment are given in Table 1.
+To exhibit the scaling as $J^{-\frac{1}{2}}$ in Theorem 2.6, we fix $p=2$ and approximate the left-hand side of (2.11) using a Monte Carlo approach with $M$ independent realizations.
+More precisely, omitting discretization for simplicity, we use the following estimator $\widehat{E}(J)$ of $E(J)$, where
+$E(J)\coloneq\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2}\right],\qquad\widehat{E}(J)\coloneq\frac{1}{MJ}\sum_{m=1}^{M}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Big{\lvert}X^{(j/J,m)}_{t}-X^{(j/J_{\infty},m)}_{t}\Big{\rvert}^{2}.$
+Here $X^{(j/J,m)}_{t}$ denotes the position of the $j$-th particle in the system of size $J$, for the $m$-th independent realization of the particle systems.
+Parameter Notation Value
+$\int$ Dimension $d$ 2
+$\int$ Inverse temperature $\beta$ $3$
+$\int$ Diffusion coefficient $\sigma$ $\frac{1}{5}$
+$\int$ Initial law $\overline{\rho}_{0}$ $\mathcal{N}(0,\operatorname{I}_{2})$
+$\int$ Time step $\Delta t$ 0.01
+$\int$ Final time $T$ 1
+$\int$ Size for mean field approximation $J_{\infty}$ $10^{6}$
+$\int$ Number of independent simulations $M$ $100$
+Table 1: Parameters of the numerical experiment described in Subsection 2.6.
+The results of the numerical experiment are illustrated in the right panel Figure 2.
+The expected scaling $E(J)\propto\frac{1}{J}$ is clearly observed for large values of $J$.
+We note, however, that this scaling of the error holds only for relatively large $J$.
+For small $J$, the effective number of particles contributing to the weighted average, as quantified by the effective sample size [MR3696003], is close to 1 in the early stages of the simulation.
+Therefore, the Monte Carlo scaling is not expected in this regime.
+Figure 2: Left: Ackley function.
+Right: Monte Carlo estimator $\widehat{E}(J)$ of $E(J)$ for $J\in\{10\cdot 2^{k}\}_{k=0}^{14}$, and straight line exhibiting the asymptotic scaling as $J^{-1}$ of $E(J)$.
+Error bars corresponding to two standard deviations are depicted on both sides of the Monte Carlo estimator.
 3 Auxiliary results
@@
     For all $R>0$ and for all $p\geqslant p_{\mathcal{M}}(s,\ell)$, where $p_{\mathcal{M}}(s,\ell)$ is given in (2.3), there exists $L_{\mathcal{M}}=L_{\mathcal{M}}(p,R)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant L_{\mathcal{M}}\,W_{p}(\mu,\nu).$ (3.1)
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant L_{\mathcal{M}}\,\mathcal{W}_{p}(\mu,\nu).$ (3.1)
   2. (b)
     For all $R>0$ and for all $p\geqslant p_{\mathcal{C}}(s,\ell)$, where $p_{\mathcal{C}}(s,\ell)$ is given in (2.3), there exists $L_{\mathcal{C}}=L_{\mathcal{C}}(p,R)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L_{\mathcal{C}}\,W_{p}(\mu,\nu).$ (3.2)
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L_{\mathcal{C}}\,\mathcal{W}_{p}(\mu,\nu).$ (3.2)
 Proof.
@@
 In the notation of Lemma A.1 we can write
-$\mathcal{M}_{\beta}(\mu)=\frac{\mu[g]}{\mu[h]},\qquad g(x):=\operatorname{e}^{-\beta f(x)}x,\qquad h(x):=\operatorname{e}^{-\beta f(x)}.$
+$\mathcal{M}_{\beta}(\mu)=\frac{\mu[g]}{\mu[h]},\qquad g(x):=\operatorname{e}^{-\beta f(x)}x,\qquad h(x)\coloneq\operatorname{e}^{-\beta f(x)}.$
 Condition A.1 is satisfied for $\varsigma:=s+1$ if $\ell=0$ or $\varsigma:=0$ if $\ell>0$. Therefore, the claim follows from Lemma A.1.
@@
 $\Bigl{\lVert}\mathcal{M}_{\beta}(\mu)\otimes\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\otimes\mathcal{M}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}$ $\leqslant\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\cdot\bigl{\lvert}\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\bigr{\rvert}$
-$\leqslant L_{\mathcal{M}}\,\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\,W_{p}(\mu,\nu)$
-$\leqslant L_{\mathcal{M}}^{2}\,\Bigl{(}W_{p}(\mu,\delta_{0})+W_{p}(\nu,\delta_{0})\Bigr{)}\,W_{p}(\mu,\nu)\leqslant 2L_{\mathcal{M}}^{2}R\,W_{p}(\mu,\nu).$
+$\leqslant L_{\mathcal{M}}\,\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\,\mathcal{W}_{p}(\mu,\nu)$
+$\leqslant L_{\mathcal{M}}^{2}\,\Bigl{(}\mathcal{W}_{p}(\mu,\delta_{0})+W_{p}(\nu,\delta_{0})\Bigr{)}\,\mathcal{W}_{p}(\mu,\nu)\leqslant 2L_{\mathcal{M}}^{2}R\,\mathcal{W}_{p}(\mu,\nu).$
 In order to bound the second term in (3.3), we again use (3.4) to deduce that the function $g$ satisfies the local Lipschitz estimate
@@
 We deduce that the functions $g$ and $h$ satisfy the assumption (A.1) of Lemma A.1 with this $\varsigma$. Therefore, it holds that
-$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\frac{\mu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\mu[\operatorname{e}^{-\beta f(x)}]}-\frac{\nu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\nu[\operatorname{e}^{-\beta f(x)}]}\right\rVert_{\rm F}\leqslant CW_{p}(\mu,\nu),$
+$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\frac{\mu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\mu[\operatorname{e}^{-\beta f(x)}]}-\frac{\nu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\nu[\operatorname{e}^{-\beta f(x)}]}\right\rVert_{\rm F}\leqslant C\mathcal{W}_{p}(\mu,\nu),$
 which completes the proof. ∎
@@
 Then for all $R>0$ and all $p\geqslant 2p_{\mathcal{M}}(s,\ell)$, there exists $L_{\sqrt{\mathcal{C}}}=L_{\sqrt{\mathcal{C}}}(p,R)>0$ such that
-$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant L_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$ (3.5)
+$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant L_{\sqrt{\mathcal{C}}}\,\mathcal{W}_{p}(\mu,\nu).$ (3.5)
 Proof.
@@
 Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$.
-By [MR2459454, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$.
-Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [MR2459454, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$.
+By [MR2459454, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $\mathcal{W}_{p}(\mu^{J},\mu)\to 0$ and $\mathcal{W}_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$.
+Since convergence in $\mathcal{W}_{p}$ implies convergence of the moments of order up to $p$ by [MR2459454, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$.
 Then
@@
 $\qquad+\left\lVert\sqrt{\mathcal{C}_{\beta}(\nu)}-\sqrt{\mathcal{C}_{\beta}(\nu^{J})}\right\rVert_{\rm F}+\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\mu^{J})}\right\rVert_{\rm F}.$
-The first term is bounded from above by $CW_{p}(\mu^{J},\nu^{J})$ by the base case, while the other two terms converge to 0 in the limit as $J\to\infty$ by Proposition 3.1.
+The first term is bounded from above by $C\mathcal{W}_{p}(\mu^{J},\nu^{J})$ by the base case, while the other two terms converge to 0 in the limit as $J\to\infty$ by Proposition 3.1.
 Taking the limit $J\to\infty$, we deduce that
-$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant CW_{p}(\mu,\nu).$
+$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant C\mathcal{W}_{p}(\mu,\nu).$
 Proof of the statement for empirical measures
 Fix $R>0$. By [MR1964483, p.5], the Wasserstein distance between empirical measures $\mu^{J}$ and $\nu^{J}$ of the form (3.6) is equal to
-$W_{p}(\mu^{J},\nu^{J})=\min_{\sigma\in\mathcal{S}_{J}}\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{\sigma(j)}\right\rvert^{p}\right)^{\frac{1}{p}},$
+$\mathcal{W}_{p}(\mu^{J},\nu^{J})=\min_{\sigma\in\mathcal{S}_{J}}\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{\sigma(j)}\right\rvert^{p}\right)^{\frac{1}{p}},$
 where $\mathcal{S}_{J}$ denotes the set of permutations in $\{1,\dotsc,J\}$.
@@
     For all $R>0$ and for all $p\geqslant p_{\mathcal{M}}$, there exists $\widetilde{L}_{\mathcal{M}}=\widetilde{L}_{\mathcal{M}}(R,p)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,W_{p}(\mu,\nu).$
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,\mathcal{W}_{p}(\mu,\nu).$
   2. (b)
     For all $R>0$ and $p\geqslant 2p_{\mathcal{M}}$, there exists $\widetilde{L}_{\mathcal{C}}=\widetilde{L}_{\mathcal{C}}(R,p)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\Bigr{\rVert}_{\rm F}\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\Bigr{\rVert}_{\rm F}\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,\mathcal{W}_{p}(\mu,\nu).$
 Remark 3.4.
@@
   • For all $R>0$ and $p\geqslant p_{\mathcal{T}}$, there exists $L_{\mathcal{T}}(R)>0$ such that
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)-\mathcal{T}(\nu)\bigr{\rVert}\leqslant L_{\mathcal{T}}(R)\,W_{p}(\mu,\nu).$ (3.8)
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)-\mathcal{T}(\nu)\bigr{\rVert}\leqslant L_{\mathcal{T}}(R)\,\mathcal{W}_{p}(\mu,\nu).$ (3.8)
 […]
   • For all $p\geqslant p_{\mathcal{T}}$, there is $K>0$ such that
-    $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)\bigr{\rVert}\leqslant KW_{p}(\mu,\delta_{0}).$ (3.9)
+    $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)\bigr{\rVert}\leqslant K\mathcal{W}_{p}(\mu,\delta_{0}).$ (3.9)
     When $\mathcal{T}=\mathcal{M}_{\beta}$, this follows immediately from the inequality $\operatorname{e}^{-\beta f^{*}}\leqslant\operatorname{e}^{-\beta f(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})}\leqslant\operatorname{e}^{-\beta f_{*}}$ if $\ell=0$ and from Proposition A.4 if $\ell>0$, where $f^{*}$ the supremum of $f$ over d.
@@
 Now, assume for contradiction that there exists $R>0$ and $\{(\mu_{i},\nu_{i})\}_{i\in\mathbf{N}}$ in $\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}$ such that
-$\forall i\in\mathbf{N},\qquad\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{W_{p}\left(\mu_{i},\nu_{i}\right)}\geqslant i.$ (3.11)
-In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of 3.11 would be bounded from above uniformly in $i$ by (3.8).
-Therefore, $W_{p}(\nu_{i},\delta_{0})\to\infty$.
+$\forall i\in\mathbf{N},\qquad\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{\mathcal{W}_{p}\left(\mu_{i},\nu_{i}\right)}\geqslant i.$ (3.11)
+In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of (3.11) would be bounded from above uniformly in $i$ by (3.8).
+Therefore, $\mathcal{W}_{p}(\nu_{i},\delta_{0})\to\infty$.
 On the other hand, by the triangle inequality,
-$W_{p}\left(\mu_{i},\nu_{i}\right)\geqslant W_{p}\left(\nu_{i},\delta_{0}\right)-W_{p}\left(\mu_{i},\delta_{0}\right)\geqslant W_{p}\left(\nu_{i},\delta_{0}\right)-R.$
+$\mathcal{W}_{p}\left(\mu_{i},\nu_{i}\right)\geqslant\mathcal{W}_{p}\left(\nu_{i},\delta_{0}\right)-\mathcal{W}_{p}\left(\mu_{i},\delta_{0}\right)\geqslant\mathcal{W}_{p}\left(\nu_{i},\delta_{0}\right)-R.$
 Therefore, by (3.9), we deduce that
-$\limsup_{i\to\infty}\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{W_{p}\left(\mu_{i},\nu_{i}\right)}\leqslant K\limsup_{i\to\infty}\frac{W_{p}(\nu_{i},\delta_{0})+R}{W_{p}\left(\nu_{i},\delta_{0}\right)-R}=K,$
+$\limsup_{i\to\infty}\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{\mathcal{W}_{p}\left(\mu_{i},\nu_{i}\right)}\leqslant K\limsup_{i\to\infty}\frac{\mathcal{W}_{p}(\nu_{i},\delta_{0})+R}{\mathcal{W}_{p}\left(\nu_{i},\delta_{0}\right)-R}=K,$
 which contradicts (3.11). ∎
@@
 $\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}\right)-\mathcal{M}_{\beta}\bigl{(}\mu\bigr{)}\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
-Proof of Lemma 3.7.
+Proof.
 We apply Theorem 3.6 with $q=\frac{p(r+2)}{r-p}$, in which case $\frac{p(q+2)}{q-p}=r.$ In our setting
@@
 For a collection of particles $\mathbf{X}\in\mathbf{R}^{dJ}$, the notation $\lvert\mathbf{X}\rvert$ denotes the usual Euclidean norm; that is
-$\lvert\mathbf{X}\rvert:=\sqrt{\sum_{j=1}^{J}\left\lvert X^{(j)}\right\rvert^{2}}.$
+$\lvert\mathbf{X}\rvert\coloneq\sqrt{\sum_{j=1}^{J}\left\lvert X^{(j)}\right\rvert^{2}}.$
 For the proof of Theorem 2.2 we need the following stability result, which only requires local Lipschitz continuity of $f$, in contrast to the results in Subsection 3.1.
@@
 where we used convexity of the square function and the fact that the weights, given hereafter, are bounded from above by 1:
-$w_{j}(\mathbf{X}):=\frac{\operatorname{e}^{-\beta f\left(X^{(j)}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}\right)}}.$
+$w_{j}(\mathbf{X})\coloneq\frac{\operatorname{e}^{-\beta f\left(X^{(j)}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}\right)}}.$
 Similarly, for the weighted covariance, it holds by the triangle inequality that
@@
 $\mathrm{d}\mathbf{X}_{t}=F(\mathbf{X}_{t})\,\mathrm{d}t+G(\mathbf{X}_{t})\,\mathrm{d}\mathbf{W}_{t}$
-where $\mathbf{X}_{t}:=\Bigl{(}X^{(1)}_{t},\dots,X^{(J)}_{t}\Bigr{)}$. The drift and diffusion coefficients are given by
-$F(\mathbf{X}_{t}):=-\begin{pmatrix}X^{(1)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\\ \vdots\\ X^{(J)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\end{pmatrix}\in\mathbf{R}^{dJ},$
+where $\mathbf{X}_{t}\coloneq\Bigl{(}X^{(1)}_{t},\dots,X^{(J)}_{t}\Bigr{)}$. The drift and diffusion coefficients are given by
+$F(\mathbf{X}_{t})\coloneq-\begin{pmatrix}X^{(1)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\\ \vdots\\ X^{(J)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\end{pmatrix}\in\mathbf{R}^{dJ},$
 and
-$G(\mathbf{X}_{t}):=$ $\sqrt{2\lambda^{-1}}\operatorname{diag}\left(\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)},\dots,\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\right)\in\mathbf{R}^{dJ\times dJ}.$
+$G(\mathbf{X}_{t})\coloneq$ :- $\sqrt{2\lambda^{-1}}\operatorname{diag}\left(\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)},\dots,\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\right)\in\mathbf{R}^{dJ\times dJ}.$
 By Lemma 4.1, the functions $F$ and $G$ have sublinear growth and are locally Lipschitz continuous.
@@
   • There exists $c>0$ such that
-    $\forall\mathbf{X}\in\mathbf{R}^{dJ},\qquad\mathcal{L}\phi(\mathbf{X}):=F(\mathbf{X})\cdot\nabla\phi(\mathbf{X})+\frac{1}{2}\operatorname*{tr}\left(G^{\mathsf{T}}(\mathbf{X})\operatorname{D}^{2}\phi(\mathbf{X})G(X)\right)\leqslant c\phi(\mathbf{X}).$
-Take $\phi(\mathbf{X}):=\tfrac{1}{2}|\mathbf{X}|^{2}$. Then
+    $\forall\mathbf{X}\in\mathbf{R}^{dJ},\qquad\mathcal{L}\phi(\mathbf{X})\coloneq F(\mathbf{X})\cdot\nabla\phi(\mathbf{X})+\frac{1}{2}\operatorname*{tr}\left(G^{\mathsf{T}}(\mathbf{X})\operatorname{D}^{2}\phi(\mathbf{X})G(X)\right)\leqslant c\phi(\mathbf{X}).$
+Take $\phi(\mathbf{X})\coloneq\tfrac{1}{2}|\mathbf{X}|^{2}$. Then
 $\mathcal{L}\phi(\mathbf{X})$ $=-\sum_{j=1}^{J}\left(X^{(j)}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\cdot X^{(j)}+J\lambda^{-1}\operatorname*{tr}\Bigl{(}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}$
@@
 We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3.
-In the following, we endow the vector space $\mathcal{X}:=\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
-$\left\lVert(m,\Gamma)\right\rVert:=\left({|m|^{2}+\left\lVert\Gamma\right\rVert^{2}_{\rm F}}\right)^{\frac{1}{2}}.$
+In the following, we endow the vector space $\mathcal{X}\coloneq\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
+$\left\lVert(m,\Gamma)\right\rVert\coloneq\left({|m|^{2}+\left\lVert\Gamma\right\rVert^{2}_{\rm F}}\right)^{\frac{1}{2}}.$
 The Hölder space $C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}$ is given by
-$C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}:=\left\{\varphi\colon[0,T]\rightarrow\mathcal{X}:\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}<\infty\right\},$
+$C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}\coloneq\left\{\varphi\colon[0,T]\rightarrow\mathcal{X}:\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}<\infty\right\},$
 where
-$\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}:=\sup_{t\in[0,T]}\bigl{\lVert}\varphi(t)\bigr{\rVert}+[\varphi]_{C^{0,\frac{1}{2}}}<\infty\qquad\text{and}\qquad[\varphi]_{C^{0,\frac{1}{2}}}:=\sup_{0\leqslant s<t\leqslant T}\frac{\bigl{\lVert}\varphi(t)-\varphi(s)\bigr{\rVert}}{|t-s|^{\frac{1}{2}}}.$
+$\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}\coloneq\sup_{t\in[0,T]}\bigl{\lVert}\varphi(t)\bigr{\rVert}+[\varphi]_{C^{0,\frac{1}{2}}}<\infty\qquad\text{and}\qquad[\varphi]_{C^{0,\frac{1}{2}}}\coloneq\sup_{0\leqslant s<t\leqslant T}\frac{\bigl{\lVert}\varphi(t)-\varphi(s)\bigr{\rVert}}{|t-s|^{\frac{1}{2}}}.$
 In order to ensure that certain matrices are positive semidefinite, we use the map $\mathcal{R}\colon\mathbf{R}^{d\times d}\rightarrow\mathbf{R}^{d\times d}$ defined by $\mathcal{R}(\Gamma)=\sqrt{\Gamma\Gamma^{\mathsf{T}}}$.
@@
 admits a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$.
-Moreover, denoting by $\rho_{t}:=\operatorname{Law}(Y_{t})\in\mathcal{P}(\mathbf{R}^{d})$ the pointwise law of $Y_{t}$, the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ belong to $C([0,T],\mathbf{R}^{d})$ and $C([0,T],\mathbf{R}^{d\times d})$ respectively.
+Moreover, denoting by $\rho_{t}\coloneq\operatorname{Law}(Y_{t})\in\mathcal{P}(\mathbf{R}^{d})$ the pointwise law of $Y_{t}$, the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ belong to $C([0,T],\mathbf{R}^{d})$ and $C([0,T],\mathbf{R}^{d\times d})$ respectively.
 Proof.
@@
 $(m,\Gamma)$ $\mapsto\bigl{(}\mathcal{M}_{\beta}(\rho_{t}),\mathcal{C}_{\beta}(\rho_{t})\bigr{)},$
-where $\rho_{t}:={\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3).
+where $\rho_{t}\coloneq{\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3).
 Thanks to Lemma 4.2, the map $\mathcal{T}$ is well-defined.
@@
 In order to prove that $\mathcal{T}$ is a compact operator, fix $R>0$ and consider the ball
-$B_{R}:=\left\{(m,\Gamma)\in C([0,T],\mathcal{X}):\left\lVert(m,\Gamma)\right\rVert_{L^{\infty}([0,T])}\leqslant R\right\}.$
+$B_{R}\coloneq\left\{(m,\Gamma)\in C([0,T],\mathcal{X}):\left\lVert(m,\Gamma)\right\rVert_{L^{\infty}([0,T])}\leqslant R\right\}.$
 By the Arzelà–Ascoli theorem, we have a compact embedding
@@
 $\mathbf{E}\left\lvert Y_{t}-Y_{r}\right\rvert^{p}\leqslant L\lvert t-r\rvert^{\frac{p}{2}}.$
-It follows that $W_{p}(\rho_{t},\rho_{r})\leqslant L^{\frac{1}{p}}\lvert t-r\rvert^{\frac{1}{2}}$, which by Proposition 3.1 implies the Hölder continuity with exponent $\frac{1}{2}$ of the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$.
+It follows that $\mathcal{W}_{p}(\rho_{t},\rho_{r})\leqslant L^{\frac{1}{p}}\lvert t-r\rvert^{\frac{1}{2}}$, which by Proposition 3.1 implies the Hölder continuity with exponent $\frac{1}{2}$ of the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$.
 This completes the proof that $\mathcal{T}$ is compact.
@@
 for some $\xi\in[0,1]$, and let $(Y_{t})$ denote the corresponding solution to (4.3). By (4.7), the stochastic process $(Y_{t})$ is also a solution to
-$\mathrm{d}Y_{t}=-\Bigl{(}Y_{t}-\xi\mathcal{M}_{\beta}(\rho_{t})\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\xi\mathcal{C}_{\beta}(\rho_{t})}\,\mathrm{d}W_{t},\qquad\rho_{t}:={\rm Law}(Y_{t}).$
+$\mathrm{d}Y_{t}=-\Bigl{(}Y_{t}-\xi\mathcal{M}_{\beta}(\rho_{t})\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\xi\mathcal{C}_{\beta}(\rho_{t})}\,\mathrm{d}W_{t},\qquad\rho_{t}\coloneq{\rm Law}(Y_{t}).$
 But then, reasoning as in Subsection 3.2 using a Grönwall-type argument and Proposition A.4 with $u=\ell$, we find that
@@
 Step 4: Showing uniqueness
-For uniqueness, let $(m_{t},\Gamma_{t})$ and $(\widehat{m}_{t},\widehat{\Gamma}_{t})$ be two fixed points of $\mathcal{T}$ with corresponding solutions $Y_{t},\widehat{Y}_{t}\in C([0,T],\mathcal{X})$ of (4.3) and let $z_{t}:=Y_{t}-\widehat{Y}_{t}$.
+For uniqueness, let $(m_{t},\Gamma_{t})$ and $(\widehat{m}_{t},\widehat{\Gamma}_{t})$ be two fixed points of $\mathcal{T}$ with corresponding solutions $Y_{t},\widehat{Y}_{t}\in C([0,T],\mathcal{X})$ of (4.3) and let $z_{t}\coloneq Y_{t}-\widehat{Y}_{t}$.
 By definition of $\mathcal{T}$ and since $\mathcal{C}_{\beta}(\mu)$ is symmetric and positive semidefinite for all $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d})$, we have $\mathcal{R}\left(\Gamma_{t}\right)=\Gamma_{t}$ and $\mathcal{R}\bigl{(}\widehat{\Gamma}_{t}\bigr{)}=\widehat{\Gamma}_{t}$ for all $t\in[0,T]$.
@@
 By Proposition 3.9, we have that
-$\eta:=\inf\left\{\lambda_{\rm min}\Bigl{(}\mathcal{C}_{\beta}(\rho_{t})\Bigr{)}:t\in[0,T]\right\}>0.$
+$\eta\coloneq\inf\left\{\lambda_{\rm min}\Bigl{(}\mathcal{C}_{\beta}(\rho_{t})\Bigr{)}:t\in[0,T]\right\}>0.$
 By virtue of the van Hemmen–Ando inequality [vanHemmen1980, Proposition 3.2], it holds for any unitarily invariant norm, thus in particular for the Frobenius norm, that
 $\left\lVert\sqrt{\mathcal{C}_{\beta}(\rho_{r})}-\sqrt{\mathcal{C}_{\beta}(\widehat{\rho}_{r})}\right\rVert_{\rm F}\leqslant\eta^{-\frac{1}{2}}\left\lVert\mathcal{C}_{\beta}(\rho_{r})-\mathcal{C}_{\beta}(\widehat{\rho}_{r})\right\rVert_{\rm F}\,.$
-Using Proposition 3.1 and noting that $W_{p}(\rho_{r},\widehat{\rho}_{r})^{p}\leqslant\mathbf{E}\lvert z_{r}\rvert^{p}$ by definition of the Wasserstein distance, we conclude that
+Using Proposition 3.1 and noting that $\mathcal{W}_{p}(\rho_{r},\widehat{\rho}_{r})^{p}\leqslant\mathbf{E}\lvert z_{r}\rvert^{p}$ by definition of the Wasserstein distance, we conclude that
 $\mathbf{E}\left[\sup_{r\in[0,t]}\bigl{\lvert}z_{r}\bigr{\rvert}^{p}\right]\leqslant C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,r]}\bigl{\lvert}z_{u}\bigr{\rvert}^{p}\right]\,\mathrm{d}r.$
@@
 $\mathbf{E}\phi\left(\overline{Y}_{t}\right)-\mathbf{E}\phi\left(\overline{Y}_{r}\right)=\mathbf{E}\left[\int_{r}^{t}\mathcal{L}_{u}\phi\left(\overline{Y}_{u}\right)\,\mathrm{d}u\right],$ (4.8)
-where $\mathcal{L}_{u}\phi(y):=-\bigl{(}y-\mathcal{M}_{\beta}\left(\overline{\rho}_{u}\right)\bigr{)}\cdot\nabla\phi(y)+\lambda^{-1}\mathcal{C}_{\beta}\left(\overline{\rho}_{u}\right):\operatorname{D}^{2}\phi(y).$ It follows from (4.8) that $\overline{\rho}_{t}$ satisfies, for all smooth compactly supported $\phi\colon\mathbf{R}^{d}\to\mathbf{R}$,
+where $\mathcal{L}_{u}\phi(y)\coloneq-\bigl{(}y-\mathcal{M}_{\beta}\left(\overline{\rho}_{u}\right)\bigr{)}\cdot\nabla\phi(y)+\lambda^{-1}\mathcal{C}_{\beta}\left(\overline{\rho}_{u}\right):\operatorname{D}^{2}\phi(y).$ It follows from (4.8) that $\overline{\rho}_{t}$ satisfies, for all smooth compactly supported $\phi\colon\mathbf{R}^{d}\to\mathbf{R}$,
 $\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbf{R}^{d}}\phi(x)\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\mathcal{L}_{t}\phi\,\overline{\rho}_{t}(\mathrm{d}x),$ (4.9)
@@
 For the proof of Theorem 2.6, this flexibility is in fact not necessary, as we could have employed instead of $\Omega_{J,t}$ in (2.16) the event
-    $\Omega_{J}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\}.$ (5.1)
+    $\Omega_{J}:=\left\{\omega\in\Omega:\frac{1}{J}\sup_{t\in[0,T]}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\}.$ (5.1)
     In certain contexts, however, being able to exclude a time-dependent event could be very useful.
@@
     Using Corollary 3.3, together with the fact
-    $\mathbf{E}\Bigl{[}W_{p}\bigl{(}\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\bigr{)}^{p}\Bigr{]}\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{k=1}^{J}\left\lvert X^{(k)}_{s\wedge\theta_{J}}-\overline{X}^{(k)}_{s\wedge\theta_{J}}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p},$
+    $\mathbf{E}\Bigl{[}\mathcal{W}_{p}\bigl{(}\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\bigr{)}^{p}\Bigr{]}\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{k=1}^{J}\left\lvert X^{(k)}_{s\wedge\theta_{J}}-\overline{X}^{(k)}_{s\wedge\theta_{J}}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p},$
     we deduce that
@@
     Since $q$ was arbitrary given the assumption that $\overline{\rho}_{0}$ has infinitely many moments, it follows that for every $\varepsilon>0$, there exists $C_{\varepsilon}$ such that
-    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\frac{p}{2}\left(1-\varepsilon\right)}.$
+    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant C_{\varepsilon}J^{-\frac{p}{2}\left(1-\varepsilon\right)}.$
     Hence, we recover almost the classical Monte Carlo rate.
@@
 Then for all $p\geqslant\varsigma+1$ and all $R>0$, there exist a constant $C=C(p,R)$ such that
-$\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert\leqslant CW_{p}(\mu,\nu).\qquad$
+$\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert\leqslant C\mathcal{W}_{p}(\mu,\nu).\qquad$
 Proof.
@@
 $\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\int_{\mathbf{R}^{d}}h\,\mathrm{d}\mu\geqslant\eta.$ (A.2)
-Define $B_{K}:=\{x\in\mathbf{R}^{d}:\lvert x\rvert\leqslant K\}$ where $K:=2^{1/p}R$. By Markov’s inequality, we have
+Define $B_{K}\coloneq\{x\in\mathbf{R}^{d}:\lvert x\rvert\leqslant K\}$ where $K\coloneq 2^{1/p}R$. By Markov’s inequality, we have
 $\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\mu\left(B_{K}\right)\geqslant 1-\mu\left(\mathbf{R}^{d}\setminus B_{K}\right)\geqslant 1-\frac{R^{p}}{K^{p}}=\frac{1}{2}.$
-Therefore, we obtain
+Therefore we obtain
 $\forall\mu\in\mathcal{P}_{p,R}(\mathbf{R}^{d}),\qquad\int_{\mathbf{R}^{d}}h\,\mathrm{d}\mu\geqslant\int_{B_{K}}h\,\mathrm{d}\mu\geqslant\frac{1}{2}\min\bigl{\{}h(x):x\in B_{K}\bigr{\}}=:\eta>0,$
@@
     The map $\mathcal{S}\colon\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{n}$ satisfies
-    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\lvert\mathcal{S}(\mu)-\mathcal{S}(\nu)\rvert\leqslant L_{\mathcal{S}}W_{p}(\mu,\nu).$
+    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\lvert\mathcal{S}(\mu)-\mathcal{S}(\nu)\rvert\leqslant L_{\mathcal{S}}\mathcal{W}_{p}(\mu,\nu).$
 Then, for all $R>0$ and $p\geqslant 2\varsigma+2$, there exist a constant $C>0$ such that for all $\mu,\nu\in\mathcal{P}_{p,R}$ and all couplings $\pi\in\Pi(\mu,\nu)$, it holds that
@@
 It follows from (a) that $\left\lVert g(x)\right\rVert^{2}\leqslant C(1+|x|)^{2\varsigma+2}$ for all $x\in\mathbf{R}^{d}$, for a constant $C$. Using (b), we obtain the bound
-$\left(\iint\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{)}\right\rVert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/2}\leqslant\sqrt{\frac{\mu[\lVert g(x)\rVert^{2}]}{\mu[h]}}\bigl{\lvert}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{\rvert}\leqslant\sqrt{\frac{C}{\eta}}L_{\mathcal{S}}W_{p}(\mu,\nu).$
+$\left(\iint\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{)}\right\rVert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/2}\leqslant\sqrt{\frac{\mu[\lVert g(x)\rVert^{2}]}{\mu[h]}}\bigl{\lvert}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{\rvert}\leqslant\sqrt{\frac{C}{\eta}}L_{\mathcal{S}}\mathcal{W}_{p}(\mu,\nu).$
 Bounding the second term
@@
 where we used that $\frac{2\varsigma p}{p-2}\leqslant p$ since $p\geqslant 2+2\varsigma$. In addition, from (b) we have
-$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R,$ (A.4)
+$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}\mathcal{W}_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R,$ (A.4)
 and so we conclude that
@@
 Conclusion
-The claim follows then by combining the bounds on the three terms, and noting, for the first term, that $W_{p}(\mu,\nu)$ is bounded from above by the bracket on the right-hand side of (A.3) by definition of the Wasserstein distance.
+The claim follows then by combining the bounds on the three terms, and noting, for the first term, that $\mathcal{W}_{p}(\mu,\nu)$ is bounded from above by the bracket on the right-hand side of (A.3) by definition of the Wasserstein distance.
 ∎
@@
 $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x)$
-for a constant $C$ depending only on $\beta$ and $f$.
-When $f$ is unbounded, proving an inequality of this type is more involved.
-The following is a generalization of [carrillo2018analytical, Lemma 3.3].
-Proposition A.3.
-Suppose that $f$ is bounded from below by $f_{\star}$ and that there are real numbers $u\geqslant\ell>0$ and positive constants $c_{u},c_{\ell}$ such that
-$\forall x\in\mathbf{R}^{d},\qquad\left\{\begin{aligned} f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)}\\ f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.\end{aligned}\right.$ (A.6)
-Then for all $p,q>0$ there are constants $C_{1}$ and $C_{2}$ depending on $(p,q,u,\ell,c_{u},c_{\ell})$ such that
-$\forall\mu\in\mathcal{P}_{q}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant\left(C_{1}+C_{2}\int_{\mathbf{R}^{d}}\lvert x\rvert^{q}\,\mu(\mathrm{d}x)\right)^{\frac{pu}{q\ell}}$ (A.7)
-Furthermore, if $u=\ell$ and $p\leqslant q$, then this statement holds with $C_{1}=0$.
-The proof relies on the following lemma.
+for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved.
+To this end, we first state and prove the following auxiliary lemma.
 Lemma A.3.
@@
 $\forall\nu\in\mathcal{P}_{q}(\mathbf{R}^{+}),\qquad\frac{\int_{\mathbf{R}^{+}}y^{p}\,\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\beta y}\,\nu(\mathrm{d}y)}\leqslant C\left(1+\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{p}{q}},$ (A.6)
-where $\mathbf{R}^{+}:=[0,\infty)$. The constant is given by
+[…]
 $C(q,p,\beta)=\left(1-\frac{1}{2^{q}}\right)^{-1}(p\beta^{-1})^{p}\operatorname{e}^{2\beta-p}.$
@@
 Fix $\nu\in\mathcal{P}(\mathbf{R}^{+})$. If the right-hand side of (A.6) is infinite, then the inequality (A.6) holds with any constant $C$, so we assume from now on that
-$\lambda:=\left(\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{1}{q}}<\infty.$
+$\lambda\coloneq\left(\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{1}{q}}<\infty.$
 Step 1
@@
 In the first inequality of (A.7), we used that, for any increasing function $g\colon\mathbf{R}^{+}\to\mathbf{R}$, the function
-$F(\gamma):=\frac{\int_{\mathbf{R}^{+}}g(z)\,\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}$
+$F(\gamma)\coloneq\frac{\int_{\mathbf{R}^{+}}g(z)\,\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}$
 is a non-increasing function of $\gamma$.
@@
 A simple calculation indeed gives that
-$F^{\prime}(\gamma)=-\int_{\mathbf{R}^{+}}g(z)\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z),\qquad L_{\gamma}\pi(\mathrm{d}z):=\frac{\operatorname{e}^{-\gamma z}\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)},\qquad\overline{z}:=\int_{\mathbf{R}^{+}}z\,L_{\gamma}\pi(\mathrm{d}z).$
+$F^{\prime}(\gamma)=-\int_{\mathbf{R}^{+}}g(z)\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z),\qquad L_{\gamma}\pi(\mathrm{d}z)\coloneq\frac{\operatorname{e}^{-\gamma z}\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)},\qquad\overline{z}\coloneq\int_{\mathbf{R}^{+}}z\,L_{\gamma}\pi(\mathrm{d}z).$
 By decomposing the integration domain into the union $[0,\overline{z})\cup\{\overline{z}\}\cup(\overline{z},\infty)$, and using the fact that $g$ is non-decreasing, we have
@@
 which completes the proof. ∎
-Proof of Proposition A.3.
+Armed with this result, we are ready to prove the following generalization of [carrillo2018analytical, Lemma 3.3].
+Proposition A.4.
+Suppose that $f$ is bounded from below by $f_{\star}$ and that there are real numbers $u\geqslant\ell>0$ and positive constants $c_{u},C_{u},c_{\ell},C_{\ell}$ such that
+$\forall x\in\mathbf{R}^{d},\qquad\left\{\begin{aligned} f(x)-f_{\star}\leqslant c_{u}|x|^{u}+C_{u}\\ f(x)-f_{\star}\geqslant c_{\ell}|x|^{\ell}-C_{\ell}.\end{aligned}\right.$ (A.8)
+Then for all $p,q>0$ there are constants $C_{1}$ and $C_{2}$ depending on $(p,q,u,\ell,c_{u},C_{u},c_{\ell},C_{\ell})$ such that
+$\forall\mu\in\mathcal{P}_{q}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant\left(C_{1}+C_{2}\int_{\mathbf{R}^{d}}\lvert x\rvert^{q}\,\mu(\mathrm{d}x)\right)^{\frac{pu}{q\ell}}$ (A.9)
+Furthermore, if $u=\ell$ and $p\leqslant q$, then this statement holds with $C_{1}=0$.
+Proof.
 We assume without loss of generality that $f_{\star}=0$. By the second inequality in (A.8), we have that
-$\lvert x\rvert^{p}=\left(\lvert x\rvert^{\ell}\right)^{\frac{p}{\ell}}\leqslant\left(\frac{f(x)}{c_{\ell}}+1\right)^{\frac{p}{\ell}}.$ (A.10)
-For convenience, we denote by $C$ any constant depending only on $(p,q,\ell,u,c_{\ell},c_{u})$. From (A.10) we deduce that
+$\lvert x\rvert^{p}=\left(\lvert x\rvert^{\ell}\right)^{\frac{p}{\ell}}\leqslant\left(\frac{f(x)+C_{\ell}}{c_{\ell}}\right)^{\frac{p}{\ell}}.$ (A.10)
+For convenience, we denote by $C$ any constant depending only on $(p,q,\ell,u,c_{\ell},C_{\ell},c_{u},C_{u})$. From (A.10) we deduce that
 $\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}$ $\leqslant C\frac{\int_{\mathbf{R}^{d}}\bigl{(}f(x)+1\bigr{)}^{\frac{p}{\ell}}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}$
@@
 Suppose that $f\in\mathcal{A}(s,0,u)$. Then
-  • for all $R>0$ and $p>1$, there is $C=C(p,R)$
-    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant C\,W_{p}(\mu,\nu),\qquad\varrho:=\frac{(s+1)p}{p-1}$ (A.13)
-  • for all $R>0$ and $p>1$, there is $C=C(p,R)$
-    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\rvert\leqslant C\,W_{p}(\mu,\nu),\qquad\varrho:=\frac{(s+2)p}{p-1}$ (A.14)
+  • for all $R>0$ and $p>1$, there is $C=C(p,R)$ such that
+    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant C\,\mathcal{W}_{p}(\mu,\nu),\qquad\varrho\coloneq\frac{(s+1)p}{p-1}$ (A.13)
+  • for all $R>0$ and $p>1$, there is $C=C(p,R)$ such that
+    $\forall(\mu,\nu)\in\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{\varrho,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\rvert\leqslant C\,\mathcal{W}_{p}(\mu,\nu),\qquad\varrho\coloneq\frac{(s+2)p}{p-1}$ (A.14)
 Remark A.7.
@@
 Acknowledgments
-We are grateful to Yunan Yang, Arnaud Guyader and Louis-Pierre Chaintron for useful discussions, and to the Lorentz Center (Leiden, the Netherlands) for hosting the workshop “Purpose-driven particle systems” in March 2023, which enabled discussions about this work.
-FH is supported by start-up funds at the California Institute of Technology.
+We are grateful to Yunan Yang, Arnaud Guyader and Louis-Pierre Chaintron for useful discussions, to Hui Huang and Hicham Kouhkouh for pointing out an issue with Assumption (H2), and to the anonymous reviewers for very useful comments and suggestions.
+We also thank the Lorentz Center (Leiden, the Netherlands) for hosting the workshop “Purpose-driven particle systems” in March 2023, which enabled discussions about this work.
+FH is supported by start-up funds at the California Institute of Technology and by NSF CAREER Award 2340762.
 FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM.
+UV is partially supported by the European Research Council (ERC) under the EU Horizon 2020 programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO).
 […]